### v1

# XTransfer: Cross-Modality Model Transfer for Human Sensing with Few Data at the Edge

Abstract
Deep learning for human sensing on edge systems offers significant opportunities for smart applications. However, its training and development are hindered by the limited availability of sensor data and resource constraints of edge systems. Current methods that rely on transferring pre-trained models often encounter issues such as modality shift and high resource demands, resulting in substantial accuracy loss, resource overhead, and poor adaptability across different sensing applications. In this paper, we propose XTransfer, a first-of-its-kind method for resource-efficient, modality-agnostic model transfer. XTransfer freely leverages single or multiple pre-trained models and transfers knowledge across different modalities by (i) model repairing that safely repairs modality shift in pre-trained model layers with only few sensor data, and (ii) layer recombining that efficiently searches and recombines layers of interest from source models in a layer-wise manner to create compact models. We benchmark various baselines across diverse human sensing datasets spanning different modalities. Comprehensive results demonstrate that XTransfer achieves state-of-the-art performance on human sensing tasks while significantly reducing the costs of sensor data collection, model training, and edge deployment.
1 Introduction
Human sensing refers to the process of capturing and interpreting data related to human activities, behaviors, and physiological states using various sensors (humansensing2023, ; Ystgaard2023, ). With the proliferation of edge devices equipped with sensors, human sensing plays a vital role in edge systems to understand contexts and enable smart applications, ranging from activity recognition to emotion or vital sign detection (Ahamed2018, ). Deep learning (DL) offers robust performance in interpreting sensor data, and its deployment on edge devices improves privacy and reduces bandwidth (Xiaoqiang2019, ; zhang2020, ; Yanjiao2020, ). However, existing DL solutions are often data- and resource-intensive for training and deployment (ondevicedl, ), posing significant challenges for data collection and edge deployment in human sensing (surveyiot, ; iotdatachallenge, ).
Unlike other data modalities (e.g., text or image), collecting human sensing data for training can be costly and even impractical. This is uniquely due to the data that presents brittleness (i.e., noise sensitivity, low SNR), inconsistency (i.e., user variance, scenario changes), and heterogeneity (i.e., differences in hardware configuration) (Lane2015, ; Jeyakumar19, ; Teh2020, ). Collecting such data typically requires ethical approvals, as it involves sensitive information (e.g., location, motion patterns) that raises privacy concerns (Kroger2019, ; zhangipsn, ). Manual annotation of large volumes of sensor data introduces additional complexity and incurs high costs (Song2023fewshot, ; chen2020deep, ). Furthermore, human sensing spans various modalities (e.g., IMU, ultrasound, mmWave radar) and data types (e.g., spectrograms, Doppler profiles, time series), thereby adding complexity to model training, adaptation, and edge deployment (humansurvey, ; humansensing2023, ).
Prior research has explored Few-shot Learning (FSL) (metasense, ; Wang2024, ), transfer learning (tlhar, ), and their combination (i.e., cross-domain FSL) (crossHAR, ; FewSense, ) to enable DL-based human sensing with few data. However, they either still require large-scale labeled datasets from the same singular modality to costly train source models from scratch (Song2023fewshot, ), or face challenges such as modality shift and limited adaptability when leveraging pre-trained models (Jaehoon2022, ; Ran2022, ; Sommelier2022, ), particularly in cross-modality settings.
These limitations highlight the need for a scalable model transfer that can broadly adapt pre-trained models across modalities (e.g., vision or text) to new sensing modalities (e.g., radar or bio-signals) with few data. Recent advances in multi-modality learning (mmfi, ; chen2025xfi, ; chen2024probabilistic, ; madaan2024jointly, ; yang2024facilitating, ; baldenweg2024multimodal, ) with methods such as knowledge distillation (e.g., image-to-sensor transfer (zhao2018through, ; Xingzhe2022, ; Sevgi2020, ; Yu2023, )) and multi-modal model (e.g., CLIP-based models (girdhar2023imagebind, ; wu2023nextgpt, )), demonstrate the growing feasibility of cross-modality model transfer, driven by the increasing availability of high-quality pre-trained models from public repositories (e.g., PyTorch Hub (PyTorchhub, )), which opens new opportunities for real-world human sensing. However, due to substantial differences in data characteristics—such as shape, distribution, and feature representation—between source and target modalities (i.e., modality shift), transferred models often suffer from significantly degraded performance (i.e., negative transfer) (Sommelier2022, ; Zhang2023, ; Jaehoon2022, ). Our experiments reveal that state-of-the-art baselines (Tab. 1), though effective on vision and NLP tasks, fail to reach oracle performance (metasense, ) and suffer from substantial overfitting when applied to human sensing tasks with few data, falling notably short of the goal, as shown in Fig. 2(a). These highlight the underlying challenge of latent feature misalignment across modalities. Also, these methods often neglect resource constraints, rendering them impractical for edge deployment, especially in advanced multi-source settings (Sicheng2020, ; Xiangyumfs, ; NEURIPS20196048ff4e, ). While methods such as neural architecture search (NAS) (searchinsights, ; AdaptiveNet2023, ; LegoDNN2021, ), pruning (Anwar2017, ; frankle2018the, ; pers2020composition, ), or quantization (Polino2018ModelCV, ), aim to reduce resource demands, they often fail to maintain performance under modality shift and few data. As shown in Fig. 2(c), our experiments using structural pruning (frankle2018the, ; NEURIPS2022, ) demonstrate significant model accuracy loss (50.5% on average). These challenges point to a critical gap in developing adaptable and scalable methods.
To overcome the challenges, we propose XTransfer, a first-of-its-kind cross(X)-modality model Transfer method that freely leverages public source models and transfers knowledge across different modalities (e.g., image, text, audio, or sensing) to human sensing tasks, significantly reducing the need for extensive sensor data collection and training from scratch. It also offers resource-efficient designs to optimize edge deployment, paving the way for scalable real-world applications. XTransfer uniquely repairs, searches and recombines layers of interest from pre-trained source models under cross-modality FSL settings (see App. A), resulting in compact and adaptable models. To mitigate modality shift and prevent overfitting, it enables model repairing achieved by a novel Splice-Repair-Removal (SRR) pipeline (c.f. Sec. 4). This pipeline transforms a few sensor data to broadly adapt pre-trained model layers by repairing layer-wise misalignment and harmonizing latent feature distributions across modalities. XTransfer also features layer recombining using a novel Layer-Wise Search (LWS) control mechanism (c.f. Sec. 5) to identify and recombine layers of interest by conducting a layer-wise search. Building on the insight that not all repaired layers contribute, this mechanism discards such layers and recombines only useful ones to optimize model compactness using principles from NAS (searchinsights, ). To further accelerate the search when dealing with multiple sources, it integrates a pre-search check strategy to efficiently avoid unnecessary repairing.
Comprehensive evaluations (c.f. Sec. 6) are conducted on 8 public datasets along with their corresponding pre-trained source models, spanning diverse modalities (e.g., image, text, audio, and sensing), 4 collected target sensing datasets in real-world settings, 2 public sensing datasets and 1 standard image dataset as targets (Tab. 2) on 3 commercial edge devices. The results show that XTransfer achieves a significantly enhanced accuracy-to-resource (ATR) ratio (c.f. Sec. 6.1), outperforming single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively, while using few data with broad adaptability. XTransfer also streamlines the layer search process, reducing search time by 2.1 to 4 times. The main contributions of this work are:
  • A first-of-its-kind method, XTransfer, that enables resource-efficient, modality-agnostic model transfer for human sensing with few data at the edge, as shown in Fig. 3.
  • A novel SRR pipeline that effectively mitigates modality shift and prevents overfitting, harmonizes latent feature distributions across modalities.
  • A novel LWS control mechanism that efficiently searches layers of interest from multiple sources in a layer-wise manner, creating compact models.
  • Extensive experimental evaluations, benchmarking state-of-the-art baselines based on various source and target datasets, validating competitive scalable performance of XTransfer.
Figure 1: Preliminary study. (a) reveal baseline performance gap  [^2] We test a pre-trained ResNet18 on miniImageNet and HHAR dataset as target under a 5-shot setting (see App. A).(overfitting, ). (b) shows the average similarity and FSL difficulty across all sensing datasets (Tab. 2) to each source modality (e.g., Image, Text, Sensing) (c.f. Sec. 3). 2 distinct areas represent similarity levels (A–hard, B–normal). Key findings: 1) compared to CUB, similarity levels across modalities are notably low, e.g., Text and Sensing fall into Area A, indicating a significant modality shift; 2) compared to Image-Unpair (i.e., no class pairing) in Area A, Image surprisingly falls into Area B, indicating that pairing classes may enhance cross-modality similarity (motivating our techniques in Sec. 4.1); 3) compared to Text and Sensing, Image exhibits more stable standard deviations and lower FSL difficulty, suggesting better potential for broad adaptability (c.f. Sec. 4). (c) shows model optimisation bottlenecks.
2 Related Work
Learning with few data for human sensing. Recent works in FSL and cross-domain FSL have proposed single- and multi-source methods. Single-source methods align latent features between source models and target data via distance-based loss functions (Zhao2021, ; Yue2021Prototypical, ), whereas multi-source methods leverage the Wisdom of the Crowd principle through source distillation (Sicheng2020, ; Xiangyumfs, ) and maximal correlation analysis (NEURIPS20196048ff4e, ). Existing attempts on human sensing tasks (metasense, ; Wang2024, ; tlhar, ; crossHAR, ; FewSense, ) remain constrained by a strong reliance on learning in the same modality, resulting in costly source data collection and model training with limited adaptability. Recent efforts have also explored fine-tuning large language models (e.g., GPT2 (zhou2023one, )) for time-series analysis. However, testing on benchmarks (Boosting2021, ; Jaehoon2022, ; chen2018a, ), our results reveal that they still suffer from severe overfitting under modality shift and few data. Also, none of them explicitly considers resource constraints of edge systems. In contrast, XTransfer enables model repairing to safely mitigate modality shifts using only few sensor data, enabling modality-agnostic model transfer with broad adaptability.
Resource constraints optimization. Model compression techniques such as model pruning (Anwar2017, ; frankle2018the, ; pers2020composition, ), quantization using low-precision data types (Polino2018ModelCV, ), model merging by fine-tuning shared layers (Gemel2023, ), one-size-fits-all model (cai2020once, ), and NAS techniques such as model structure optimization (AdaptiveNet2023, ; LegoDNN2021, ) or search using few data (neuralFT, ; xu2022fewshot, ), have shown success in reducing model resource overhead. However, they strongly rely on either sufficient target datasets or minimal modality shifts to maintain model performance. XTransfer uniquely applies layer recombining to optimize model structures by efficiently searching and recombining layers of interest from pre-trained source models. It not only creates compact models for streamlined edge deployment but also further enhances model repairing.
3 Preliminary
Insights in cross-modality. We apply a standard cross-domain FSL benchmarking (Jaehoon2022, ) to gain insights. It uses the Earth Mover’s Distance (EMD) to measure similarity between source models from 3 modalities (i.e., Image, Text, Sensing) and targets, including 5 human sensing target datasets and 2 well-known image datasets (Tab. 2). It also assesses FSL difficulty by fine-tuning source models on targets. Since source models involve a large number of classes  [^3] The number of classes in the source model should be greater than that in the target. (e.g., 100 classes in miniImageNet), we propose a class pairing strategy that pairs source and target classes exhibiting similar latent feature distributions across modalities. The key insights are shown in Fig. 2(b).
Layer-wise analysis. We perform a layer-wise analysis to gain insights, focusing on accuracy changes and latent feature distributions at each layer. Since effective training relies on learning discriminative features (islam2021broad, ), we benchmark latent feature distribution clustering and use Mean Magnitude of Channels (MMC) (i.e., channel importance) (pmlrv162luo22c, ) of each layer to indicate layer-wise performance. We define a single layer or a dependent layer block (i.e., L-Blocks in Tab. 3)—where layers are not structurally separable—as the default segmentation unit (i.e., L-unit). Fig. 2(a) shows that negative transfer (Zhang2023, ) disrupts accuracy convergence across layers (i.e., layer-wise dependence), causing layer-wise misalignment and accuracy drops. These misalignments correlate with MMC shifts, where greater shifts correspond to larger accuracy drops shown in Fig. 2(b), indicating inconsistent latent feature interpretation during fine-tuning. To track MMC shifts, we essentially employ the Silhouette score (S-score) (shahapure2020cluster, ), which captures inter- and intra-class distances and reflects the discriminability of layer-wise latent features.
Challenges. We aim to minimize layer-wise MMC shifts, but several technical challenges arise. First, the pre-trained feature space at each layer is highly complex due to the multidimensional nature of MMC, and this complexity increases with layer depth. Such complexity poses challenges for "repairing" latent feature distributions across modalities using few sensor data. Second, strong layer-wise dependencies, formed during source model training, become increasingly rigid in deeper layers, posing challenges for standard techniques (e.g., N-Pair loss (Sohn2016, ), Triplet loss (Schroff2015, )) that may fail to preserve these dependencies. Third, Fig. 2(c) shows that S-scores may stagnate or even drop at certain layers (i.e., ineffective layers), impairing overall convergence, indicating that not all repaired layers contribute positively. This motivates our LWS control design, which selectively recombines effective layers (i.e., layers of interest) while discarding others. However, such selection must account for on-device resource constraints to optimize the layer growth. The search space also expands rapidly with the number of layers or source models, further complicating efficient search.
Figure 2: Design insights. (a) Layer-wise accuracy convergence using baselines is disrupted due to modality shift. (b) A notable MMC shift emerges and grows with increasing layer index, i.e., accuracy increases while MMC shift drops in Area A and begins to drop as MMC shift largely grows in Area B. (c) After repairing, layer S-score improves, but stagnation occurs at certain layers.
4 Model Repairing
Enabled by SRR pipeline, this process safely repairs layer-wise misalignment by minimizing MMC shifts, effectively restoring layer performance from source models using few sensing inputs in cross-modality settings. SRR pipeline involves three key stages shown in Fig. 3. In the Splice stage, input shapes between source models and sensor data must be aligned well, as default reshaping (c.f. App. B.1)—using fixed down/up-sampling—cannot be fine-tuned and may lose important features across channel, height, and width, especially when layer shapes vary during layer-wise search. To enhance, we design a compact connector placed between sensing inputs and layers, consisting of a Pre-header (adaptive convolutional layer), Resizer, and one encoder-decoder pair to bridge cross-modality structures. In the next Repair stage, each connector is fine-tuned to generate adaptive channels, preserve key sensing features, and minimize MMC shift via the generative transfer module. In the following Removal stage, channel removal is designed to further enhance layer-wise performance while optimizing resource usage. For more design details, see Appendix B.2.
Figure 3: XTransfer overview. LWS control segments source models into layers and uses the pre-search check to decide if repairing is needed. SRR pipeline then fine-tunes connectors to repair selected layers. Finally, LWS control selects and recombines layers of interest into a compact model.
4.1 Repairing in complex feature space
Since the original latent features from source models are well-trained for high discriminability (c.f. Sec. 3), our intuition is to safely fine-tune the misinterpreted latent features of few sensor data by aligning with the original features (i.e., anchors) to minimize MMC shifts. We hence propose anchor-based cross-modality supervision alignment, which aligns source MMCs as anchors (i.e., anchor MMCs) with sensing MMCs. To handle complex feature spaces, our insight is to reduce MMC dimensionality by preserving key channels and excluding irrelevant ones.
Reduced-dimensional feature space. We project anchor MMCs into a lower-dimensional feature space using principal component analysis (PCA) (VALPOLA2015143, ) (i.e., anchor PCA space). PCA enables a lower-dimensional subspace (i.e., adjusted by the number of PCA components, with 2 components used by default) where the projected data exhibits the highest possible variance (i.e., highly clustered). The optimized component weights (i.e., projection coefficients) highlight the most important channels that contribute to the performance of each layer. Given high accuracy achieved in each original layer, the projected anchor MMCs show high clustering performance, e.g., the centroid of the distribution of each class (i.e., Anchor-center) clustered tightly, as shown in Fig. 3(a). Once the anchor PCA space is initialized for each layer, we project sensing MMCs into its space using the same parameters.
Feature space alignment. After projecting sensing MMCs, we observe that the sample distribution for each class, along with their centroid (i.e., Sensing-center), appears disorganized compared to the well-clustered Anchor-center at each layer, as shown in Fig. 3(a). Specifically, the scale and orientation of Sensing-center are notably different from those of Anchor-center, caused by MMC shifts. To align the scale, we first define a scale function based on the mean shift in inter-class distance $\mathrm{InterD}(\cdot)$, formulated as $S={{Mean}({InterD}(Pro(fs_{ij})))}\mathbin{/}{{Mean}({InterD}(Pro(ft_{ij})))}$, where $Pro(\cdot)={PCA}({MMC}(\cdot))$ denotes the space project function, $fs$ and $ft$ represent the latent features from both source and sensing modalities, respectively. $i$ denotes the index of pre-trained models, and $j$ denotes the layer index in each model $i$. Next, to align the orientation, we design a rotation alignment function that minimizes the cosine angle (nguyen2010cosine, ) by fine-tuning a multi-dimensional rotation matrix $[M_{rot}]$ (where the dimensions equal to the number of PCA components) as below:
$\footnotesize\begin{split}\arg\min_{[M_{rot}]}Cos(Cen(Pro(fs_{ij})),Cen(Pro(ft_{ij})*S)\cdot[M_{rot}])\end{split}$ (1)
where $Cos(\cdot)$ denotes the cosine similarity function, and $Cen(\cdot)$ calculates the centroid of the projected distribution. Fig. 3(b) shows the aligned results.
Anchor class pairing. Building on key finding that class pairing boosts latent feature alignment across modalities (c.f. Sec. 3), we propose selecting source classes with the highest S-scores as anchor classes to pair with sensing classes to facilitate the supervision alignment. We minimize pairing shift (i.e., the sum distance across each pair) via a linear sum assignment solved using the standard Hungarian algorithm, outputting the optimal pairing set $pairST$.
4.2 Repairing layer-wise dependence
Generative transfer module. To effectively restore layer-wise dependence, we design a generative transfer module that enhances the supervision alignment through generative learning. It consists of a generator (i.e., connector) and a discriminator—each frozen pre-trained layer paired with its anchor PCA space (Fig. 3). Freezing preserves stable anchors and prevents overfitting. For each layer, given the optimal pairing set $pairST$, the generator is fine-tuned using the discriminator’s loss to minimize MMC shifts, and uniquely transforms sensor inputs to align with anchors.
Anchor-based repair loss. Building on the generative transfer module, we propose an anchor-based repair loss function ($loss_{srr}$) to harmonize latent feature distributions across modalities. Given each layer’s anchor PCA space, our objective is twofold: 1) minimize the Euclidean distance (ED) ($D(\cdot)$) between the centroid of projected anchor MMC distribution and that of the projected sensing MMC distribution for each class pair $(c_{s},c_{t})\in pairST$ (i.e., positive loss), and 2) maximize the distance between sensing MMC samples with different labels $(c_{t}^{-})$ (i.e., negative loss). Since the projected anchor MMCs are highly clustered, it can provide the maximum margin ($M_{max}$) between anchor clusters. To further improve the negative loss, we update the function by calculating the anchor-based margin $margin^{c}=InterD(Pro(fs^{c}_{ij}))-IntraD(Pro(fs^{c}_{ij}))$, where the $M_{max}$ is selected, and $IntraD(\cdot)$ measures the intra-class distance. Hence, our loss function is formulated as:
$\footnotesize\begin{split}loss_{srr}=&\frac{1}{N}\sum^{N}_{(c_{s},c_{t})\in pairST}D(Cen(Pro(fs^{c_{s}}_{ij})),Cen(Pro(ft^{c_{t}}_{ij})))\\ +&\frac{1}{K}\sum^{K}_{(c_{t},c_{tk}^{-})}ReLU(M_{max}-D(Pro(ft^{c_{t}}_{ij}),Pro(ft^{c_{tk}^{-}}_{ij})))\end{split}$ (2)
where $K$ denotes the number of negative samples, and $N$ denotes the number of classes. The $ReLU(\cdot)$ is used to trigger the negative loss in the condition where the distance between the samples is less than $M_{max}$. Fig. 3(c) shows the effectiveness of model repairing.
5 Layer Recombining
Working in conjunction with SRR pipeline for repairing each layer, XTransfer leverages layer recombining enabled by LWS control to: 1) perform the pre-search check that accelerates the search process, and 2) efficiently select and recombine layers of interest to form a compact model.
5.1 Search control design
Following NAS principles (searchinsights, ), poor design choices (e.g., search space, actions, strategy) can disrupt the balance between model performance and resource efficiency. As layers of interest grow, resource overhead can increase exponentially, making dynamic search control essential under on-device resource constraints. To efficiently manage, we define the search space as all layers from source models, structured into search pools containing layer candidates (i.e., repaired layers), each with width $I$ (number of models) and depth $J$ (number of L-units), forming a window of size $I\times J$. Search actions (init, continue, skip, cross) are defined to guide layer recombination within each designated search pool. Init selects the starting layer with a default fixed depth of 1. Continue and skip guide intra-model layer recombination, while cross enables inter-model layer recombination.
Search strategy under resource constraints. To effectively control search actions within each search pool, we propose a layer value function based on the S-score $Score(\cdot)$ to evaluate both individual repaired layer performance and the convergence between adjacent layers after each SRR process $SRR(\cdot)$. It is formulated as $V_{ij}=Score(SRR(Pro(ft_{ij}),MT_{ij},L_{ij}),Y_{t}))$, where $MT_{ij}$ denotes the connector and $L_{ij}$ represents each repaired layer. Under resource constraints, only the top-performing layer in each pool is selected. LWS control then checks whether its S-score exceeds that of the previous layer. If successful, the layer $L_{ij}$ is identified as the layer of interest and recombined, and the next pool window begins at layer index $j+1$. Otherwise, the current pool is discarded, and the search window shifts forward. Considering the constraints, we design a resource-constrained function that works in conjunction with $V_{ij}(\cdot)$, which is formulated as $RC(n)=exp(n/(L-2)-2)+1,n\in[1,~{}L]$, where $n$ denotes the current recombined layer index, and $L$ denotes the maximum number of L-units. Each layer’s adjusted overhead, $R(n)_{ij}$, is computed as $R(n)_{ij}=(Res(L_{ij})*RC(n))/Max(Res(P))$, where $Res(\cdot)$ calculates actual cost (e.g., FLOPs), and $Max(Res(P))$ is the peak cost in current pool $P$. The final resource-constrained layer value function becomes $VR_{ij}=V_{ij}\mathbin{/}R(n)_{ij}$. The objective is then defined as $\arg\max_{L_{ij}\in P}VR_{ij},\,s.t.\sum_{0}^{n}Res(L_{ij})\leq R_{Max}$, where $R_{Max}$ is the total allowed resource budget.
5.2 Efficient search
Pre-search check. LWS control is designed to handle large search spaces from deep or multi-source models. While fine-tuning the connectors in SRR pipeline is efficient, the large search space can considerably slow down LWS control. We hence propose the pre-search check strategy to avoid unnecessary repair steps before applying SRR. We track S-score progression across three states—after anchor pairing (Anchor) (i.e., repairing objective), before SRR (Before), and after SRR (After) (i.e., repaired). SRR pipeline consistently improves S-scores from Before to After, approaching the Anchor level, indicating a strong correlation. Based on this, our intuition is to estimate repaired S-score (i.e., repair rate) of each layer based on Before and Anchor states to decide whether to bypass unnecessary repairs. Experiments show the actual repair rate grows exponentially with the recombined layer index (see App. B.4). We hence design a repair rate growth model formulated as $rate_{n}^{est}=exp(a*n)+b,n\in[1,L]$, where $a$ and $b$ are optimized via non-linear regression to minimize MSE to fit $rate_{n}^{est}$, with the objective function defined as $\arg\min_{a,b}\frac{1}{n^{{}^{\prime}}}\sum_{n=0}^{n^{{}^{\prime}}}(rate_{n}^{est}-rate_{n})^{2}$, where $n^{{}^{\prime}}$ is the number of recombined layers. The estimation begins after init. Once a new layer is selected, LWS control updates the estimation. Due to the limited updates, the estimation may be inaccurate. To mitigate, we also develop a dynamic search range mechanism that dynamically adjusts the search range in each pool (details in App. B.5).
6 Experiment
6.1 Experiment setup
Table 1: State-of-the-art baselines.
  Baselines Group Description
  TL (TransferabilityID, ) Single-source A standard transfer learning method by freezing all layers
FT (Jaehoon2022, ) Single-source A standard fine-tuning method by updating all layers
ProtoNet (Snell2017, ) Single-source A distance-based FSL method
MAML (Finn2017, ) Single-source A learning-based FSL method
DAPN (Zhao2021, ) Single-source A cross-domain FSL method using adversarial ProtoNet
GPT2 (zhou2023one, ) Single-(LLM) A large language model based method using fine-tuning
MDDA (Sicheng2020, ) Multi-source A cross-domain FSL method using knowledge distillation
MCW (NEURIPS20196048ff4e, ) Multi-source A cross-domain FSL method using maximal correlation analysis
MetaSense (metasense, ) Oracle-(GT) An advanced FSL method based on MAML with sufficient target data
Table 2: Source and target datasets specifications.
  Dataset Modality Subject Class Type Shape Privacy
  Source
miniImageNet (miniimagenet, ) Image - 100 - 3x84x84 Public
miniDomainNet (domainnet, ) Image - 126 - 3x224x224 Public
Office-31 (office31, ) Image - 31 - 3x224x224 Public
Office-Home (officehome, ) Image - 65 - 3x224x224 Public
Caltech-101 (caltech101, ) Image - 101 - 3x224x224 Public
Newsgroup (text2021, ) Text - 20 - 100x1x1000 Public
VoxCeleb (VoxCeleb2017, ) Audio - 100 Time-S 1x512x300 Public
OPPORTUNITY (Roggen2010, ) IMU 12 11 Time-S 77x1x100 Public
Target
HHAR (Stisen2015, ) IMU 9 6 Time-S 6x1x256 Public
WESAD (Schmidt2018, ) ECG 15 3 Time-S 10x1x256 Public
Gesture IMU/PPG 10 8 Time-S 7x1x200 Private
Writing Ultrasound 10 10 Time-S 1x1x256 Private
Emotion mmWave 10 7 Spectrum 3x224x224 Private
BP mmWave 10 - Doppler 1x1x499 Private
ChestX (Wang2017, ) Image - 5 - 1x1024x1024 Public
  Backbone Layers L-Blocks GFLOPs Params Kernel Shape
  ResNet18 (chen2018a, ) 18 9 3.67 11.18M 2D 3x224x224
ResNet18-1D (chen2018a, ) 18 9 0.39 3.9M 1D 100x1x1000
Multi-ResNet18 (Sicheng2020, ) 18x5 9 18.35 55.9M 2D 3x224x224
Conv4 (chen2018a, ) 4 4 0.02 0.07 1D 1x1x500
Conv5 (metasense, ) 5 5 0.19 30.42M 1D By Input
GPT2 (zhou2023one, ) 4x6 6 0.41 82.9M 1D By Input
Table 3: DL model backbone specifications.
We benchmark XTransfer against state-of-the-art baselines, grouped into single- and multi-source methods (Tab. 1). Evaluation is conducted on 8 widely-used public datasets for cross-modality model transfer (Tab. 2), using their corresponding public pre-trained models as sources, from diverse modalities (e.g., image, text, audio, and sensing) with standard backbones such as ResNet18 from PyTorch Hub (PyTorchhub, ) (Tab. 3). To evaluate XTransfer, we develop 4 real-world applications and build three testbeds on 3 commercial edge devices (e.g., smartwatch, smartphone, raspberry pi). Based on these, we collect 4 human sensing datasets (marked as Private in Tab. 2) as targets, from 40 participants in real-world settings, with ethics approval from the Human Research Ethics Committee of our institute. Additionally, we include 2 public sensing datasets and 1 image dataset as targets for standard benchmarking. We apply standard Leave-One-Out Cross-Validation (LOOCV) (metasense, ) for evaluating human sensing datasets, and adopt accuracy-to-resource (ATR) ratio, defined as $ATR=Accuracy\mathbin{/}(\alpha Norm(FLOPs))+(1-\alpha)Norm(Params))$ for evaluating resource-accuracy efficiency, where $\alpha$ is set to 0.5 by default. For more details, see Appendix A.
6.2 XTransfer performance
Figure 4: Ablation study evaluating the performance of model repairing and layer recombining.
Repair performance. We first evaluate how our Repair loss reduces MMC shift to enhance the S-score of each layer with different MMC dimensionalities in the SRR pipeline. We use HHAR and the standard pre-trained ResNet18 on miniImageNet in a 5-shot setting. The channel removal is disabled in this experiment. We test with 4 different levels of MMC dimensionality (i.e., number of channels) ranging from 64 to 512. We employ N-Pair loss (Sohn2016, ) and Triplet loss (Schroff2015, ) as baselines. We run 15 rounds for each setting by default throughout the evaluation. Fig. 4(a) presents that our Repair loss consistently achieves the highest S-score of 0.2 on average, outperforming both N-Pair loss (0.05) and Triplet loss (0.07), across all levels of MMC dimensionality. It indicates that Repair loss effectively enhances layer performance.
Efficient search performance. We now evaluate LWS control performance regarding model accuracy and layer search efficiency. We employ HHAR as target, ResNet18 on miniImageNet and Multi-ResNet18 on 5 image datasets as sources shown in Tab. 2, with a default search depth of 3 in 3-, 5- and 10-shot settings. We also utilize MetaSense in Tab. 1 as the oracle baseline. We compare different configurations: ResNet18 (Single), Multi-ResNet18 (Multi) without efficient search enabled, and Multi-ResNet18 (Multi-Efficient) with efficient search enabled. As expected, Fig. 4(b) shows that Multi-Efficient achieves a substantial time reduction ranging from 2.1 to 4 times compared to Multi in 5- and 10-shot settings, while outperforming the oracle baseline. Additionally, Multi-Efficient improves average accuracy by 7.8% over Multi across the three settings. While Multi-Efficient takes a longer search time than Single, it successfully achieves superior accuracy, highlighting the benefits of using multiple source models in XTransfer. The results also reveal that both Multi-Efficient and Multi achieve lower standard deviation on average compared to Single, suggesting that using multiple source models offers commendable search stability. In short, the results prove that the proposed efficient search significantly accelerates the layer search process and achieves better model accuracy.
Impact of different sources. Since the selection of pre-trained source models plays a crucial role in XTransfer, we evaluate how different source models from various modalities affect the accuracy and resource overhead of the output models. We use HHAR, 2 pre-trained ResNet18s on both miniImageNet (Image) and VoxCeleb (Audio), and 2 pre-trained ResNet18-1Ds on both Newsgroup (Text) and OPPORTUNITY (Sensing), as shown in Tab. 2. We keep the other experimental setups unchanged. Fig. 4(c) shows that the output models using Image source achieve the highest accuracy of 70.5% on average, compared to that of 69.6%, 64.3%, and 67% achieved by the models using Text, Audio, Sensing as sources, respectively. This suggests that the latent features learned from Image have better discriminability than those from other modalities. Since XTransfer prioritizes accuracy, we select Image source models by default. More ablation study results refer to Appendix C.
6.3 Comprehensive comparison
Cross-modality model transfer performance. We evaluate the overall performance of XTransfer regarding both accuracy and ATR across 5 sensing datasets and 1 image dataset as targets, as shown in Tab. 2. We compare XTransfer against a range of state-of-the-art baselines shown in Tab. 1. We also mark our methods as Our-Single and Our-Multi, using pre-trained ResNet18 on miniImageNet and Multi-ResNet18 with 5 corresponding image datasets as sources. In addition, we evaluate the fine-tuning performance of GPT2-small for comparison and align the parameters with the work (zhou2023one, ). All methods on target datasets are evaluated under the same cross-modality FSL settings (see App. A), with MetaSense (metasense, ) serving as the oracle baseline, trained on Conv5 with sufficient target sensor data (Tab. 3). For consistency, we also train the oracle using ResNet18 on the Emotion dataset to match 2D kernels. As shown in Tab.4 and 5, both our methods outperform the baselines regarding model accuracy and ATR, and successfully reach or surpass the oracle in model accuracy across all target datasets in 5- and 10-shot settings. Specifically, both Our-Single and Our-Multi achieve a significantly higher ATR on average, surpassing the existing single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively. Interestingly, in 3-shot setting, the averaged model accuracy of our methods is slightly lower than that of the oracle on the HHAR, Gesture, and Emotion datasets. This suggests that XTransfer’s Repair loss may lead to less stable improvements for latent feature alignment when using extremely low sample sizes, but its performance remains strong and stable in 5- and 10-shot settings. In summary, the comprehensive results show that XTransfer achieves state-of-the-art model accuracy, substantially reducing the cost of data collection and model resource overhead for scalable deployment of human sensing applications.
Training resource usage. We also examine the training resource usage of XTransfer by measuring GPU utilization (Util) on the cloud. We focus on the above baselines that require model re-training or adaptation for transfer, e.g., MAML (Finn2017, ) and DAPN (Zhao2021, )), excluding methods that rely solely on fine-tuning or distance-based techniques. The experimental setup remains the same as above. As shown in Tab. 6, both our methods achieve the lowest GPU Util across all n-shot settings compared to baselines. Notably, both Our-Single and Our-Multi achieve a substantial reduction in GPU Util by 2.6 to 3.6 times and 2.5 to 3.5 times on average, respectively, compared to baselines. This is due to our layer-wise method, which progressively repairs and recombines only a subset of layers rather than training the entire model at once. It also highlights the effectiveness of our compact connector design. Despite leveraging multiple source models, Our-Multi shows no significant increase (e.g., only 3.67%) in training resource usage, further highlighting its efficiency. Hence, XTransfer significantly reduces cloud resource usage, enabling scalable model transfer with low computational overhead.
HHAR (%) WESAD (%) Gesture (%) Writing (%) Emotion (%) ChestX (%)
Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
MetaSense (metasense, ) 63.2 69.0 75.2 55.1 64.0 68.3 66.3 73.4 79.4 72.1 83.3 89.6 55.5 56.3 65.5 24.8 28.1 31.7
ProtoNet (Snell2017, ) 50.6 45.7 49.8 43.9 49.3 52.0 50.8 55.8 59.2 69.5 78.7 73.4 51.3 49.2 54.8 21.7 23.4 24.5
DAPN (Zhao2021, ) 48.7 51.2 54.7 56.9 61.2 63.1 45.8 49.4 50.5 75.1 78.6 79.6 48.0 50.0 58.7 24.0 24.4 25.5
MAML (Finn2017, ) 36.6 42.3 42.4 39.6 57.9 65.6 37.2 41.3 39.4 31.6 39.7 35.2 22.3 26.9 31.1 22.5 20.4 21.9
GPT2 (zhou2023one, ) 41.0 34.0 45.0 48.0 35.0 36.0 49.0 54.0 63.0 52.0 71.0 75.0 - - - - - -
Our-Single 59.0 71.8 80.7 77.9 78.4 81.8 45.2 69.6 77.8 72.8 87.0 91.3 43.6 55.6 61.4 27.7 28.6 30.4
MDDA (Sicheng2020, ) 13.2 17.7 18.8 40.0 39.3 41.9 11.2 13.9 15.2 12.5 10.9 9.0 14.4 15.3 16.0 20.6 19.6 19.9
MCW (NEURIPS20196048ff4e, ) 27.6 25.4 26.9 30.7 32.7 32.4 22.3 23.4 25.6 32.3 31.6 30.2 25.3 27.2 26.5 22.4 21.5 22.8
Our-Multi 59.3 74.3 78.4 76.4 77.8 78.7 48.7 73.1 78.6 74.1 86.1 90.4 42.3 55.1 58.6 29.0 30.0 30.2
Table 4: Comparison in model accuracy. Top-2 results are highlighted in bold.
HHAR WESAD Gesture Writing Emotion ChestX
Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
ProtoNet (Snell2017, ) 0.51 0.46 0.50 0.44 0.49 0.52 0.51 0.56 0.59 0.69 0.79 0.73 0.51 0.49 0.55 0.22 0.23 0.25
DAPN (Zhao2021, ) 0.49 0.51 0.55 0.57 0.61 0.63 0.46 0.49 0.51 0.75 0.79 0.80 0.48 0.50 0.59 0.24 0.24 0.26
MAML (Finn2017, ) 0.37 0.42 0.42 0.40 0.58 0.66 0.37 0.41 0.39 0.32 0.40 0.35 0.22 0.27 0.31 0.22 0.20 0.22
GPT2 (zhou2023one, ) 0.11 0.09 0.12 0.13 0.09 0.09 0.13 0.14 0.17 0.14 0.19 0.20 - - - - - -
Our-Single 1.42 1.57 2.09 2.76 2.61 2.51 1.55 1.90 1.76 2.11 1.88 1.84 0.82 1.07 1.26 0.83 1.04 1.24
MDDA (Sicheng2020, ) 0.03 0.04 0.04 0.08 0.08 0.08 0.02 0.03 0.03 0.03 0.02 0.02 0.03 0.03 0.03 0.04 0.04 0.04
MCW (NEURIPS20196048ff4e, ) 0.06 0.05 0.05 0.06 0.07 0.06 0.04 0.05 0.05 0.06 0.06 0.06 0.05 0.05 0.05 0.04 0.04 0.05
Our-Multi 1.71 1.47 1.64 3.00 2.70 2.75 1.20 1.25 1.63 1.93 1.96 1.53 0.96 1.23 1.73 0.68 1.24 0.83
Table 5: Comparison in ATR ratio. Top-2 results are highlighted in bold
HHAR (%) WESAD (%) Gesture (%) Writing (%) Emotion (%) ChestX (%)
Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
DAPN (Zhao2021, ) 48.85 49.19 51.79 44.62 49.49 48.81 42.73 48.71 50.59 43.88 56.63 57.51 44.67 59.73 63.12 46.66 47.32 48.48
MAML (Finn2017, ) 59.59 79.32 89.72 52.58 70.27 77.52 53.11 66.75 73.21 72.56 85.47 92.22 63.39 72.78 75.01 44.52 55.46 61.36
Our-Single 18.27 18.90 19.48 15.95 16.46 16.98 20.67 21.71 22.51 19.92 20.67 21.08 18.24 19.20 19.96 16.97 17.58 18.80
Our-Multi 18.31 20.77 22.06 18.00 19.42 23.10 20.49 22.49 24.49 18.68 19.25 21.69 16.45 18.18 19.85 15.33 17.69 19.62
Table 6: Comparison in GPU utilization of each method during training. Top-2 results are highlighted in bold
7 Conclusion
This paper presents XTransfer, a novel method that enables resource-efficient, modality-agnostic model transfer for human sensing applications at the edge. XTransfer harnesses the power of pre-trained models as free sources and requires only few sensor data to create compact models. It significantly reduces the costs associated with sensor data collection, model training, and edge deployment, offering a practical and scalable method for advancing human sensing.
Limitations and future work. Since XTransfer reuses layers from one or more pre-trained source models, its performance depends on their quality (c.f. Sec. 4). While performance varies with different source models, XTransfer’s modality-agnostic design ensures broad adaptability, enabling flexible use of diverse source models without concern for modality shift or resource constraints. For fair benchmarking, we primarily use the pre-trained ResNet18 from PyTorch Hub (PyTorchhub, ) (c.f. Sec. 6.1), but XTransfer’s layer-wise repairing and recombining support flexible backbone choices. While current version requires source models with the same backbone, future versions will support mixed backbones to further enhance flexibility.
Appendix
The appendix is organized as follows:
  • $\bullet$
    Section A illustrates more details on our experiment setup, including implementation details, developed real-world sensing applications, experimental setups, dataset statistics, evaluation metrics, and computation resources we used for our experiments.
  • $\bullet$
    Section B introduces additional technique details of different components in XTransfer.
  • $\bullet$
    Section C gives more experimental results about ablation study and benchmarking.
Appendix A Experiment setup details
A.1 Settings for learning with few data
A typical FSL method consists of two stages–meta-training and meta-testing [56]. In meta-training stage, FSL requires a large-scale source dataset from the same singular modality (i.e., typically specific to the application) to train source models from scratch. The source dataset is divided into a support set and a query set  [^4] The support set and query set are assumed to be drawn from the same data distribution [82].. Data for each set is loaded according to the n-way (i.e., number of classes) and n-shot (i.e., number of samples per class) format [15]. The meta-testing stage follows the same structure but uses unseen target data. Specifically, the target support set is used to adapt the source models, which are then evaluated on the target query set. Leave-One-Out Cross-Validation (LOOCV) is widely used as the de facto standard for evaluating datasets involving multiple users in human sensing [18]. LOOCV works by leaving one user out as the target for meta-testing while using the remaining users to train the source models during meta-training in each round. The MetaSense [18] is used as an oracle baseline under FSL settings, where sufficient target sensing datasets are provided during meta-training stage. Unlike FSL, cross-domain FSL focuses primarily on meta-testing and transfers pre-trained source models across domains (i.e., different users), while remaining within the same singular modality [23]. In this work, the cross-modality FSL settings are defined similarly to cross-domain FSL, but with pre-trained source models selected across different modalities without constraints.
A.2 Setups in preliminary study
As shown in Fig. 2(a), we test a standard overfitting metric, defined as $\frac{\left|Accuracy_{train}-Accuracy_{test}\right|}{Accuracy_{train}}$ [51]. In Fig. 2(b), we test 3 source models (using a pre-trained ResNet18 from PyTorch Hub [38]) from 3 modalities, e.g., miniImageNet (Image), Newsgroup (Text), and OPPORTUNITY (Sensing), shown in Tab. 2. We also use 5 sensing datasets (HHAR, WESAD, Writing, Emotion, and Gesture) as targets, and employ 2 well-known image datasets (CUB and ChestX) as reference targets to the Image, detailed in Tab. 2. We use a default 5-shot cross-modality FSL setting with LOOCV enabled.
A.3 Setups in layer-wise analysis
In Section 3, we conduct a layer-wise analysis to gain insights into cross-modality model transfer with few sensor data. To set up, we continue to use the standard ResNet18 [38] and segment it as 9 L-Blocks, as shown in Tab. 3. We also employ the default reshaping (c.f. App. B.1) on HHAR as target with 5-shot setting, and use the original source model input (miniImageNet) as reference, marked as Original. We examine the single-source baselines including Transfer Learning (TL) [68], Fine-tuning (FT) [23], MAML [70], DAPN [52], and MetaSense [18] as the oracle baseline, shown in Tab. 1. To precisely track the MMC shift changes at each layer during fine-tuning,
A.4 Layer-wise metric
To precisely track the MMC shift changes at each layer during fine-tuning, we essentially use the Silhouette score (S-score) [63] as the layer-wise metric to evaluate how well the features are being fine-tuned. The S-score is in the range from -1 to 1 and being better if closing to 1. To verify whether the S-score indicates the layer performance, we observe the changes in the S-score and accuracy at each layer using the TL [68] and Original methods. Fig. 6(a) also shows that S-score and accuracy by Original are positively correlated, converging in a layer-wise manner. TL’s S-score also performs equivalently to its accuracy (i.e., S-score drop along with accuracy drop) in detecting the negative transfer across different layers.
A.5 Implementation
We implement XTransfer using PyTorch version 1.4.0. We also benchmark XTransfer with a full development cycle of cross-modality FSL tasks, including pre- and post-processing, training, and inference on commercial edge devices. Once pre-trained source models and few sensor data are processed in cloud via XTransfer, compact models can be seamlessly converted to various DL platforms (e.g., PyTorch Mobile, TFLite) for broader edge deployments. To enable reproducibility of the results, we fix all random seeds in the code, set the evaluation mode consistently when evaluating output models, and enable CUDA convolution determinism. To measure model resource overhead (e.g., FLOPs and parameters), we use the standard PyTorch OpCounter. For regression tasks, class-based transformation [83] enables compatibility.
A.6 Developed sensing applications
To evaluate XTransfer on real-world human sensing targets, we essentially develop 4 human sensing applications, as shown in Fig. 5. These include: (1) emotion recognition via facial expressions (i.e., Emotion) and (2) non-contact blood pressure monitoring (i.e., BP) based on an embedded mmWave radar setup, (3) handwriting tracking (i.e., Writing) using ultrasound based on a smartphone, and (4) gesture recognition (i.e., Gesture) using IMU and PPG sensors based on a smartwatch. The data collection is approved with ethics approval and involves 40 subjects. The corresponding target sensing datasets are shown in Tab. 2 with detailed specification.
A.7 Hardware setup and computation environment
We evaluate XTransfer on real-world testbeds using 3 commercial edge devices, ranging from smartwatch to Raspberry Pi (Tab. 7). For example, Fig. 5(a) presents an embedded mmWave radar setup which consists of a TI IWR1843Boost and a Raspberry Pi 4 using ROS Melodic on Ubuntu 18.04.4. In addition, we implement both smartwatch and smartphone testbeds using PyTorch Android version 1.10.0. We also employ servers equipped with NVIDIA GeForce RTX 3090 GPU to build our cloud environment.
A.8 Training details
In the Repair stage, we fine-tune the generative transfer module (c.f. Sec. 4) for 100 episodes using SGD with a learning rate of 1e-2 and momentum 0.95. We apply a step-based learning rate scheduler with a step size of 20 and a decay factor of 0.5. We also use standard early stopping with a patience of 20. We retain the training parameters of the baselines in Tab. 1 as specified in their original implementations.
Figure 5: (a) Embedded mmWave radar testbed setup; (b)-(e) Built human sensing applications across different real-world settings.
  Device ROM RAM CPU Battery OS
  Galaxy Watch 4 16GB 1.5GB Exynos W920 247mAh Wear OS 3.5
Raspberry Pi 4 64GB 8GB Broadcom BCM2711 - Ubuntu 18.04.4
Redmi K20 Pro 128GB 8GB Snapdragon 855 4000mAh Android 11
Table 7: Edge devices specification.
Appendix B Technical details
Figure 6: Design insights. (a) shows layer-wise metric correlation. (b)(c) present efficient search insights into LWS control using multiple source models.
B.1 Default reshaping
To align with source model input shape, we develop a default reshaping to transform sensor data shape. It uses bilinear interpolation [84] (i.e., Resizer) to resize the height and width, and a fixed convolutional layer (i.e., Fixed header) with a $1\times 1$ kernel to match the input channels [85], placed between source models and sensor data.
B.2 PCA-based channel removal
In the Removal stage, channel removal is designed to further enhance layer performance while optimizing resource usage. This is achieved by removing unnecessary channels at each layer. Existing pruning techniques [46, 47] typically rely on channel importance metrics, such as the L2 norm, calculated from MMC. However, due to MMC shift, these metrics may inaccurately assess channel importance, resulting in degraded layer performance. To address this, we propose using PCA-based channel importance, derived from the optimized component weights during the repair process, as a more reliable metric for channel removal. Our objective is to remove unnecessary channels $\widetilde{ch}$ while preserving the maximum S-score ($Score(\cdot)$), formulated as $\arg\max_{(\widetilde{ch}\,\in CH)}Score(PCA(MMC(ft^{r}_{ij}),\widetilde{ch}),Y_{t})$, where $CH$ denotes the set of all channels, $r$ denotes the repaired latent features $ft_{ij}$, and $Y_{t}$ denotes the target labels.
B.3 Setups for efficient search insights
To gain design insights of efficient search for LWS control (c.f. Sec. 5), we obverse a complete process where all layers are repaired using the HHAR dataset and a Multi-ResNet18 model (i.e., 5 source models with 45 L-unit), as shown in Tab. 3. Fig. 6(b) shows the changes in the S-score across three states: after pairing anchors (i.e., Anchor), before and after applying the SRR pipeline (i.e., Before and After, respectively), in a layer-wise manner. The results show that the SRR pipeline successfully increases the S-score from Before state (i.e., lower bound) to align with the Anchor state (i.e., the objective). This indicates a strong correlation between the After state’s S-score and both Before and Anchor values. These insights lead to the observation for designing the repair rate growth model.
B.4 Observation for designing repair rate growth model
Our key idea is to estimate the repaired S-score (i.e., repair value) of each layer before initiating SRR processing, allowing us to decide whether to bypass unnecessary repairs. To empirically examine this, we first define a repair rate function to capture this correlation as $rate_{n}=(After-Before)/Anchor=(V_{ij}^{n}-Score(Pro(ft_{ij}^{n}),Y_{t}))/Score(Pro(fs_{ij}^{n}),Y_{s})$, where $Y_{s}$ denotes the source labels, and we normalize $rate_{n}\in[0,1]$. We then repeat the process multiple times to observe the changes in $rate_{n}$. Fig. 6(c) shows that the repair rate demonstrates exponential growth. Thus, we design the repair rate growth model based on an exponential growth model (c.f. Sec. 5.2).
B.5 Dynamic search range mechanism
Based on the estimated repair rate, LWS control can initially use a small, fixed search range to efficiently filter out most layer candidates before initiating SRR processing. However, because the updates to $rate_{n}^{est}$ are limited, the estimation may be inaccurate, potentially affecting selection decisions (e.g., missing the optimal layer candidate). To mitigate this risk, we develop a dynamic search range mechanism that adjusts the search range to balance accuracy and efficiency. This mechanism is formulated as $range_{n}=(1+\hat{u}^{\pm}rate_{n}^{est})*range_{n-1},\,range_{n}\in[0.2,1]$, where the range starts at 0.5 and adjusts based on the previous range. Here, $\hat{u}$ denotes a signed unit vector. If $rate_{n-1}^{est}$ is less than $rate_{n-1}$, it indicates that the estimation might be inaccurate and $\hat{u}$ will be positive to increase the range in searching for the optimal layer. Otherwise, the range will be reduced to improve search efficiency.
B.6 Post fine-tuning
After LWS control is completed, we add either a linear classifier or a regression layer (e.g., a fully connected layer) at the end of the model to finalize its backbone. To strengthen layer-wise dependence, we perform a quick post fine-tuning process, updating only the connectors and fully connected layers.
Appendix C Additional experiment results
Figure 7: Ablation study evaluating the performance of components, search parameters, and applications.
C.1 Removal performance
We evaluate how channel removal enhances the average S-score of selected layers. We use HHAR and the standard pre-trained ResNet18 on miniImageNet in a 5-shot setting, and enable channel removal with a default search depth of 3. We employ L2 norm-based (i.e., L2-based) model pruning [47, 50] as the baseline and compare it with the setting where channel removal is disabled (i.e., Non-Pruning). Fig. 7(a) demonstrates that the proposed channel removal outperforms both Non-Pruning and L2-based pruning across all settings, achieving an average S-score of 0.1 by selected layers. L2-based pruning achieves a lower average S-score than Non-Pruning, indicating that the channel importance may be inaccurately estimated, resulting in degraded performance. These results show that the proposed channel removal further improves layer performance by accurately removing unnecessary channels at each layer.
C.2 Impact of search depth
In this experiment, we investigate how different search depths affect the accuracy and resource overhead (e.g., FLOPs) of output models. Varying the search depth alters the search window sizes with the number of layer candidates, which can lead to different output model backbones and accuracy. To quantify these differences, we employ HHAR as target, ResNet18 on miniImageNet (Single) and Multi-ResNet18 (Multi) on 5 image datasets as sources (shown in Tab. 2) in 3-, 5- and 10-shot settings. We also utilize MetaSense in Tab. 1 as the oracle baseline. Both Single and Multi models have up to 9 L-unit, hence we set three different search depths as D1, D3, D9. Fig. 7(b) shows that the output models of both Single-D1 and Multi-D1 result in high FLOPs, while that of both Single-D9 and Multi-D9 achieves low FLOPs. As search depth decreases, the search process becomes longer, and it triggers LWS control to make more decisions on recombining layers, which results in a suboptimal model backbone with high FLOPs. Although both Single-D3 and Multi-D3 do not produce the lowest FLOPs on average, they outperform others and achieve the highest average model accuracy of 70.52% and 70.69%, respectively, across the three settings. Therefore, we set D3 as the default search depth, as it provides the best balance between resource overhead and accuracy of the output models.
C.3 Regression task performance
We also evaluate the effectiveness of XTransfer in a regression task using the BP dataset, which includes both systolic (SBP) and diastolic (DBP) sensor data shown in Tab. 2 collected by the blood pressure monitoring application. We employ the pre-trained ResNet18 on miniImageNet as the source model and use the 5-shot setting. To integrate this into the SRR pipeline, we transform the regression task by treating individual regression labels as distinct classes for anchor class pairing. For benchmarking, we use mean error (ME) and standard deviation (STD), following the Advancement of Medical Instrumentation (AAMI) standard, which accepts an ME within 5 mmHg and STD within 8 mmHg [86]. We employ a random forest (RF) based method [87] as the oracle baseline. Besides, we train a Conv4 backbone in Tab. 3 (Conv) using LOOCV, and the fine-tuned Conv (ConvFT) as the baselines. Fig. 7(c) shows that both XTransfer and the RF methods surpass the AAMI standard. In particular, XTransfer outperforms the baselines and RF in estimating both SBP (ME: -0.64mmHg, STD: 7.2mmHg) and DBP (ME: 0.07mmHg, STD: 7.47mmHg) tasks. These results demonstrate that XTransfer is highly efficient in handling regression tasks with few sensor data.
C.4 On-device performance
We now evaluate the on-device performance of deploying XTransfer’s output models on 3 commercial edge devices (Tab. 7), focusing on resource overhead metrics such as FLOPs, parameters, latency per call, and peak memory footprint. For each target dataset, we test both types of output models created by our single-source (Single) and multi-source (Multi) methods. We also use ResNet18 and Multi-ResNet18 (Tab. 3) as baselines. As shown in Tab. 8, all output models achieve significant reductions in model resource overhead, e.g., FLOPs is reduced by 2.4 to 5.7$\times$ for Single models and by 10.1 to 16.5$\times$ for Multi models, compared to the baselines. These recombined models also demonstrate substantial improvement in on-device performance. Specifically, in the Single method, latency is reduced by 1.4 to 4.1 $\times$ and peak memory footprint decreases by 1.8 to 3.9$\times$. In the Multi method, the reductions are even more pronounced, with latency dropping by 7.4 to 21$\times$ and memory footprint by 2.2 to 8.4$\times$, compared to the baselines. These results prove that XTransfer successfully achieves streamlined edge deployment.
Watch Pi Phone
Dataset Type FLOPs Params Latency (ms) Mem (MB) Latency (ms) Mem (MB) Latency (ms) Mem (MB)
HHAR Single 1.51G 0.75M 625.5 125.5 274.18 246.62 26.86 179.5
Multi 1.46G 2.16M 634.5 140 249.26 247.83 28.34 209
WESAD Single 1.19G 0.74M 604.5 120 240.05 244.64 25.77 168
Multi 0.89G 0.34M 460.5 122 209.52 235.55 21.57 181
Gesture Single 1.21G 2.25M 706.5 135 259.25 251.26 31.5 192
Multi 1.45G 2.09M 692 137 268.79 253.7 30.96 190
Writing Single 1.32G 0.87M 467.5 126 261.71 242.33 29.79 183.5
Multi 1.28G 0.73M 513 124.5 285.2 247.01 34.67 185
Emotion Single 1.51G 0.49M 767.5 130.5 390.36 271.02 41.1 171.5
Multi 1.43G 2.14M 781 142 285.16 251.48 37.9 189.5
ChestX Single 0.64G 0.72M 690.5 126 210.45 237.45 29.57 165
Multi 1.06G 0.7M 670 130.5 261.14 236.96 30.14 183.5
ResNet18 - 3.67G 11.18M 1060 189 703.52 303.75 89.8 311
Multi-ResNet18 - 14.68G 55.9M 5816 356 4242.27 482.26 454.44 436
Table 8: On-device Performance.

### v2

@@
-XTransfer: Cross-Modality Model Transfer for Human Sensing with Few Data at the Edge
+XTransfer: Modality-Agnostic Few-Shot Model Transfer for Human Sensing at the Edge
 Abstract
-Deep learning for human sensing on edge systems offers significant opportunities for smart applications.
+Deep learning for human sensing on edge systems presents significant potential for smart applications.
 However, its training and development are hindered by the limited availability of sensor data and resource constraints of edge systems.
-Current methods that rely on transferring pre-trained models often encounter issues such as modality shift and high resource demands, resulting in substantial accuracy loss, resource overhead, and poor adaptability across different sensing applications.
-In this paper, we propose XTransfer, a first-of-its-kind method for resource-efficient, modality-agnostic model transfer.
-XTransfer freely leverages single or multiple pre-trained models and transfers knowledge across different modalities by (i) model repairing that safely repairs modality shift in pre-trained model layers with only few sensor data, and (ii) layer recombining that efficiently searches and recombines layers of interest from source models in a layer-wise manner to create compact models.
+While transferring pre-trained models to different sensing applications is promising, existing methods often require extensive sensor data and computational resources, resulting in high costs and poor adaptability in practice.
+In this paper, we propose XTransfer, a first-of-its-kind method enabling modality-agnostic, few-shot model transfer with resource-efficient design.
+XTransfer flexibly uses single or multiple pre-trained models and transfers knowledge across different modalities by (i) model repairing that safely mitigates modality shift by adapting pre-trained layers with only few sensor data, and (ii) layer recombining that efficiently searches and recombines layers of interest from source models in a layer-wise manner to create compact models.
 We benchmark various baselines across diverse human sensing datasets spanning different modalities.
-Comprehensive results demonstrate that XTransfer achieves state-of-the-art performance on human sensing tasks while significantly reducing the costs of sensor data collection, model training, and edge deployment.
+Comprehensive results demonstrate that XTransfer achieves state-of-the-art performance while significantly reducing the costs of sensor data collection, model training, and edge deployment.
 1 Introduction
-Human sensing refers to the process of capturing and interpreting data related to human activities, behaviors, and physiological states using various sensors (humansensing2023, ; Ystgaard2023, ).
-With the proliferation of edge devices equipped with sensors, human sensing plays a vital role in edge systems to understand contexts and enable smart applications, ranging from activity recognition to emotion or vital sign detection (Ahamed2018, ).
-Deep learning (DL) offers robust performance in interpreting sensor data, and its deployment on edge devices improves privacy and reduces bandwidth (Xiaoqiang2019, ; zhang2020, ; Yanjiao2020, ).
-However, existing DL solutions are often data- and resource-intensive for training and deployment (ondevicedl, ), posing significant challenges for data collection and edge deployment in human sensing (surveyiot, ; iotdatachallenge, ).
-Unlike other data modalities (e.g., text or image), collecting human sensing data for training can be costly and even impractical.
-This is uniquely due to the data that presents brittleness (i.e., noise sensitivity, low SNR), inconsistency (i.e., user variance, scenario changes), and heterogeneity (i.e., differences in hardware configuration) (Lane2015, ; Jeyakumar19, ; Teh2020, ).
-Collecting such data typically requires ethical approvals, as it involves sensitive information (e.g., location, motion patterns) that raises privacy concerns (Kroger2019, ; zhangipsn, ).
-Manual annotation of large volumes of sensor data introduces additional complexity and incurs high costs (Song2023fewshot, ; chen2020deep, ).
-Furthermore, human sensing spans various modalities (e.g., IMU, ultrasound, mmWave radar) and data types (e.g., spectrograms, Doppler profiles, time series), thereby adding complexity to model training, adaptation, and edge deployment (humansurvey, ; humansensing2023, ).
-Prior research has explored Few-shot Learning (FSL) (metasense, ; Wang2024, ), transfer learning (tlhar, ), and their combination (i.e., cross-domain FSL) (crossHAR, ; FewSense, ) to enable DL-based human sensing with few data.
-However, they either still require large-scale labeled datasets from the same singular modality to costly train source models from scratch (Song2023fewshot, ), or face challenges such as modality shift and limited adaptability when leveraging pre-trained models (Jaehoon2022, ; Ran2022, ; Sommelier2022, ), particularly in cross-modality settings.
-These limitations highlight the need for a scalable model transfer that can broadly adapt pre-trained models across modalities (e.g., vision or text) to new sensing modalities (e.g., radar or bio-signals) with few data.
-Recent advances in multi-modality learning (mmfi, ; chen2025xfi, ; chen2024probabilistic, ; madaan2024jointly, ; yang2024facilitating, ; baldenweg2024multimodal, ) with methods such as knowledge distillation (e.g., image-to-sensor transfer (zhao2018through, ; Xingzhe2022, ; Sevgi2020, ; Yu2023, )) and multi-modal model (e.g., CLIP-based models (girdhar2023imagebind, ; wu2023nextgpt, )), demonstrate the growing feasibility of cross-modality model transfer, driven by the increasing availability of high-quality pre-trained models from public repositories (e.g., PyTorch Hub (PyTorchhub, )), which opens new opportunities for real-world human sensing.
-However, due to substantial differences in data characteristics—such as shape, distribution, and feature representation—between source and target modalities (i.e., modality shift), transferred models often suffer from significantly degraded performance (i.e., negative transfer) (Sommelier2022, ; Zhang2023, ; Jaehoon2022, ).
-Our experiments reveal that state-of-the-art baselines (Tab.
-1), though effective on vision and NLP tasks, fail to reach oracle performance (metasense, ) and suffer from substantial overfitting when applied to human sensing tasks with few data, falling notably short of the goal, as shown in Fig.
-2(a).
+Human sensing refers to the process of capturing and interpreting data related to human activities, behaviors, and physiological states using various sensors (Zhang et al., 2023a; Ystgaard et al., 2023).
+With the proliferation of edge devices equipped with sensors, human sensing plays a vital role in edge systems to understand contexts and enable smart applications, ranging from activity recognition to emotion or vital sign detection (Ahamed & Farid, 2018).
+Deep learning (DL) offers robust performance in interpreting sensor data, and its deployment on edge devices improves privacy and reduces bandwidth (Zhang et al., 2020; Chen et al., 2020).
+However, existing DL solutions are often data- and resource-intensive for training and deployment (Dhar et al., 2021), posing significant challenges for data collection and edge deployment in human sensing (Sobin, 2020; Liu et al., 2019).
+Unlike other data modalities (e.g., vision or text), collecting human sensing data for training can be costly and even impractical.
+This is uniquely due to the data that presents brittleness (i.e., noise sensitivity, low SNR), inconsistency (i.e., user variance, scenario changes), and heterogeneity (i.e., differences in hardware configuration) (Lane & Georgiev, 2015; Jeyakumar et al., 2019; Teh et al., 2020).
+Collecting such data typically requires ethical approvals, as it involves sensitive information (e.g., location, motion patterns) that raises privacy concerns (Zhang et al., 2022).
+Manual annotation of large volumes of sensor data is labor-intensive and incurs high costs (Song et al., 2023).
+Importantly, human sensing spans various modalities (e.g., IMU, ultrasound, mmWave radar) and data types (e.g., spectrograms, Doppler profiles, time series), further amplifying the costs of applying DL solutions and naturally leading to cross-modality scenarios (Xiao et al., 2022; Zhang et al., 2023a).
+Prior research has explored Few-shot Learning (FSL) (Wang et al., 2024), transfer learning (Dhekane & Ploetz, 2024), and their combination (i.e., cross-domain FSL) (Thukral et al., 2025; Yin et al., 2024) to notably reduce data collection costs by adapting pre-trained models to sensing applications.
+However, they still rely on large-scale labeled datasets from the same modality to costly train source models from scratch (Song et al., 2023), or face challenges such as modality shift and poor adaptability when leveraging pre-trained models (Oh et al., 2022; Tao et al., 2022; Guo et al., 2022), particularly in cross-modality settings.
+Recent advances in multi-modal learning (Yang et al., 2023; Chen & Yang, 2025; mengxi Chen et al., 2024; Madaan et al., 2024; Yang et al., 2024; Baldenweg et al., 2024) (e.g., CLIP-based models (Girdhar et al., 2023; Wu et al., 2023; Moon et al., 2023)) and cross-modal learning (e.g., image-to-sensor distillation (Zhao et al., 2018; Song et al., 2022; Gurbuz et al., 2020; Zhang et al., 2023c) or alignment-based transfer (Kamboj et al., 2025; Chen et al., 2024)), demonstrate the feasibility of knowledge transfer across modalities, but they remain modality-specific (i.e., relying on shared semantic feature spaces), and typically require large-scale paired data with resource-intensive training and deployment, leading to high costs and making them impractical for few-shot adaptation to support human sensing at the edge.
+These limitations motivate the need for a scalable and adaptable model transfer that can broadly adapt pre-trained models across modalities (e.g., vision or text) to new sensing modalities (e.g., radar or bio-signals) with few data.
+In practice, however, addressing this need is uniquely challenging.
+Due to substantial differences in data characteristics—such as shape, distribution, and feature representation—between source and target modalities (i.e., modality shift), transferred models often suffer from significantly degraded performance (i.e., negative transfer) (Guo et al., 2022; Oh et al., 2022).
+Our experiments reveal that state-of-the-art (SOTA) baselines (Table 1) fail to reach oracle performance (Gong et al., 2019) and suffer from substantial overfitting when applied to human sensing tasks with few data, falling notably short of the goal, as shown in Figure 2(a).
 These highlight the underlying challenge of latent feature misalignment across modalities.
-Also, these methods often neglect resource constraints, rendering them impractical for edge deployment, especially in advanced multi-source settings (Sicheng2020, ; Xiangyumfs, ; NEURIPS20196048ff4e, ).
-While methods such as neural architecture search (NAS) (searchinsights, ; AdaptiveNet2023, ; LegoDNN2021, ), pruning (Anwar2017, ; frankle2018the, ; pers2020composition, ), or quantization (Polino2018ModelCV, ), aim to reduce resource demands, they often fail to maintain performance under modality shift and few data.
-As shown in Fig.
-2(c), our experiments using structural pruning (frankle2018the, ; NEURIPS2022, ) demonstrate significant model accuracy loss (50.5% on average).
-These challenges point to a critical gap in developing adaptable and scalable methods.
-To overcome the challenges, we propose XTransfer, a first-of-its-kind cross(X)-modality model Transfer method that freely leverages public source models and transfers knowledge across different modalities (e.g., image, text, audio, or sensing) to human sensing tasks, significantly reducing the need for extensive sensor data collection and training from scratch.
-It also offers resource-efficient designs to optimize edge deployment, paving the way for scalable real-world applications.
-XTransfer uniquely repairs, searches and recombines layers of interest from pre-trained source models under cross-modality FSL settings (see App.
-A), resulting in compact and adaptable models.
-To mitigate modality shift and prevent overfitting, it enables model repairing achieved by a novel Splice-Repair-Removal (SRR) pipeline (c.f.
-Sec.
-4).
-This pipeline transforms a few sensor data to broadly adapt pre-trained model layers by repairing layer-wise misalignment and harmonizing latent feature distributions across modalities.
-XTransfer also features layer recombining using a novel Layer-Wise Search (LWS) control mechanism (c.f.
-Sec.
-5) to identify and recombine layers of interest by conducting a layer-wise search.
-Building on the insight that not all repaired layers contribute, this mechanism discards such layers and recombines only useful ones to optimize model compactness using principles from NAS (searchinsights, ).
-To further accelerate the search when dealing with multiple sources, it integrates a pre-search check strategy to efficiently avoid unnecessary repairing.
-Comprehensive evaluations (c.f.
-Sec.
-6) are conducted on 8 public datasets along with their corresponding pre-trained source models, spanning diverse modalities (e.g., image, text, audio, and sensing), 4 collected target sensing datasets in real-world settings, 2 public sensing datasets and 1 standard image dataset as targets (Tab.
-2) on 3 commercial edge devices.
-The results show that XTransfer achieves a significantly enhanced accuracy-to-resource (ATR) ratio (c.f.
-Sec.
-6.1), outperforming single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively, while using few data with broad adaptability.
-XTransfer also streamlines the layer search process, reducing search time by 2.1 to 4 times.
-The main contributions of this work are:
-  • A first-of-its-kind method, XTransfer, that enables resource-efficient, modality-agnostic model transfer for human sensing with few data at the edge, as shown in Fig. 3.
-  • A novel SRR pipeline that effectively mitigates modality shift and prevents overfitting, harmonizes latent feature distributions across modalities.
-  • A novel LWS control mechanism that efficiently searches layers of interest from multiple sources in a layer-wise manner, creating compact models.
-  • Extensive experimental evaluations, benchmarking state-of-the-art baselines based on various source and target datasets, validating competitive scalable performance of XTransfer.
+Also, these methods often neglect resource constraints, rendering them impractical for edge deployment, especially in advanced multi-source settings (Zhao et al., 2020; Yue et al., 2021a; Lee et al., 2019).
+While methods such as neural architecture search (NAS) (Wen et al., 2023; Han et al., 2021), pruning (Frankle & Carbin, 2019; Shen et al., 2022), or quantization (Polino et al., 2018), aim to reduce resource demands, they often fail to maintain performance under modality shift and few data.
+As shown in Figure 2(c), our experiments using pruning (Shen et al., 2022) demonstrate significant model accuracy loss by 50.5% on average.
+Figure 1: Overview.
+XTransfer transfers source models across modalities with few sensor data through model repairing (SRR pipeline) and layer recombining (LWS control).
+LWS control first segments source models into layers and operates layer-wise search across pools.
+At each pool, the pre-search check decides which layers need repairing, then SRR pipeline repairs them and LWS control selects layers of interest.
+These layers are incrementally recombined during the search, resulting in a compact model for enabling human sensing at the edge.
+Subfigures (a)–(c) illustrate the feature space evolution before and after repairing.
+Driven by the increasing availability of high-quality pre-trained models from public repositories (e.g., PyTorch Hub (Hub, 2022)), we propose XTransfer, a first-of-its-kind method enabling modality-agnostic, few-shot model transfer.
+It flexibly leverages public pre-trained models and transfers knowledge across different modalities (e.g., image, text, audio, or sensing) to human sensing tasks using only few sensor data, significantly reducing the costs of large-scale data collection and training from scratch.
+Its resource-efficient design further optimizes edge deployment, enabling both scalability and adaptability to diverse human sensing applications.
+XTransfer uniquely repairs, searches and recombines layers of interest from pre-trained models under cross-modality FSL settings (see Appendix A), resulting in compact models.
+To mitigate modality shift and prevent overfitting, it enables model repairing using a novel Splice-Repair-Removal (SRR) pipeline (see Section 4).
+SRR pipeline transforms a few sensor data to broadly adapt pre-trained layers by repairing layer-wise misalignment and harmonizing latent feature distributions across modalities.
+XTransfer also features layer recombining using a novel Layer-Wise Search (LWS) control mechanism (see Section 5) to identify and recombine layers of interest by conducting a layer-wise search.
+Building on the insight that not all repaired layers contribute, LWS control discards such layers and recombines only useful ones to optimize model compactness using principles from NAS (White et al., 2023).
+To further accelerate the search in multi-source settings, a pre-search check strategy is designed to efficiently avoid unnecessary repairing.
+The main contributions are summarized as: 1) a first-of-its-kind method, XTransfer, that enables modality-agnostic, few-shot model transfer with resource-efficient design for human sensing on edge systems, as shown in Figure 1; 2) a novel SRR pipeline that effectively mitigates modality shift and prevents overfitting, harmonizes latent feature distributions across modalities; 3) a novel LWS control mechanism that efficiently searches layers of interest from multiple sources in a layer-wise manner to create compact models; and 4) extensive experimental evaluations (see Section 6) under cross-modality FSL settings, benchmarking SOTA baselines (Table 1) on 8 public datasets with corresponding pre-trained source models across diverse modalities (e.g., image, text, audio, and sensing) and 7 target datasets (Table 2) using 3 commercial edge devices.
+The results show that XTransfer achieves a significantly enhanced accuracy-to-resource (ATR) ratio (see Section 6.1), outperforming single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively, while reducing layer search time by 2.1 to 4 times and substantially lowering the costs of data collection, model training, and edge deployment, demonstrating broad adaptability.
 Figure 2: Preliminary study.
-(a) reveal baseline performance gap  [^2] We test a pre-trained ResNet18 on miniImageNet and HHAR dataset as target under a 5-shot setting (see App.
-A).(overfitting, ).
-(b) shows the average similarity and FSL difficulty across all sensing datasets (Tab.
-2) to each source modality (e.g., Image, Text, Sensing) (c.f.
-Sec.
-3).
+(a) reveal baseline performance gap  [^2] We test a pre-trained ResNet18 on miniImageNet as source and HHAR dataset as target under a 5-shot setting (see Appendix A)..
+(b) shows the average similarity and FSL difficulty across all sensing datasets to source modalities (e.g., Image, Text, Sensing) using default reshaping (see Appendix B.1).
 2 distinct areas represent similarity levels (A–hard, B–normal).
-Key findings: 1) compared to CUB, similarity levels across modalities are notably low, e.g., Text and Sensing fall into Area A, indicating a significant modality shift; 2) compared to Image-Unpair (i.e., no class pairing) in Area A, Image surprisingly falls into Area B, indicating that pairing classes may enhance cross-modality similarity (motivating our techniques in Sec.
-4.1); 3) compared to Text and Sensing, Image exhibits more stable standard deviations and lower FSL difficulty, suggesting better potential for broad adaptability (c.f.
-Sec.
-4).
-(c) shows model optimisation bottlenecks.
+Key findings: 1) compared to CUB, similarity levels across modalities are notably low, e.g., Text and Sensing fall into Area A, indicating a significant modality shift; 2) compared to Image-Unpair (i.e., no class pairing) in Area A, Image surprisingly falls into Area B, indicating that pairing classes may enhance cross-modality similarity (motivating our techniques in Section 4.1); 3) compared to Text and Sensing, Image exhibits more stable standard deviations and lower FSL difficulty, suggesting better potential for broad adaptability.
+(c) shows model pruning bottlenecks.
 2 Related Work
@@
 Recent works in FSL and cross-domain FSL have proposed single- and multi-source methods.
-Single-source methods align latent features between source models and target data via distance-based loss functions (Zhao2021, ; Yue2021Prototypical, ), whereas multi-source methods leverage the Wisdom of the Crowd principle through source distillation (Sicheng2020, ; Xiangyumfs, ) and maximal correlation analysis (NEURIPS20196048ff4e, ).
-Existing attempts on human sensing tasks (metasense, ; Wang2024, ; tlhar, ; crossHAR, ; FewSense, ) remain constrained by a strong reliance on learning in the same modality, resulting in costly source data collection and model training with limited adaptability.
-Recent efforts have also explored fine-tuning large language models (e.g., GPT2 (zhou2023one, )) for time-series analysis.
-However, testing on benchmarks (Boosting2021, ; Jaehoon2022, ; chen2018a, ), our results reveal that they still suffer from severe overfitting under modality shift and few data.
-Also, none of them explicitly considers resource constraints of edge systems.
-In contrast, XTransfer enables model repairing to safely mitigate modality shifts using only few sensor data, enabling modality-agnostic model transfer with broad adaptability.
+Single-source methods align latent features between source models and target data via distance-based objectives (Zhao et al., 2021; Yue et al., 2021b), whereas multi-source methods leverage the Wisdom of the Crowd principle through unsupervised source distillation (Zhao et al., 2020; Yue et al., 2021a) or maximal correlation analysis (Lee et al., 2019).
+Recent efforts have also explored fine-tuning LLMs (e.g., GPT2 (Zhou et al., 2023)) for time-series analysis.
+However, testing on benchmarks (Liang et al., 2021; Oh et al., 2022), our results reveal that they suffer from severe overfitting under cross-modality FSL settings.
+LLM-based data augmentation (Leng et al., 2024; 2023) helps reduce data collection costs, but remains modality-specific and orthogonal to our focus.
+Existing attempts on human sensing tasks (Gong et al., 2019; Wang et al., 2024; Thukral et al., 2025; Yin et al., 2024) remain constrained by a strong reliance on learning in the same modality, resulting in costly source data collection and training with limited adaptability.
+In contrast, XTransfer enables layer-wise model repairing to safely mitigate modality shifts using only few sensor data, achieving significant cost reduction with broad adaptability.
 Resource constraints optimization.
-Model compression techniques such as model pruning (Anwar2017, ; frankle2018the, ; pers2020composition, ), quantization using low-precision data types (Polino2018ModelCV, ), model merging by fine-tuning shared layers (Gemel2023, ), one-size-fits-all model (cai2020once, ), and NAS techniques such as model structure optimization (AdaptiveNet2023, ; LegoDNN2021, ) or search using few data (neuralFT, ; xu2022fewshot, ), have shown success in reducing model resource overhead.
+Model compression techniques such as model pruning (Frankle & Carbin, 2019; Shen et al., 2022), quantization using low-precision data types (Polino et al., 2018), model merging by fine-tuning shared layers (Padmanabhan et al., 2023), one-size-fits-all model (Cai et al., 2020), and NAS techniques such as model structure optimization (Wen et al., 2023; Han et al., 2021) or search using few data (Eustratiadis et al., 2024; Xu et al., 2022), have shown success in reducing model resource overhead.
 However, they strongly rely on either sufficient target datasets or minimal modality shifts to maintain model performance.
-XTransfer uniquely applies layer recombining to optimize model structures by efficiently searching and recombining layers of interest from pre-trained source models.
+XTransfer uniquely applies layer recombining to optimize model structures while strengthening layer-wise dependence, by efficiently searching and recombining layers of interest from pre-trained source models at scale.
 It not only creates compact models for streamlined edge deployment but also further enhances model repairing.
 3 Preliminary
-Insights in cross-modality.
-We apply a standard cross-domain FSL benchmarking (Jaehoon2022, ) to gain insights.
-It uses the Earth Mover’s Distance (EMD) to measure similarity between source models from 3 modalities (i.e., Image, Text, Sensing) and targets, including 5 human sensing target datasets and 2 well-known image datasets (Tab.
-2).
-It also assesses FSL difficulty by fine-tuning source models on targets.
-Since source models involve a large number of classes  [^3] The number of classes in the source model should be greater than that in the target.
-(e.g., 100 classes in miniImageNet), we propose a class pairing strategy that pairs source and target classes exhibiting similar latent feature distributions across modalities.
-The key insights are shown in Fig.
-2(b).
-Layer-wise analysis.
-We perform a layer-wise analysis to gain insights, focusing on accuracy changes and latent feature distributions at each layer.
-Since effective training relies on learning discriminative features (islam2021broad, ), we benchmark latent feature distribution clustering and use Mean Magnitude of Channels (MMC) (i.e., channel importance) (pmlrv162luo22c, ) of each layer to indicate layer-wise performance.
-We define a single layer or a dependent layer block (i.e., L-Blocks in Tab.
-3)—where layers are not structurally separable—as the default segmentation unit (i.e., L-unit).
-Fig.
-2(a) shows that negative transfer (Zhang2023, ) disrupts accuracy convergence across layers (i.e., layer-wise dependence), causing layer-wise misalignment and accuracy drops.
-These misalignments correlate with MMC shifts, where greater shifts correspond to larger accuracy drops shown in Fig.
-2(b), indicating inconsistent latent feature interpretation during fine-tuning.
-To track MMC shifts, we essentially employ the Silhouette score (S-score) (shahapure2020cluster, ), which captures inter- and intra-class distances and reflects the discriminability of layer-wise latent features.
-Challenges.
-We aim to minimize layer-wise MMC shifts, but several technical challenges arise.
-First, the pre-trained feature space at each layer is highly complex due to the multidimensional nature of MMC, and this complexity increases with layer depth.
-Such complexity poses challenges for "repairing" latent feature distributions across modalities using few sensor data.
-Second, strong layer-wise dependencies, formed during source model training, become increasingly rigid in deeper layers, posing challenges for standard techniques (e.g., N-Pair loss (Sohn2016, ), Triplet loss (Schroff2015, )) that may fail to preserve these dependencies.
-Third, Fig.
-2(c) shows that S-scores may stagnate or even drop at certain layers (i.e., ineffective layers), impairing overall convergence, indicating that not all repaired layers contribute positively.
-This motivates our LWS control design, which selectively recombines effective layers (i.e., layers of interest) while discarding others.
-However, such selection must account for on-device resource constraints to optimize the layer growth.
-The search space also expands rapidly with the number of layers or source models, further complicating efficient search.
 Figure 3: Design insights.
@@
 (c) After repairing, layer S-score improves, but stagnation occurs at certain layers.
+Insights in cross-modality FSL.
+We apply a standard cross-domain FSL benchmarking (Oh et al., 2022) to gain insights.
+It uses Earth Mover’s Distance (EMD) to measure similarity between source models from 3 modalities (i.e., Image, Text, Sensing) and targets including 5 human sensing and 2 well-known image datasets (Table 2).
+It also assesses FSL difficulty by fine-tuning source models on targets.
+Since source models involve a large number of classes  [^3] The number of classes in the source model should be greater than that in the target.
+(e.g., 100 classes in miniImageNet), we propose a class pairing strategy that pairs source and target classes exhibiting similar latent feature distributions across modalities.
+The key insights are shown in Figure 2(b).
+Layer-wise analysis.
+We perform a layer-wise analysis to gain insights, focusing on accuracy changes and latent feature distributions at each layer.
+Since effective training relies on learning discriminative features (Islam et al., 2021), we benchmark latent feature distribution clustering and use Mean Magnitude of Channels (MMC) (Luo et al., 2022) of each layer as the default metric to quantify layer-wise performance, as it offers a lightweight measure for channel importance.
+We define a single layer or a dependent layer block (i.e., L-Blocks in Table 3)—where layers are not structurally separable—as the default segmentation unit (i.e., L-unit).
+Figure 3(a) shows that negative transfer (Zhang et al., 2023b) disrupts accuracy convergence across layers (i.e., layer-wise dependence), causing layer-wise misalignment and accuracy drops.
+These misalignments correlate with MMC shifts, where greater shifts correspond to larger accuracy drops shown in Figure 3(b), indicating inconsistent latent feature interpretation during fine-tuning.
+To track MMC shifts, we essentially employ the Silhouette score (S-score) (Shahapure & Nicholas, 2020), which captures inter- and intra-class distances and reflects the discriminability of layer-wise latent features.
+Challenges.
+We aim to minimize layer-wise MMC shifts under cross-modality FSL settings, but several technical challenges arise.
+First, it is hindered by structural and semantic gaps between sources and target data, where classes and semantics often differ.
+Second, the pre-trained feature space at each layer is highly complex due to the multidimensional and noisy nature of MMC across source models from different modalities.
+This complexity increases with layer depth, as the strong layer-wise dependence formed during source model training becomes increasingly rigid, making it difficult to repair layer-wise misalignment using only few sensor data.
+Third, Figure 3(c) shows that S-scores may stagnate or even drop at certain layers (i.e., ineffective layers), impairing overall convergence.
+It indicates that not all repaired layers contribute positively, motivating our LWS control design to selectively recombine layers of interest while discarding others.
+The selection must also account for on-device resource constraints to control layer growth, while the rapidly expanding search space with increasing layers and source models further challenges the layer search.
 4 Model Repairing
-Enabled by SRR pipeline, this process safely repairs layer-wise misalignment by minimizing MMC shifts, effectively restoring layer performance from source models using few sensing inputs in cross-modality settings.
-SRR pipeline involves three key stages shown in Fig.
-3.
-In the Splice stage, input shapes between source models and sensor data must be aligned well, as default reshaping (c.f.
-App.
-B.1)—using fixed down/up-sampling—cannot be fine-tuned and may lose important features across channel, height, and width, especially when layer shapes vary during layer-wise search.
-To enhance, we design a compact connector placed between sensing inputs and layers, consisting of a Pre-header (adaptive convolutional layer), Resizer, and one encoder-decoder pair to bridge cross-modality structures.
-In the next Repair stage, each connector is fine-tuned to generate adaptive channels, preserve key sensing features, and minimize MMC shift via the generative transfer module.
-In the following Removal stage, channel removal is designed to further enhance layer-wise performance while optimizing resource usage.
-For more design details, see Appendix B.2.
-Figure 3: XTransfer overview.
-LWS control segments source models into layers and uses the pre-search check to decide if repairing is needed.
-SRR pipeline then fine-tunes connectors to repair selected layers.
-Finally, LWS control selects and recombines layers of interest into a compact model.
+Enabled by SRR pipeline, this process safely repairs layer-wise misalignment by minimizing MMC shifts, effectively restoring layer performance from source models using few sensing inputs under cross-modality FSL settings.
+Three stages are designed, as shown in Figure 1.
+In the Splice stage, default reshaping (see Appendix B.1)—using fixed down/up-sampling—cannot be fine-tuned and may lose important features, especially when layer shapes vary during layer-wise search.
+We propose a compact connector placed between sensing inputs and layers, consisting of a Pre-header (adaptive convolutional layer), Resizer, and one encoder-decoder pair to bridge cross-modality structural gaps.
+In the Repair stage, each connector is fine-tuned by the generative transfer module to generate adaptive channels, preserve key sensing features, and minimize MMC shifts.
+Finally, the Removal stage applies channel removal (see Appendix B.2) to further enhance layer-wise performance.
 4.1 Repairing in complex feature space
-Since the original latent features from source models are well-trained for high discriminability (c.f.
-Sec.
-3), our intuition is to safely fine-tune the misinterpreted latent features of few sensor data by aligning with the original features (i.e., anchors) to minimize MMC shifts.
-We hence propose anchor-based cross-modality supervision alignment, which aligns source MMCs as anchors (i.e., anchor MMCs) with sensing MMCs.
-To handle complex feature spaces, our insight is to reduce MMC dimensionality by preserving key channels and excluding irrelevant ones.
-Reduced-dimensional feature space.
-We project anchor MMCs into a lower-dimensional feature space using principal component analysis (PCA) (VALPOLA2015143, ) (i.e., anchor PCA space).
-PCA enables a lower-dimensional subspace (i.e., adjusted by the number of PCA components, with 2 components used by default) where the projected data exhibits the highest possible variance (i.e., highly clustered).
-The optimized component weights (i.e., projection coefficients) highlight the most important channels that contribute to the performance of each layer.
-Given high accuracy achieved in each original layer, the projected anchor MMCs show high clustering performance, e.g., the centroid of the distribution of each class (i.e., Anchor-center) clustered tightly, as shown in Fig.
-3(a).
-Once the anchor PCA space is initialized for each layer, we project sensing MMCs into its space using the same parameters.
+Since the original latent features from source models are well-trained for high discriminability, our intuition is to safely fine-tune few sensor data by aligning the misinterpreted latent features with the original features (i.e., anchors) to minimize MMC shifts.
+We hence propose cross-modality anchor-based alignment, which aligns source MMCs as anchors (i.e., anchor MMCs) with sensing MMCs.
+To handle complex feature spaces, our insight is to reduce MMC dimensionality by preserving key channels and excluding irrelevant or noisy ones.
+Reduced-orthogonal feature space.
+We project anchor MMCs into a lower-dimensional, orthogonal feature space using principal component analysis (PCA) (Valpola, 2015) (i.e., anchor PCA space).
+Unlike the raw MMC feature space, where channel dimensions are often noisy and correlated, PCA enforces orthogonality across components, removing redundancy and suppressing noise.
+This enables a compact subspace (i.e., defaulting to 2 components) where the projected data captures maximum variance and forms highly clustered distributions.
+The optimized component weights (i.e., projection coefficients) highlight the most important channels contributing to layer performance.
+Given high discriminability of the original layers, the projected anchor MMCs preserve high clustering performance, with class distributions tightly centered (i.e., Anchor-center), as shown in Figure 1(a).
+Once initialized for each layer, this anchor PCA space is reused to project sensing MMCs.
 Feature space alignment.
-After projecting sensing MMCs, we observe that the sample distribution for each class, along with their centroid (i.e., Sensing-center), appears disorganized compared to the well-clustered Anchor-center at each layer, as shown in Fig.
-3(a).
+After projecting sensing MMCs, we observe that the sample distribution for each class, along with their centroid (i.e., Sensing-center), appears disorganized compared to the well-clustered Anchor-center at each layer, as shown in Figure 1(a).
 Specifically, the scale and orientation of Sensing-center are notably different from those of Anchor-center, caused by MMC shifts.
@@
 $i$ denotes the index of pre-trained models, and $j$ denotes the layer index in each model $i$.
-Next, to align the orientation, we design a rotation alignment function that minimizes the cosine angle (nguyen2010cosine, ) by fine-tuning a multi-dimensional rotation matrix $[M_{rot}]$ (where the dimensions equal to the number of PCA components) as below:
-$\footnotesize\begin{split}\arg\min_{[M_{rot}]}Cos(Cen(Pro(fs_{ij})),Cen(Pro(ft_{ij})*S)\cdot[M_{rot}])\end{split}$ (1)
-where $Cos(\cdot)$ denotes the cosine similarity function, and $Cen(\cdot)$ calculates the centroid of the projected distribution. Fig. 3(b) shows the aligned results.
+Next, to align the orientation, we design a rotation alignment function that minimizes the cosine angle (Nguyen & Bai, 2010) by fine-tuning a multi-dimensional rotation matrix $M_{rot}]$ with dimensions matching the number of PCA components:
+$\begin{split}\arg\min_{[M_{rot}]}Cos(Cen(Pro(fs_{ij})),Cen(Pro(ft_{ij})*S)\cdot[M_{rot}])\end{split}$ (1)
+where $Cos(\cdot)$ denotes the cosine similarity function, and $Cen(\cdot)$ calculates the centroid of the projected distribution. Figure 1(b) shows the aligned results.
 Anchor class pairing.
-Building on key finding that class pairing boosts latent feature alignment across modalities (c.f.
-Sec.
-3), we propose selecting source classes with the highest S-scores as anchor classes to pair with sensing classes to facilitate the supervision alignment.
+Building on the key finding that class pairing boosts latent feature alignment across modalities (see Section 3), we propose selecting source classes with the highest S-scores as anchor classes to pair with sensing classes to facilitate the anchor-based alignment.
 We minimize pairing shift (i.e., the sum distance across each pair) via a linear sum assignment solved using the standard Hungarian algorithm, outputting the optimal pairing set $pairST$.
@@
 Generative transfer module.
-To effectively restore layer-wise dependence, we design a generative transfer module that enhances the supervision alignment through generative learning.
-It consists of a generator (i.e., connector) and a discriminator—each frozen pre-trained layer paired with its anchor PCA space (Fig.
-3).
+To effectively restore layer-wise dependence, we design a generative transfer module that enhances the anchor-based alignment through generative learning.
+It consists of a generator (i.e., connector) and a discriminator—each frozen pre-trained layer paired with its anchor PCA space (Figure 1).
 Freezing preserves stable anchors and prevents overfitting.
@@
 Building on the generative transfer module, we propose an anchor-based repair loss function ($loss_{srr}$) to harmonize latent feature distributions across modalities.
-Given each layer’s anchor PCA space, our objective is twofold: 1) minimize the Euclidean distance (ED) ($D(\cdot)$) between the centroid of projected anchor MMC distribution and that of the projected sensing MMC distribution for each class pair $(c_{s},c_{t})\in pairST$ (i.e., positive loss), and 2) maximize the distance between sensing MMC samples with different labels $(c_{t}^{-})$ (i.e., negative loss).
-Since the projected anchor MMCs are highly clustered, it can provide the maximum margin ($M_{max}$) between anchor clusters.
+Given each layer’s anchor PCA space, our objective is twofold: 1) minimize the distance ($D(\cdot)$) between the centroid of projected anchor MMC distribution and that of the projected sensing MMC distribution for each class pair $(c_{s},c_{t})\in pairST$ (i.e., positive loss), and 2) maximize the distance between sensing MMC samples with different labels $(c_{t}^{-})$ (i.e., negative loss).
+Since the projected anchor MMCs are highly clustered, they can provide the maximum margin ($M_{max}$) between anchor clusters.
 To further improve the negative loss, we update the function by calculating the anchor-based margin $margin^{c}=InterD(Pro(fs^{c}_{ij}))-IntraD(Pro(fs^{c}_{ij}))$, where the $M_{max}$ is selected, and $IntraD(\cdot)$ measures the intra-class distance.
 Hence, our loss function is formulated as:
-$\footnotesize\begin{split}loss_{srr}=&\frac{1}{N}\sum^{N}_{(c_{s},c_{t})\in pairST}D(Cen(Pro(fs^{c_{s}}_{ij})),Cen(Pro(ft^{c_{t}}_{ij})))\\ +&\frac{1}{K}\sum^{K}_{(c_{t},c_{tk}^{-})}ReLU(M_{max}-D(Pro(ft^{c_{t}}_{ij}),Pro(ft^{c_{tk}^{-}}_{ij})))\end{split}$ (2)
+$\begin{split}loss_{srr}=&\frac{1}{N}\sum^{N}_{(c_{s},c_{t})\in pairST}D(Cen(Pro(fs^{c_{s}}_{ij})),Cen(Pro(ft^{c_{t}}_{ij})))\\ +&\frac{1}{K}\sum^{K}_{(c_{t},c_{tk}^{-})}ReLU(M_{max}-D(Pro(ft^{c_{t}}_{ij}),Pro(ft^{c_{tk}^{-}}_{ij})))\end{split}$ (2)
 where $K$ denotes the number of negative samples, and $N$ denotes the number of classes.
 The $ReLU(\cdot)$ is used to trigger the negative loss in the condition where the distance between the samples is less than $M_{max}$.
-Fig.
-3(c) shows the effectiveness of model repairing.
+Figure 1(c) shows the effectiveness of model repairing.
 5 Layer Recombining
-Working in conjunction with SRR pipeline for repairing each layer, XTransfer leverages layer recombining enabled by LWS control to: 1) perform the pre-search check that accelerates the search process, and 2) efficiently select and recombine layers of interest to form a compact model.
+Working in conjunction with SRR pipeline for repairing each layer, XTransfer leverages layer recombining enabled by LWS control to: 1) perform the pre-search check that accelerates the search process, and 2) effectively select and recombine layers of interest to form a compact model.
 5.1 Search control design
-Following NAS principles (searchinsights, ), poor design choices (e.g., search space, actions, strategy) can disrupt the balance between model performance and resource efficiency.
+Following NAS principles (White et al., 2023), poor design choices (e.g., search space, actions, strategy) can disrupt the balance between model performance and resource efficiency.
 As layers of interest grow, resource overhead can increase exponentially, making dynamic search control essential under on-device resource constraints.
-To efficiently manage, we define the search space as all layers from source models, structured into search pools containing layer candidates (i.e., repaired layers), each with width $I$ (number of models) and depth $J$ (number of L-units), forming a window of size $I\times J$.
+To effectively manage, we define the search space as all layers from source models, structured into search pools containing layer candidates (i.e., repaired layers), each with width $I$ (number of models) and depth $J$ (number of L-units), forming a window of size $I\times J$.
 Search actions (init, continue, skip, cross) are defined to guide layer recombination within each designated search pool.
@@
 Search strategy under resource constraints.
-To effectively control search actions within each search pool, we propose a layer value function based on the S-score $Score(\cdot)$ to evaluate both individual repaired layer performance and the convergence between adjacent layers after each SRR process $SRR(\cdot)$.
+To effectively control search actions within each search pool, we propose a layer value function based on the S-score $Score(\cdot)$ to evaluate each repaired layer performance and the convergence between adjacent layers after each SRR process $SRR(\cdot)$.
 It is formulated as $V_{ij}=Score(SRR(Pro(ft_{ij}),MT_{ij},L_{ij}),Y_{t}))$, where $MT_{ij}$ denotes the connector and $L_{ij}$ represents each repaired layer.
-Under resource constraints, only the top-performing layer in each pool is selected.
-LWS control then checks whether its S-score exceeds that of the previous layer.
+Under resource constraints, only the highest-valued repaired layer in each pool is selected.
+LWS control then checks whether its S-score exceeds that of the previous layer to maintain layer-wise dependence.
 If successful, the layer $L_{ij}$ is identified as the layer of interest and recombined, and the next pool window begins at layer index $j+1$.
@@
 Each layer’s adjusted overhead, $R(n)_{ij}$, is computed as $R(n)_{ij}=(Res(L_{ij})*RC(n))/Max(Res(P))$, where $Res(\cdot)$ calculates actual cost (e.g., FLOPs), and $Max(Res(P))$ is the peak cost in current pool $P$.
-The final resource-constrained layer value function becomes $VR_{ij}=V_{ij}\mathbin{/}R(n)_{ij}$.
-The objective is then defined as $\arg\max_{L_{ij}\in P}VR_{ij},\,s.t.\sum_{0}^{n}Res(L_{ij})\leq R_{Max}$, where $R_{Max}$ is the total allowed resource budget.
+The resource-constrained layer value function becomes $VR_{ij}=V_{ij}\mathbin{/}R(n)_{ij}$.
+The objective is defined as $\arg\max_{L_{ij}\in P}VR_{ij},\,s.t.\sum_{0}^{n}Res(L_{ij})\leq R_{Max}$, where $R_{Max}$ is device-specific resource budget.
 5.2 Efficient search
@@
 Based on this, our intuition is to estimate repaired S-score (i.e., repair rate) of each layer based on Before and Anchor states to decide whether to bypass unnecessary repairs.
-Experiments show the actual repair rate grows exponentially with the recombined layer index (see App.
-B.4).
+Experiments show the actual repair rate grows exponentially with the recombined layer index (see Appendix B.4).
 We hence design a repair rate growth model formulated as $rate_{n}^{est}=exp(a*n)+b,n\in[1,L]$, where $a$ and $b$ are optimized via non-linear regression to minimize MSE to fit $rate_{n}^{est}$, with the objective function defined as $\arg\min_{a,b}\frac{1}{n^{{}^{\prime}}}\sum_{n=0}^{n^{{}^{\prime}}}(rate_{n}^{est}-rate_{n})^{2}$, where $n^{{}^{\prime}}$ is the number of recombined layers.
@@
 Due to the limited updates, the estimation may be inaccurate.
-To mitigate, we also develop a dynamic search range mechanism that dynamically adjusts the search range in each pool (details in App.
-B.5).
+To mitigate, we also develop a dynamic search range mechanism that dynamically adjusts the search range in each pool (details in Appendix B.5).
 6 Experiment
-6.1 Experiment setup
-Table 1: State-of-the-art baselines.
+Table 1: State-of-the-art (SOTA) baselines.
   Baselines Group Description
-  TL (TransferabilityID, ) Single-source A standard transfer learning method by freezing all layers
-FT (Jaehoon2022, ) Single-source A standard fine-tuning method by updating all layers
-ProtoNet (Snell2017, ) Single-source A distance-based FSL method
-MAML (Finn2017, ) Single-source A learning-based FSL method
-DAPN (Zhao2021, ) Single-source A cross-domain FSL method using adversarial ProtoNet
-GPT2 (zhou2023one, ) Single-(LLM) A large language model based method using fine-tuning
-MDDA (Sicheng2020, ) Multi-source A cross-domain FSL method using knowledge distillation
-MCW (NEURIPS20196048ff4e, ) Multi-source A cross-domain FSL method using maximal correlation analysis
-MetaSense (metasense, ) Oracle-(GT) An advanced FSL method based on MAML with sufficient target data
+  TL (Jiang et al., 2022) Single-source A standard transfer learning method by freezing all layers
+FT (Oh et al., 2022) Single-source A standard fine-tuning method by updating all layers
+ProtoNet (Chen et al., 2019) Single-source A distance-based FSL method
+MAML (Chen et al., 2019) Single-source A learning-based FSL method
+DAPN (Zhao et al., 2021) Single-source A cross-domain FSL method using adversarial ProtoNet
+SemiCMT (Chen et al., 2024) Single-source A cross-modal method using contrastive alignment
+GPT2 (Zhou et al., 2023) Single-(LLM) A large language model based method using fine-tuning
+MDDA (Zhao et al., 2020) Multi-source A cross-domain FSL method using knowledge distillation
+MCW (Lee et al., 2019) Multi-source A cross-domain FSL method using maximal correlation analysis
+MetaSense (Gong et al., 2019) Oracle-(GT) An advanced FSL method based on MAML with sufficient target data
 Table 2: Source and target datasets specifications.
@@
   Source
-miniImageNet (miniimagenet, ) Image - 100 - 3x84x84 Public
-miniDomainNet (domainnet, ) Image - 126 - 3x224x224 Public
-Office-31 (office31, ) Image - 31 - 3x224x224 Public
-Office-Home (officehome, ) Image - 65 - 3x224x224 Public
-Caltech-101 (caltech101, ) Image - 101 - 3x224x224 Public
-Newsgroup (text2021, ) Text - 20 - 100x1x1000 Public
-VoxCeleb (VoxCeleb2017, ) Audio - 100 Time-S 1x512x300 Public
-OPPORTUNITY (Roggen2010, ) IMU 12 11 Time-S 77x1x100 Public
+miniImageNet (Vinyals et al., 2016) Image - 100 - 3x84x84 Public
+miniDomainNet (Zhou et al., 2021) Image - 126 - 3x224x224 Public
+Office-31 (Saenko et al., 2010) Image - 31 - 3x224x224 Public
+Office-Home (Venkateswara et al., 2017) Image - 65 - 3x224x224 Public
+Caltech-101 (Fei-Fei et al., 2004) Image - 101 - 3x224x224 Public
+Newsgroup (Thulasidasan et al., 2021) Text - 20 - 100x1x1000 Public
+VoxCeleb (Nagrani et al., 2017) Audio - 100 Time-S 1x512x300 Public
+OPPORTUNITY (Roggen et al., 2010) IMU 12 11 Time-S 77x1x100 Public
 Target
-HHAR (Stisen2015, ) IMU 9 6 Time-S 6x1x256 Public
-WESAD (Schmidt2018, ) ECG 15 3 Time-S 10x1x256 Public
+HHAR (Stisen et al., 2015) IMU 9 6 Time-S 6x1x256 Public
+WESAD (Schmidt et al., 2018) ECG 15 3 Time-S 10x1x256 Public
 Gesture IMU/PPG 10 8 Time-S 7x1x200 Private
@@
 BP mmWave 10 - Doppler 1x1x499 Private
-ChestX (Wang2017, ) Image - 5 - 1x1024x1024 Public
+ChestX (Wang et al., 2017) Image - 5 - 1x1024x1024 Public
   Backbone Layers L-Blocks GFLOPs Params Kernel Shape
-  ResNet18 (chen2018a, ) 18 9 3.67 11.18M 2D 3x224x224
-ResNet18-1D (chen2018a, ) 18 9 0.39 3.9M 1D 100x1x1000
-Multi-ResNet18 (Sicheng2020, ) 18x5 9 18.35 55.9M 2D 3x224x224
-Conv4 (chen2018a, ) 4 4 0.02 0.07 1D 1x1x500
-Conv5 (metasense, ) 5 5 0.19 30.42M 1D By Input
-GPT2 (zhou2023one, ) 4x6 6 0.41 82.9M 1D By Input
+  ResNet18 (Chen et al., 2019) 18 9 3.67 11.18M 2D 3x224x224
+ResNet18-1D (Chen et al., 2019) 18 9 0.39 3.9M 1D 100x1x1000
+Multi-Conv4 (Zhao et al., 2020) 4x5 4 0.3 1.15M 2D 3x32x32
+Multi-ResNet10 (Zhao et al., 2020) 10x5 5 0.7 24.5M 2D 3x84x84
+Multi-ResNet18 (Zhao et al., 2020) 18x5 9 18.35 55.9M 2D 3x224x224
+Conv4 (Chen et al., 2019) 4 4 0.02 0.07 1D 1x1x500
+Conv5 (Gong et al., 2019) 5 5 0.19 30.42M 1D By Input
+GPT2 (Zhou et al., 2023) 4x6 6 0.41 82.9M 1D By Input
 Table 3: DL model backbone specifications.
-We benchmark XTransfer against state-of-the-art baselines, grouped into single- and multi-source methods (Tab.
-1).
-Evaluation is conducted on 8 widely-used public datasets for cross-modality model transfer (Tab.
-2), using their corresponding public pre-trained models as sources, from diverse modalities (e.g., image, text, audio, and sensing) with standard backbones such as ResNet18 from PyTorch Hub (PyTorchhub, ) (Tab.
-3).
-To evaluate XTransfer, we develop 4 real-world applications and build three testbeds on 3 commercial edge devices (e.g., smartwatch, smartphone, raspberry pi).
-Based on these, we collect 4 human sensing datasets (marked as Private in Tab.
-2) as targets, from 40 participants in real-world settings, with ethics approval from the Human Research Ethics Committee of our institute.
+6.1 Experiment setup
+We benchmark XTransfer against SOTA baselines, grouped into single- and multi-source methods (Table 1).
+Evaluation is conducted on 8 widely-used public datasets under cross-modality FSL settings (Table 2), using their corresponding public pre-trained models as sources, from diverse modalities (e.g., image, text, audio, and sensing) with various backbones such as ResNet18 from PyTorch Hub (Hub, 2022) (Table 3).
+To evaluate XTransfer, we develop 4 real-world applications and build testbeds on 3 commercial edge devices (e.g., smartwatch, smartphone, Raspberry Pi).
+Based on these, we collect 4 human sensing datasets (marked as Private in Table 2) as targets, from 40 participants in real-world settings, with ethics approval from the Human Research Ethics Committee of our institute.
 Additionally, we include 2 public sensing datasets and 1 image dataset as targets for standard benchmarking.
-We apply standard Leave-One-Out Cross-Validation (LOOCV) (metasense, ) for evaluating human sensing datasets, and adopt accuracy-to-resource (ATR) ratio, defined as $ATR=Accuracy\mathbin{/}(\alpha Norm(FLOPs))+(1-\alpha)Norm(Params))$ for evaluating resource-accuracy efficiency, where $\alpha$ is set to 0.5 by default.
-For more details, see Appendix A.
+We apply standard Leave-One-Out Cross-Validation (LOOCV) (Gong et al., 2019) for evaluating human sensing datasets, and adopt accuracy-to-resource (ATR) ratio, defined as $ATR=Accuracy\mathbin{/}(\alpha Norm(FLOPs))+(1-\alpha)Norm(Params))$ for evaluating resource-accuracy efficiency, where $\alpha$ is set to 0.5 by default (see Appendix A for details).
 6.2 XTransfer performance
@@
 We test with 4 different levels of MMC dimensionality (i.e., number of channels) ranging from 64 to 512.
-We employ N-Pair loss (Sohn2016, ) and Triplet loss (Schroff2015, ) as baselines.
+We employ N-Pair loss (Zhang et al., 2023c) and Triplet loss (Weller et al., 2022) in FSL sensing tasks as baselines.
 We run 15 rounds for each setting by default throughout the evaluation.
-Fig.
-4(a) presents that our Repair loss consistently achieves the highest S-score of 0.2 on average, outperforming both N-Pair loss (0.05) and Triplet loss (0.07), across all levels of MMC dimensionality.
+Figure 4(a) presents that our Repair loss consistently achieves the highest S-score of 0.2 on average, outperforming both N-Pair loss (0.05) and Triplet loss (0.07), across all levels of MMC dimensionality.
 It indicates that Repair loss effectively enhances layer performance.
@@
 We now evaluate LWS control performance regarding model accuracy and layer search efficiency.
-We employ HHAR as target, ResNet18 on miniImageNet and Multi-ResNet18 on 5 image datasets as sources shown in Tab.
-2, with a default search depth of 3 in 3-, 5- and 10-shot settings.
-We also utilize MetaSense in Tab.
-1 as the oracle baseline.
+We employ HHAR as target, ResNet18 on miniImageNet and Multi-ResNet18 on 5 image datasets as sources shown in Table 2, with a default search depth of 3 in 3-, 5- and 10-shot settings.
+We also utilize MetaSense in Table 1 as the oracle baseline.
 We compare different configurations: ResNet18 (Single), Multi-ResNet18 (Multi) without efficient search enabled, and Multi-ResNet18 (Multi-Efficient) with efficient search enabled.
-As expected, Fig.
-4(b) shows that Multi-Efficient achieves a substantial time reduction ranging from 2.1 to 4 times compared to Multi in 5- and 10-shot settings, while outperforming the oracle baseline.
+As expected, Figure 4(b) shows that Multi-Efficient achieves a substantial time reduction ranging from 2.1 to 4 times compared to Multi in 5- and 10-shot settings, while outperforming the oracle baseline.
 Additionally, Multi-Efficient improves average accuracy by 7.8% over Multi across the three settings.
@@
 Since the selection of pre-trained source models plays a crucial role in XTransfer, we evaluate how different source models from various modalities affect the accuracy and resource overhead of the output models.
-We use HHAR, 2 pre-trained ResNet18s on both miniImageNet (Image) and VoxCeleb (Audio), and 2 pre-trained ResNet18-1Ds on both Newsgroup (Text) and OPPORTUNITY (Sensing), as shown in Tab.
-2.
+We use HHAR, 2 pre-trained ResNet18s on both miniImageNet (Image) and VoxCeleb (Audio), and 2 pre-trained ResNet18-1Ds on both Newsgroup (Text) and OPPORTUNITY (Sensing), as shown in Table 2.
 We keep the other experimental setups unchanged.
-Fig.
-4(c) shows that the output models using Image source achieve the highest accuracy of 70.5% on average, compared to that of 69.6%, 64.3%, and 67% achieved by the models using Text, Audio, Sensing as sources, respectively.
+Figure 4(c) shows that the output models using Image source achieve the highest accuracy of 70.5% on average, compared to that of 69.6%, 64.3%, and 67% achieved by the models using Text, Audio, Sensing as sources, respectively.
 This suggests that the latent features learned from Image have better discriminability than those from other modalities.
@@
 6.3 Comprehensive comparison
-Cross-modality model transfer performance.
-We evaluate the overall performance of XTransfer regarding both accuracy and ATR across 5 sensing datasets and 1 image dataset as targets, as shown in Tab.
-2.
-We compare XTransfer against a range of state-of-the-art baselines shown in Tab.
-1.
+Cross-modality FSL performance.
+We evaluate the overall performance of XTransfer regarding both accuracy and ATR across 5 sensing datasets and 1 image dataset as targets (Table 2).
+We compare XTransfer against a range of SOTA baselines shown in Table 1.
 We also mark our methods as Our-Single and Our-Multi, using pre-trained ResNet18 on miniImageNet and Multi-ResNet18 with 5 corresponding image datasets as sources.
-In addition, we evaluate the fine-tuning performance of GPT2-small for comparison and align the parameters with the work (zhou2023one, ).
-All methods on target datasets are evaluated under the same cross-modality FSL settings (see App.
-A), with MetaSense (metasense, ) serving as the oracle baseline, trained on Conv5 with sufficient target sensor data (Tab.
-3).
+In addition, we evaluate the fine-tuning performance of GPT2-small for comparison and align the parameters with the work (Zhou et al., 2023).
+All methods on target datasets are evaluated under the same cross-modality FSL settings (see Appendix A), with MetaSense (Gong et al., 2019) serving as the oracle baseline, trained on Conv5 with sufficient target sensor data (Table 3).
 For consistency, we also train the oracle using ResNet18 on the Emotion dataset to match 2D kernels.
-As shown in Tab.4 and 5, both our methods outperform the baselines regarding model accuracy and ATR, and successfully reach or surpass the oracle in model accuracy across all target datasets in 5- and 10-shot settings.
+As shown in Table 4 and 5, both our methods outperform the baselines regarding model accuracy and ATR, and successfully reach or surpass the oracle in model accuracy across all target datasets in 5- and 10-shot settings.
 Specifically, both Our-Single and Our-Multi achieve a significantly higher ATR on average, surpassing the existing single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively.
-Interestingly, in 3-shot setting, the averaged model accuracy of our methods is slightly lower than that of the oracle on the HHAR, Gesture, and Emotion datasets.
-This suggests that XTransfer’s Repair loss may lead to less stable improvements for latent feature alignment when using extremely low sample sizes, but its performance remains strong and stable in 5- and 10-shot settings.
-In summary, the comprehensive results show that XTransfer achieves state-of-the-art model accuracy, substantially reducing the cost of data collection and model resource overhead for scalable deployment of human sensing applications.
+In 3-shot setting, our method achieves slightly lower accuracy than the oracle on HHAR, Gesture, and Emotion.
+This indicates that extremely low samples may not fully represent the test data distribution, leading to less stable feature alignment.
+Since the layers are reused from pre-trained sources, the performance also depends on their quality and semantic overlap with targets.
+Notably, the flexibility of XTransfer allows further gains in such low-shot cases by using additional source models with higher quality or closer relevance.
+In summary, the results show that XTransfer achieves SOTA accuracy while substantially reducing data and resource costs, providing a scalable and adaptable method for human sensing at the edge.
 Training resource usage.
-We also examine the training resource usage of XTransfer by measuring GPU utilization (Util) on the cloud.
-We focus on the above baselines that require model re-training or adaptation for transfer, e.g., MAML (Finn2017, ) and DAPN (Zhao2021, )), excluding methods that rely solely on fine-tuning or distance-based techniques.
+We also examine the training resource usage of XTransfer by measuring GPU utilization (Util) in the cloud.
+We focus on the above baselines that require model re-training or adaptation for transfer, e.g., MAML (Chen et al., 2019), DAPN (Zhao et al., 2021), and SemiCMT (Chen et al., 2024)), excluding methods that rely solely on fine-tuning or distance-based techniques.
 The experimental setup remains the same as above.
-As shown in Tab.
-6, both our methods achieve the lowest GPU Util across all n-shot settings compared to baselines.
+As shown in Table 6, both our methods achieve the lowest GPU Util across all n-shot settings compared to baselines.
 Notably, both Our-Single and Our-Multi achieve a substantial reduction in GPU Util by 2.6 to 3.6 times and 2.5 to 3.5 times on average, respectively, compared to baselines.
@@
 It also highlights the effectiveness of our compact connector design.
-Despite leveraging multiple source models, Our-Multi shows no significant increase (e.g., only 3.67%) in training resource usage, further highlighting its efficiency.
-Hence, XTransfer significantly reduces cloud resource usage, enabling scalable model transfer with low computational overhead.
+Despite leveraging multiple source models, Our-Multi shows no significant increase (e.g., only 3.67%) in training resource usage, further highlighting its efficiency and scalability.
+Hence, XTransfer significantly reduces cloud resource usage, enabling scalable model transfer.
 HHAR (%) WESAD (%) Gesture (%) Writing (%) Emotion (%) ChestX (%)
 Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
-MetaSense (metasense, ) 63.2 69.0 75.2 55.1 64.0 68.3 66.3 73.4 79.4 72.1 83.3 89.6 55.5 56.3 65.5 24.8 28.1 31.7
-ProtoNet (Snell2017, ) 50.6 45.7 49.8 43.9 49.3 52.0 50.8 55.8 59.2 69.5 78.7 73.4 51.3 49.2 54.8 21.7 23.4 24.5
-DAPN (Zhao2021, ) 48.7 51.2 54.7 56.9 61.2 63.1 45.8 49.4 50.5 75.1 78.6 79.6 48.0 50.0 58.7 24.0 24.4 25.5
-MAML (Finn2017, ) 36.6 42.3 42.4 39.6 57.9 65.6 37.2 41.3 39.4 31.6 39.7 35.2 22.3 26.9 31.1 22.5 20.4 21.9
-GPT2 (zhou2023one, ) 41.0 34.0 45.0 48.0 35.0 36.0 49.0 54.0 63.0 52.0 71.0 75.0 - - - - - -
+MetaSense (Gong et al., 2019) 63.2 69.0 75.2 55.1 64.0 68.3 66.3 73.4 79.4 72.1 83.3 89.6 55.5 56.3 65.5 24.8 28.1 31.7
+ProtoNet (Chen et al., 2019) 50.6 45.7 49.8 43.9 49.3 52.0 50.8 55.8 59.2 69.5 78.7 73.4 51.3 49.2 54.8 21.7 23.4 24.5
+DAPN (Zhao et al., 2021) 48.7 51.2 54.7 56.9 61.2 63.1 45.8 49.4 50.5 75.1 78.6 79.6 48.0 50.0 58.7 24.0 24.4 25.5
+MAML (Chen et al., 2019) 36.6 42.3 42.4 39.6 57.9 65.6 37.2 41.3 39.4 31.6 39.7 35.2 22.3 26.9 31.1 22.5 20.4 21.9
+SemiCMT (Chen et al., 2024) 33.3 38.9 48.9 37.3 40.0 44.4 23.5 34.4 42.5 27.3 38.6 48.0 24.0 33.6 46.7 25.3 25.6 28.0
+GPT2 (Zhou et al., 2023) 41.0 34.0 45.0 48.0 35.0 36.0 49.0 54.0 63.0 52.0 71.0 75.0 - - - - - -
 Our-Single 59.0 71.8 80.7 77.9 78.4 81.8 45.2 69.6 77.8 72.8 87.0 91.3 43.6 55.6 61.4 27.7 28.6 30.4
-MDDA (Sicheng2020, ) 13.2 17.7 18.8 40.0 39.3 41.9 11.2 13.9 15.2 12.5 10.9 9.0 14.4 15.3 16.0 20.6 19.6 19.9
-MCW (NEURIPS20196048ff4e, ) 27.6 25.4 26.9 30.7 32.7 32.4 22.3 23.4 25.6 32.3 31.6 30.2 25.3 27.2 26.5 22.4 21.5 22.8
+MDDA (Zhao et al., 2020) 13.2 17.7 18.8 40.0 39.3 41.9 11.2 13.9 15.2 12.5 10.9 9.0 14.4 15.3 16.0 20.6 19.6 19.9
+MCW (Lee et al., 2019) 27.6 25.4 26.9 30.7 32.7 32.4 22.3 23.4 25.6 32.3 31.6 30.2 25.3 27.2 26.5 22.4 21.5 22.8
 Our-Multi 59.3 74.3 78.4 76.4 77.8 78.7 48.7 73.1 78.6 74.1 86.1 90.4 42.3 55.1 58.6 29.0 30.0 30.2
@@
 Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
-ProtoNet (Snell2017, ) 0.51 0.46 0.50 0.44 0.49 0.52 0.51 0.56 0.59 0.69 0.79 0.73 0.51 0.49 0.55 0.22 0.23 0.25
-DAPN (Zhao2021, ) 0.49 0.51 0.55 0.57 0.61 0.63 0.46 0.49 0.51 0.75 0.79 0.80 0.48 0.50 0.59 0.24 0.24 0.26
-MAML (Finn2017, ) 0.37 0.42 0.42 0.40 0.58 0.66 0.37 0.41 0.39 0.32 0.40 0.35 0.22 0.27 0.31 0.22 0.20 0.22
-GPT2 (zhou2023one, ) 0.11 0.09 0.12 0.13 0.09 0.09 0.13 0.14 0.17 0.14 0.19 0.20 - - - - - -
+ProtoNet (Chen et al., 2019) 0.51 0.46 0.50 0.44 0.49 0.52 0.51 0.56 0.59 0.69 0.79 0.73 0.51 0.49 0.55 0.22 0.23 0.25
+DAPN (Zhao et al., 2021) 0.49 0.51 0.55 0.57 0.61 0.63 0.46 0.49 0.51 0.75 0.79 0.80 0.48 0.50 0.59 0.24 0.24 0.26
+MAML (Chen et al., 2019) 0.37 0.42 0.42 0.40 0.58 0.66 0.37 0.41 0.39 0.32 0.40 0.35 0.22 0.27 0.31 0.22 0.20 0.22
+SemiCMT (Chen et al., 2024) 0.33 0.39 0.49 0.37 0.40 0.44 0.24 0.34 0.43 0.27 0.39 0.48 0.24 0.34 0.47 0.25 0.26 0.28
+GPT2 (Zhou et al., 2023) 0.11 0.09 0.12 0.13 0.09 0.09 0.13 0.14 0.17 0.14 0.19 0.20 - - - - - -
 Our-Single 1.42 1.57 2.09 2.76 2.61 2.51 1.55 1.90 1.76 2.11 1.88 1.84 0.82 1.07 1.26 0.83 1.04 1.24
-MDDA (Sicheng2020, ) 0.03 0.04 0.04 0.08 0.08 0.08 0.02 0.03 0.03 0.03 0.02 0.02 0.03 0.03 0.03 0.04 0.04 0.04
-MCW (NEURIPS20196048ff4e, ) 0.06 0.05 0.05 0.06 0.07 0.06 0.04 0.05 0.05 0.06 0.06 0.06 0.05 0.05 0.05 0.04 0.04 0.05
+MDDA (Zhao et al., 2020) 0.03 0.04 0.04 0.08 0.08 0.08 0.02 0.03 0.03 0.03 0.02 0.02 0.03 0.03 0.03 0.04 0.04 0.04
+MCW (Lee et al., 2019) 0.06 0.05 0.05 0.06 0.07 0.06 0.04 0.05 0.05 0.06 0.06 0.06 0.05 0.05 0.05 0.04 0.04 0.05
 Our-Multi 1.71 1.47 1.64 3.00 2.70 2.75 1.20 1.25 1.63 1.93 1.96 1.53 0.96 1.23 1.73 0.68 1.24 0.83
@@
 Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
-DAPN (Zhao2021, ) 48.85 49.19 51.79 44.62 49.49 48.81 42.73 48.71 50.59 43.88 56.63 57.51 44.67 59.73 63.12 46.66 47.32 48.48
-MAML (Finn2017, ) 59.59 79.32 89.72 52.58 70.27 77.52 53.11 66.75 73.21 72.56 85.47 92.22 63.39 72.78 75.01 44.52 55.46 61.36
+DAPN (Zhao et al., 2021) 48.85 49.19 51.79 44.62 49.49 48.81 42.73 48.71 50.59 43.88 56.63 57.51 44.67 59.73 63.12 46.66 47.32 48.48
+MAML (Chen et al., 2019) 59.59 79.32 89.72 52.58 70.27 77.52 53.11 66.75 73.21 72.56 85.47 92.22 63.39 72.78 75.01 44.52 55.46 61.36
+SemiCMT (Chen et al., 2024) 55.60 73.43 78.91 38.83 50.33 67.89 57.02 72.16 98.04 75.14 79.12 81.79 60.30 63.43 68.60 41.54 50.05 63.43
 Our-Single 18.27 18.90 19.48 15.95 16.46 16.98 20.67 21.71 22.51 19.92 20.67 21.08 18.24 19.20 19.96 16.97 17.58 18.80
@@
 7 Conclusion
-This paper presents XTransfer, a novel method that enables resource-efficient, modality-agnostic model transfer for human sensing applications at the edge.
+This paper presents XTransfer, a novel method that enables modality-agnostic few-shot model transfer with resource-efficient design for human sensing applications.
 XTransfer harnesses the power of pre-trained models as free sources and requires only few sensor data to create compact models.
-It significantly reduces the costs associated with sensor data collection, model training, and edge deployment, offering a practical and scalable method for advancing human sensing.
-Limitations and future work.
-Since XTransfer reuses layers from one or more pre-trained source models, its performance depends on their quality (c.f.
-Sec.
-4).
-While performance varies with different source models, XTransfer’s modality-agnostic design ensures broad adaptability, enabling flexible use of diverse source models without concern for modality shift or resource constraints.
-For fair benchmarking, we primarily use the pre-trained ResNet18 from PyTorch Hub (PyTorchhub, ) (c.f.
-Sec.
-6.1), but XTransfer’s layer-wise repairing and recombining support flexible backbone choices.
-While current version requires source models with the same backbone, future versions will support mixed backbones to further enhance flexibility.
+It significantly reduces the costs associated with sensor data collection, model training, and edge deployment, offering a scalable and adaptable method for advancing human sensing.
 Appendix
@@
 A.1 Settings for learning with few data
-A typical FSL method consists of two stages–meta-training and meta-testing [56].
-In meta-training stage, FSL requires a large-scale source dataset from the same singular modality (i.e., typically specific to the application) to train source models from scratch.
-The source dataset is divided into a support set and a query set  [^4] The support set and query set are assumed to be drawn from the same data distribution [82]..
-Data for each set is loaded according to the n-way (i.e., number of classes) and n-shot (i.e., number of samples per class) format [15].
+A typical FSL method consists of two stages–meta-training and meta-testing (Chen et al., 2019).
+In meta-training stage, FSL requires a large-scale source dataset from the same modality (i.e., typically specific to the application) to train source models from scratch.
+The source dataset is divided into a support set and a query set  [^4] The support set and query set are assumed to be drawn from the same data distribution (Triantafillou et al., 2020)..
+Data for each set is loaded according to the n-way (i.e., number of classes) and n-shot (i.e., number of samples per class) format (Song et al., 2023).
 The meta-testing stage follows the same structure but uses unseen target data.
 Specifically, the target support set is used to adapt the source models, which are then evaluated on the target query set.
-Leave-One-Out Cross-Validation (LOOCV) is widely used as the de facto standard for evaluating datasets involving multiple users in human sensing [18].
+Leave-One-Out Cross-Validation (LOOCV) is widely used as the de facto standard for evaluating datasets involving multiple users in human sensing (Gong et al., 2019).
 LOOCV works by leaving one user out as the target for meta-testing while using the remaining users to train the source models during meta-training in each round.
-The MetaSense [18] is used as an oracle baseline under FSL settings, where sufficient target sensing datasets are provided during meta-training stage.
-Unlike FSL, cross-domain FSL focuses primarily on meta-testing and transfers pre-trained source models across domains (i.e., different users), while remaining within the same singular modality [23].
-In this work, the cross-modality FSL settings are defined similarly to cross-domain FSL, but with pre-trained source models selected across different modalities without constraints.
+The MetaSense (Gong et al., 2019) is used as an oracle baseline under FSL settings, where sufficient target sensing datasets are provided during meta-training stage.
+Unlike FSL, cross-domain FSL focuses primarily on meta-testing and transfers pre-trained source models across domains (i.e., different users), while remaining within the same modality (Oh et al., 2022).
+In this work, the cross-modality FSL settings are defined similarly to cross-domain FSL, but extend it to transfer pre-trained source models across different modalities.
 A.2 Setups in preliminary study
-As shown in Fig.
-2(a), we test a standard overfitting metric, defined as $\frac{\left|Accuracy_{train}-Accuracy_{test}\right|}{Accuracy_{train}}$ [51].
-In Fig.
-2(b), we test 3 source models (using a pre-trained ResNet18 from PyTorch Hub [38]) from 3 modalities, e.g., miniImageNet (Image), Newsgroup (Text), and OPPORTUNITY (Sensing), shown in Tab.
-2.
-We also use 5 sensing datasets (HHAR, WESAD, Writing, Emotion, and Gesture) as targets, and employ 2 well-known image datasets (CUB and ChestX) as reference targets to the Image, detailed in Tab.
-2.
+As shown in Figure 2(a), we test a standard overfitting metric, defined as $\frac{\left|Accuracy_{train}-Accuracy_{test}\right|}{Accuracy_{train}}$ (López et al., 2022).
+In Figure 2(b), we test 3 source models (using a pre-trained ResNet18 from PyTorch Hub (Hub, 2022)) from 3 modalities, e.g., miniImageNet (Image), Newsgroup (Text), and OPPORTUNITY (Sensing), shown in Table 2.
+We also use 5 sensing datasets (HHAR, WESAD, Writing, Emotion, and Gesture) as targets, and employ 2 well-known image datasets (CUB and ChestX) as reference targets to the Image, detailed in Table 2.
 We use a default 5-shot cross-modality FSL setting with LOOCV enabled.
@@
 In Section 3, we conduct a layer-wise analysis to gain insights into cross-modality model transfer with few sensor data.
-To set up, we continue to use the standard ResNet18 [38] and segment it as 9 L-Blocks, as shown in Tab.
-3.
-We also employ the default reshaping (c.f.
-App.
-B.1) on HHAR as target with 5-shot setting, and use the original source model input (miniImageNet) as reference, marked as Original.
-We examine the single-source baselines including Transfer Learning (TL) [68], Fine-tuning (FT) [23], MAML [70], DAPN [52], and MetaSense [18] as the oracle baseline, shown in Tab.
-1.
+To set up, we continue to use the standard ResNet18 (Hub, 2022) and segment it as 9 L-Blocks, as shown in Table 3.
+We also employ the default reshaping (see Section B.1) on HHAR as target with 5-shot setting, and use the original source model input (miniImageNet) as reference, marked as Original.
+We examine the single-source baselines, including Transfer Learning (TL) (Jiang et al., 2022), Fine-tuning (FT) (Oh et al., 2022), MAML (Chen et al., 2019), DAPN (Zhao et al., 2021), and MetaSense (Gong et al., 2019) as the oracle baseline, shown in Table 1.
 To precisely track the MMC shift changes at each layer during fine-tuning,
 A.4 Layer-wise metric
-To precisely track the MMC shift changes at each layer during fine-tuning, we essentially use the Silhouette score (S-score) [63] as the layer-wise metric to evaluate how well the features are being fine-tuned.
+To precisely track the MMC shift changes at each layer during fine-tuning, we essentially use the Silhouette score (S-score) (Shahapure & Nicholas, 2020) as the layer-wise metric to evaluate how well the features are being fine-tuned.
 The S-score is in the range from -1 to 1 and being better if closing to 1.
-To verify whether the S-score indicates the layer performance, we observe the changes in the S-score and accuracy at each layer using the TL [68] and Original methods.
-Fig.
-6(a) also shows that S-score and accuracy by Original are positively correlated, converging in a layer-wise manner.
+To verify whether the S-score indicates the layer performance, we observe the changes in the S-score and accuracy at each layer using the TL (Jiang et al., 2022) and Original methods.
+Figure 6(a) also shows that S-score and accuracy by Original are positively correlated, converging in a layer-wise manner.
 TL’s S-score also performs equivalently to its accuracy (i.e., S-score drop along with accuracy drop) in detecting the negative transfer across different layers.
@@
 We also benchmark XTransfer with a full development cycle of cross-modality FSL tasks, including pre- and post-processing, training, and inference on commercial edge devices.
-Once pre-trained source models and few sensor data are processed in cloud via XTransfer, compact models can be seamlessly converted to various DL platforms (e.g., PyTorch Mobile, TFLite) for broader edge deployments.
+Once pre-trained source models and few sensor data are processed in the cloud via XTransfer, compact models can be seamlessly converted to various DL platforms (e.g., PyTorch Mobile, TFLite) for broader edge deployments.
 To enable reproducibility of the results, we fix all random seeds in the code, set the evaluation mode consistently when evaluating output models, and enable CUDA convolution determinism.
 To measure model resource overhead (e.g., FLOPs and parameters), we use the standard PyTorch OpCounter.
-For regression tasks, class-based transformation [83] enables compatibility.
+For regression tasks, class-based transformation (Shi et al., 2022) enables compatibility.
 A.6 Developed sensing applications
-To evaluate XTransfer on real-world human sensing targets, we essentially develop 4 human sensing applications, as shown in Fig.
-5.
+To evaluate XTransfer on real-world human sensing targets, we essentially develop 4 human sensing applications, as shown in Figure 5.
 These include: (1) emotion recognition via facial expressions (i.e., Emotion) and (2) non-contact blood pressure monitoring (i.e., BP) based on an embedded mmWave radar setup, (3) handwriting tracking (i.e., Writing) using ultrasound based on a smartphone, and (4) gesture recognition (i.e., Gesture) using IMU and PPG sensors based on a smartwatch.
 The data collection is approved with ethics approval and involves 40 subjects.
-The corresponding target sensing datasets are shown in Tab.
-2 with detailed specification.
+The corresponding target sensing datasets are shown in Table 2 with detailed specifications.
 A.7 Hardware setup and computation environment
-We evaluate XTransfer on real-world testbeds using 3 commercial edge devices, ranging from smartwatch to Raspberry Pi (Tab.
-7).
-For example, Fig.
-5(a) presents an embedded mmWave radar setup which consists of a TI IWR1843Boost and a Raspberry Pi 4 using ROS Melodic on Ubuntu 18.04.4.
+We evaluate XTransfer on real-world testbeds using 3 commercial edge devices, ranging from smartwatch to Raspberry Pi (Table 7).
+For example, Figure 5(a) presents an embedded mmWave radar setup which consists of a TI IWR1843Boost and a Raspberry Pi 4 using ROS Melodic on Ubuntu 18.04.4.
 In addition, we implement both smartwatch and smartphone testbeds using PyTorch Android version 1.10.0.
@@
 A.8 Training details
-In the Repair stage, we fine-tune the generative transfer module (c.f.
-Sec.
-4) for 100 episodes using SGD with a learning rate of 1e-2 and momentum 0.95.
+In the Repair stage, we fine-tune the generative transfer module (see Section 4) for 100 episodes using SGD with a learning rate of 1e-2 and momentum 0.95.
 We apply a step-based learning rate scheduler with a step size of 20 and a decay factor of 0.5.
 We also use standard early stopping with a patience of 20.
-We retain the training parameters of the baselines in Tab.
-1 as specified in their original implementations.
+We retain the training parameters of the baselines in Table 1 as specified in their original implementations.
 Figure 5: (a) Embedded mmWave radar testbed setup; (b)-(e) Built human sensing applications across different real-world settings.
@@
 To align with source model input shape, we develop a default reshaping to transform sensor data shape.
-It uses bilinear interpolation [84] (i.e., Resizer) to resize the height and width, and a fixed convolutional layer (i.e., Fixed header) with a $1\times 1$ kernel to match the input channels [85], placed between source models and sensor data.
+It uses bilinear interpolation (Mastyło, 2013) (i.e., Resizer) to resize the height and width, and a fixed convolutional layer (i.e., Fixed header) with a $1\times 1$ kernel to match the input channels (Szegedy et al., 2014), placed between source models and sensor data.
 B.2 PCA-based channel removal
-In the Removal stage, channel removal is designed to further enhance layer performance while optimizing resource usage.
+In the Removal stage, channel removal is designed to further enhance layer-wise performance while optimizing resource usage.
 This is achieved by removing unnecessary channels at each layer.
-Existing pruning techniques [46, 47] typically rely on channel importance metrics, such as the L2 norm, calculated from MMC.
+Existing pruning techniques (Shen et al., 2022; Frankle & Carbin, 2019) typically rely on channel importance metrics, such as the L2 norm, calculated from MMC.
 However, due to MMC shift, these metrics may inaccurately assess channel importance, resulting in degraded layer performance.
@@
 B.3 Setups for efficient search insights
-To gain design insights of efficient search for LWS control (c.f.
-Sec.
-5), we obverse a complete process where all layers are repaired using the HHAR dataset and a Multi-ResNet18 model (i.e., 5 source models with 45 L-unit), as shown in Tab.
-3.
-Fig.
-6(b) shows the changes in the S-score across three states: after pairing anchors (i.e., Anchor), before and after applying the SRR pipeline (i.e., Before and After, respectively), in a layer-wise manner.
+To gain design insights of efficient search for LWS control (see Section 5), we observe a complete process where all layers are repaired using the HHAR dataset and a Multi-ResNet18 model (i.e., 5 source models with 45 L-unit), as shown in Table 3.
+Figure 6(b) shows the changes in the S-score across three states: after pairing anchors (i.e., Anchor), before and after applying the SRR pipeline (i.e., Before and After, respectively), in a layer-wise manner.
 The results show that the SRR pipeline successfully increases the S-score from Before state (i.e., lower bound) to align with the Anchor state (i.e., the objective).
@@
 We then repeat the process multiple times to observe the changes in $rate_{n}$.
-Fig.
-6(c) shows that the repair rate demonstrates exponential growth.
-Thus, we design the repair rate growth model based on an exponential growth model (c.f.
-Sec.
-5.2).
+Figure 6(c) shows that the repair rate demonstrates exponential growth.
+Thus, we design the repair rate growth model based on an exponential growth model (see Section 5.2).
 B.5 Dynamic search range mechanism
@@
 Appendix C Additional experiment results
+C.1 Feature space alignment performance
+In the process of feature space alignment (Section 4.1), we use a standard Hungarian algorithm based on default distance metrics to select and pair classes.
+To verify its effectiveness, we keep using the setup of HHAR, the standard pre-trained ResNet18 on miniImageNet, and the standard PCA under 5-shot settings to examine how our repair loss performs with/without the feature space alignment, especially in the initial layer repairing.
+Repairing the initial layer is critical, as it often introduces a large, disorganized feature space and significantly affects the repair of subsequent layers.
+In Table 8, the results show that with the alignment, the repair loss decreases considerably faster by 1.87 times on average, compared to the case without alignment.
+This indicates that the design of feature space alignment notably contributes to the repairing process.
+Epoch (#) #1 #20 #40 #60 #80 #100
+With Alignment 0.486 0.355 0.350 0.339 0.333 0.331
+Without Alignment 0.476 0.406 0.393 0.383 0.378 0.387
+Table 8: Feature space alignment performance across training epochs.
+C.2 Impact of PCA variations
+Since our method uses the repair learning mechanism that relies on backpropagation (BP) through the reduced-orthogonal feature space, standard PCA can project data into a linear subspace and maintain a closed-form, differentiable structure, which enables smooth integration with BP in our end-to-end learning pipeline.
+Compared to advanced choices such as kernel PCA, we find that it introduces a non-linear mapping into a high-dimensional feature space using a kernel function, and its projections are generally not differentiable with respect to input features.
+This makes it incompatible with our pipeline.
+To test whether linear PCA variations affect the overall performance of XTransfer, we compare standard PCA against sparse PCA using the setup of HHAR based on the ResNet18 under 3 few-shot settings with a default search depth of 3.
+As shown in Table 9 (#2 denotes 2-component setting), sparse PCA reduces learning time slightly by 11.6% on average, but suffers from a notable accuracy drop by 4.4% on average.
+These results support the choice of standard PCA as a balanced and practical solution.
+PCA Variation 3-shot 5-shot 10-shot
+Accuracy (%) Time (mins) Accuracy (%) Time (mins) Accuracy (%) Time (mins)
+Standard PCA-#2 58.96 3.9 71.85 5.6 80.74 10.3
+Sparse PCA-#2 58.52 3.8 64.67 5.3 75.11 8.8
+Table 9: Impact of PCA variations on accuracy and training time.
+C.3 Impact of PCA components
+We then evaluate the impact of PCA components on the model accuracy and time efficiency of XTransfer.
+Since the feature space with different dimensionalities adjusted by PCA components may impact our method performance, we test 3 different levels of PCA components, including PCA-#2 (top two components), PCA-80% (selected components covering 80% of the feature variance), and PCA-Max (maximum components available).
+We continue to use the same setup as above with the standard PCA.
+The results in Table 10 show that as the number of PCA components increases, the training time rises notably.
+In fact, PCA-Max results in an average accuracy drop of 4.5% compared to PCA-80%.
+By contrast, PCA-#2 achieves the best accuracy-to-time ratio across all n-shot settings, with competitive accuracy and the lowest training time, hence we use PCA-#2 as the default throughout the evaluation.
+PCA Components 3-shot 5-shot 10-shot
+Accuracy (%) Time (mins) Accuracy (%) Time (mins) Accuracy (%) Time (mins)
+PCA-#2 58.96 3.9 71.85 5.6 80.74 10.3
+PCA-80% 63.61 5.2 74.72 7.6 80.00 12.9
+PCA-Max 60.00 7.2 75.56 8.1 70.28 16.9
+Table 10: Impact of PCA components on accuracy and training time.
 Figure 7: Ablation study evaluating the performance of components, search parameters, and applications.
-C.1 Removal performance
-We evaluate how channel removal enhances the average S-score of selected layers.
-We use HHAR and the standard pre-trained ResNet18 on miniImageNet in a 5-shot setting, and enable channel removal with a default search depth of 3.
-We employ L2 norm-based (i.e., L2-based) model pruning [47, 50] as the baseline and compare it with the setting where channel removal is disabled (i.e., Non-Pruning).
-Fig.
-7(a) demonstrates that the proposed channel removal outperforms both Non-Pruning and L2-based pruning across all settings, achieving an average S-score of 0.1 by selected layers.
+C.4 Removal performance
+We next evaluate how channel removal enhances the average S-score of selected layers.
+We continue to use the same setup as above under 5-shot settings.
+We employ L2 norm-based (i.e., L2-based) model pruning (Frankle & Carbin, 2019; Shen et al., 2022) as the baseline and compare it with the setting where channel removal is disabled (i.e., Non-Pruning).
+Figure 7(a) demonstrates that the proposed channel removal outperforms both Non-Pruning and L2-based pruning across all settings, achieving an average S-score of 0.1 by selected layers.
 L2-based pruning achieves a lower average S-score than Non-Pruning, indicating that the channel importance may be inaccurately estimated, resulting in degraded performance.
 These results show that the proposed channel removal further improves layer performance by accurately removing unnecessary channels at each layer.
-C.2 Impact of search depth
+C.5 Impact of search depth
 In this experiment, we investigate how different search depths affect the accuracy and resource overhead (e.g., FLOPs) of output models.
 Varying the search depth alters the search window sizes with the number of layer candidates, which can lead to different output model backbones and accuracy.
-To quantify these differences, we employ HHAR as target, ResNet18 on miniImageNet (Single) and Multi-ResNet18 (Multi) on 5 image datasets as sources (shown in Tab.
-2) in 3-, 5- and 10-shot settings.
-We also utilize MetaSense in Tab.
-1 as the oracle baseline.
+To quantify these differences, we employ HHAR as target, ResNet18 on miniImageNet (Single) and Multi-ResNet18 (Multi) on 5 image datasets as sources (shown in Table 2) in 3-, 5- and 10-shot settings.
+We also utilize MetaSense in Table 1 as the oracle baseline.
 Both Single and Multi models have up to 9 L-unit, hence we set three different search depths as D1, D3, D9.
-Fig.
-7(b) shows that the output models of both Single-D1 and Multi-D1 result in high FLOPs, while that of both Single-D9 and Multi-D9 achieves low FLOPs.
+Figure 7(b) shows that the output models of both Single-D1 and Multi-D1 result in high FLOPs, while that of both Single-D9 and Multi-D9 achieve low FLOPs.
 As search depth decreases, the search process becomes longer, and it triggers LWS control to make more decisions on recombining layers, which results in a suboptimal model backbone with high FLOPs.
@@
 Therefore, we set D3 as the default search depth, as it provides the best balance between resource overhead and accuracy of the output models.
-C.3 Regression task performance
-We also evaluate the effectiveness of XTransfer in a regression task using the BP dataset, which includes both systolic (SBP) and diastolic (DBP) sensor data shown in Tab.
-2 collected by the blood pressure monitoring application.
+C.6 Impact of model backbones
+The current implementation of XTransfer uses homogeneous pre-trained source models.
+Technically, the layer-wise repairing and recombination techniques impose no fundamental restrictions on the choice of model backbone and work with various architectural variants.
+To verify this, we evaluate the trade-off between accuracy and computational cost using different source backbones.
+We continue to use HHAR, and test Multi-Conv4, Multi-ResNet10, and Multi-ResNet18 in a 5-shot, multi-source setting shown in Table 3.
+The results in Table 11 demonstrate that the average accuracy declined, transitioning from Multi-Conv4 to Multi-ResNet10, and then increasing with Multi-ResNet18.
+Interestingly, Multi-Conv4, despite having fewer parameters, outperforms Multi-ResNet10, which may be attributed to differences in model quality or distributional divergence between source and target.
+Multi-ResNet18, though with higher FLOPs, achieves the best accuracy.
+It suggests that larger models trained on large-scale cross-modality datasets exhibit superior latent features, which notably contribute to such model transfer.
+Our techniques such as the connectors are designed to flexibly handle the layers with different shapes/structures, and can be extended to heterogeneous model backbones.
+In addition, our method focuses on latent feature alignment rather than specific feature extraction.
+It remains fully compatible with advanced source models (e.g., with temporal encoders), which can be seamlessly integrated into our pipelines with minimal or no modifications.
+Backbone Accuracy (%) FLOPs (G)
+Multi-Conv4 65.56 0.01
+Multi-ResNet10 63.61 0.14
+Multi-ResNet18 78.44 2.19
+Table 11: Performance of different source model backbones in terms of accuracy and FLOPs.
+C.7 Training efficiency
+To test the practical training efficiency in the cloud, we further evaluate training time across the selected baselines.
+We use the same experimental setups as in Table 6, Section 6.3.
+For MAML (Chen et al., 2019) and SemiCMT (Chen et al., 2024), which require model re-training from pre-trained source models to adapt to the target modality classes, training time includes both re-training and fine-tuning.
+For DAPN (Zhao et al., 2021), it reflects model adaptation process on few target data.
+The results in Table 12 show that both Our-Single and Our-Multi achieve significant training time reductions compared to MAML.
+Our-Single also matches or exceeds DAPN and SemiCMT in training efficiency.
+Notably, even when using up to 5 source models, Our-Multi remains practical training efficiency, requiring only 2.05 times more training time on average compared to Our-Single across all datasets and shot settings.
+This further supports the practicability and scalability of our method.
+HHAR (mins) WESAD (mins) Gesture (mins) Writing (mins) Emotion (mins) ChestX (mins)
+Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
+DAPN (Zhao et al., 2021) 2.38 4.62 11.78 1.49 2.54 6.78 3.36 6.12 15.72 4.12 7.71 20.97 2.99 7.34 18.86 2.14 4.57 12.04
+MAML (Chen et al., 2019) 124.28 125.66 183.20 71.71 77.07 115.87 118.79 129.79 191.95 165.06 129.27 194.36 94.81 98.40 153.20 66.20 72.17 106.97
+SemiCMT (Chen et al., 2024) 3.02 8.27 10.23 1.82 2.48 5.63 4.81 9.42 12.83 10.10 13.92 18.75 3.39 7.32 11.48 2.84 5.28 8.33
+Our-Single 3.98 5.57 7.95 2.12 2.99 4.33 5.64 7.80 11.24 7.94 10.54 14.69 5.58 7.43 9.88 3.65 4.75 6.84
+Our-Multi 9.25 15.33 15.25 4.40 5.38 7.21 12.49 15.12 24.37 15.35 17.36 32.38 11.72 13.39 19.75 8.41 9.33 13.61
+Table 12: Comparison in training time of each method.
+C.8 Regression task performance
+We also evaluate the effectiveness of XTransfer in a regression task using the BP dataset, which includes both systolic (SBP) and diastolic (DBP) sensor data shown in Table 2 collected by the blood pressure monitoring application.
 We employ the pre-trained ResNet18 on miniImageNet as the source model and use the 5-shot setting.
 To integrate this into the SRR pipeline, we transform the regression task by treating individual regression labels as distinct classes for anchor class pairing.
-For benchmarking, we use mean error (ME) and standard deviation (STD), following the Advancement of Medical Instrumentation (AAMI) standard, which accepts an ME within 5 mmHg and STD within 8 mmHg [86].
-We employ a random forest (RF) based method [87] as the oracle baseline.
-Besides, we train a Conv4 backbone in Tab.
-3 (Conv) using LOOCV, and the fine-tuned Conv (ConvFT) as the baselines.
-Fig.
-7(c) shows that both XTransfer and the RF methods surpass the AAMI standard.
+For benchmarking, we use mean error (ME) and standard deviation (STD), following the Advancement of Medical Instrumentation (AAMI) standard, which accepts an ME within 5 mmHg and STD within 8 mmHg (Stergiou et al., 2018).
+We employ a random forest (RF) based method (Shi et al., 2023) as the oracle baseline.
+Besides, we train a Conv4 backbone in Table 3 (Conv) using LOOCV, and the fine-tuned Conv (ConvFT) as the baselines.
+Figure 7(c) shows that both XTransfer and the RF methods surpass the AAMI standard.
 In particular, XTransfer outperforms the baselines and RF in estimating both SBP (ME: -0.64mmHg, STD: 7.2mmHg) and DBP (ME: 0.07mmHg, STD: 7.47mmHg) tasks.
 These results demonstrate that XTransfer is highly efficient in handling regression tasks with few sensor data.
-C.4 On-device performance
-We now evaluate the on-device performance of deploying XTransfer’s output models on 3 commercial edge devices (Tab.
-7), focusing on resource overhead metrics such as FLOPs, parameters, latency per call, and peak memory footprint.
-For each target dataset, we test both types of output models created by our single-source (Single) and multi-source (Multi) methods.
-We also use ResNet18 and Multi-ResNet18 (Tab.
-3) as baselines.
-As shown in Tab.
-8, all output models achieve significant reductions in model resource overhead, e.g., FLOPs is reduced by 2.4 to 5.7$\times$ for Single models and by 10.1 to 16.5$\times$ for Multi models, compared to the baselines.
-These recombined models also demonstrate substantial improvement in on-device performance.
-Specifically, in the Single method, latency is reduced by 1.4 to 4.1 $\times$ and peak memory footprint decreases by 1.8 to 3.9$\times$.
-In the Multi method, the reductions are even more pronounced, with latency dropping by 7.4 to 21$\times$ and memory footprint by 2.2 to 8.4$\times$, compared to the baselines.
-These results prove that XTransfer successfully achieves streamlined edge deployment.
+C.9 On-device performance
+[…]
 Watch Pi Phone

### v3

@@
 However, its training and development are hindered by the limited availability of sensor data and resource constraints of edge systems.
-While transferring pre-trained models to different sensing applications is promising, existing methods often require extensive sensor data and computational resources, resulting in high costs and poor adaptability in practice.
+While transferring pre-trained models to different sensing applications is promising, existing methods often require extensive sensor data and computational resources, resulting in high costs and limited transferability.
 In this paper, we propose XTransfer, a first-of-its-kind method enabling modality-agnostic, few-shot model transfer with resource-efficient design.
-XTransfer flexibly uses single or multiple pre-trained models and transfers knowledge across different modalities by (i) model repairing that safely mitigates modality shift by adapting pre-trained layers with only few sensor data, and (ii) layer recombining that efficiently searches and recombines layers of interest from source models in a layer-wise manner to create compact models.
+XTransfer flexibly uses pre-trained models and transfers knowledge across different modalities by (i) model repairing that safely mitigates modality shift by adapting pre-trained layers with only few sensor data, and (ii) layer recombining that efficiently searches and recombines layers of interest from source models in a layer-wise manner to restructure models.
 We benchmark various baselines across diverse human sensing datasets spanning different modalities.
-Comprehensive results demonstrate that XTransfer achieves state-of-the-art performance while significantly reducing the costs of sensor data collection, model training, and edge deployment.
+The results show that XTransfer achieves state-of-the-art performance while significantly reducing the costs of sensor data collection, model training, and edge deployment.
+Human Sensing, Model Transfer, Edge Deployment
 1 Introduction
-Human sensing refers to the process of capturing and interpreting data related to human activities, behaviors, and physiological states using various sensors (Zhang et al., 2023a; Ystgaard et al., 2023).
-With the proliferation of edge devices equipped with sensors, human sensing plays a vital role in edge systems to understand contexts and enable smart applications, ranging from activity recognition to emotion or vital sign detection (Ahamed & Farid, 2018).
+Human sensing refers to the process of capturing and interpreting data related to human activities, behaviors, and physiological states using various sensors (Zhang et al., 2023a).
+With the proliferation of edge devices equipped with sensors, human sensing plays a vital role in edge systems to understand contexts and enable smart applications, ranging from activity recognition to emotion or vital sign detection (Ahamed and Farid, 2018).
 Deep learning (DL) offers robust performance in interpreting sensor data, and its deployment on edge devices improves privacy and reduces bandwidth (Zhang et al., 2020; Chen et al., 2020).
@@
 Unlike other data modalities (e.g., vision or text), collecting human sensing data for training can be costly and even impractical.
-This is uniquely due to the data that presents brittleness (i.e., noise sensitivity, low SNR), inconsistency (i.e., user variance, scenario changes), and heterogeneity (i.e., differences in hardware configuration) (Lane & Georgiev, 2015; Jeyakumar et al., 2019; Teh et al., 2020).
+This is uniquely due to the data that presents brittleness (i.e., noise sensitivity, low SNR), inconsistency (i.e., user variance, scenario changes), and heterogeneity (i.e., differences in hardware configuration) (Lane and Georgiev, 2015; Jeyakumar et al., 2019; Teh et al., 2020).
 Collecting such data typically requires ethical approvals, as it involves sensitive information (e.g., location, motion patterns) that raises privacy concerns (Zhang et al., 2022).
@@
 Importantly, human sensing spans various modalities (e.g., IMU, ultrasound, mmWave radar) and data types (e.g., spectrograms, Doppler profiles, time series), further amplifying the costs of applying DL solutions and naturally leading to cross-modality scenarios (Xiao et al., 2022; Zhang et al., 2023a).
-Prior research has explored Few-shot Learning (FSL) (Wang et al., 2024), transfer learning (Dhekane & Ploetz, 2024), and their combination (i.e., cross-domain FSL) (Thukral et al., 2025; Yin et al., 2024) to notably reduce data collection costs by adapting pre-trained models to sensing applications.
-However, they still rely on large-scale labeled datasets from the same modality to costly train source models from scratch (Song et al., 2023), or face challenges such as modality shift and poor adaptability when leveraging pre-trained models (Oh et al., 2022; Tao et al., 2022; Guo et al., 2022), particularly in cross-modality settings.
-Recent advances in multi-modal learning (Yang et al., 2023; Chen & Yang, 2025; mengxi Chen et al., 2024; Madaan et al., 2024; Yang et al., 2024; Baldenweg et al., 2024) (e.g., CLIP-based models (Girdhar et al., 2023; Wu et al., 2023; Moon et al., 2023)) and cross-modal learning (e.g., image-to-sensor distillation (Zhao et al., 2018; Song et al., 2022; Gurbuz et al., 2020; Zhang et al., 2023c) or alignment-based transfer (Kamboj et al., 2025; Chen et al., 2024)), demonstrate the feasibility of knowledge transfer across modalities, but they remain modality-specific (i.e., relying on shared semantic feature spaces), and typically require large-scale paired data with resource-intensive training and deployment, leading to high costs and making them impractical for few-shot adaptation to support human sensing at the edge.
-These limitations motivate the need for a scalable and adaptable model transfer that can broadly adapt pre-trained models across modalities (e.g., vision or text) to new sensing modalities (e.g., radar or bio-signals) with few data.
+Prior works on Few-shot Learning (FSL) (Wang et al., 2024), transfer learning (Dhekane and Ploetz, 2024), and cross-domain FSL (Thukral et al., 2025; Yin et al., 2024) reduce data collection costs by adapting pre-trained models to sensing applications.
+However, they typically rely on large-scale labeled datasets within the same modality to costly train source models from scratch (Song et al., 2023), or face challenges such as modality shift with limited transferability when leveraging pre-trained models (Oh et al., 2022; Tao et al., 2022; Guo et al., 2022), particularly in cross-modality settings.
+Recent advances in multi-modal learning (Yang et al., 2023; Chen and Yang, 2025; mengxi Chen et al., 2024; Madaan et al., 2024; Yang et al., 2024; Baldenweg et al., 2024) (e.g., CLIP-based models (Girdhar et al., 2023; Wu et al., 2023; Moon et al., 2023)) and cross-modal learning (e.g., image-to-sensor distillation (Zhao et al., 2018; Song et al., 2022; Gurbuz et al., 2020; Zhang et al., 2023c) or alignment-based transfer (Kamboj et al., 2025; Chen et al., 2024)), further demonstrate the feasibility of knowledge transfer across modalities.
+However, they remain modality-specific (i.e., relying on shared semantic feature spaces), and typically require large-scale unlabeled and paired data with resource-intensive training and deployment, leading to high costs and making them impractical for few-shot adaptation to support human sensing at the edge.
+These gaps motivate a scalable and adaptable model transfer that can broadly reuse pre-trained models across modalities (e.g., vision or text) for new sensing modalities (e.g., radar or bio-signals) using only few data.
 In practice, however, addressing this need is uniquely challenging.
-Due to substantial differences in data characteristics—such as shape, distribution, and feature representation—between source and target modalities (i.e., modality shift), transferred models often suffer from significantly degraded performance (i.e., negative transfer) (Guo et al., 2022; Oh et al., 2022).
+Due to substantial differences in data characteristics, such as shape, distribution, and feature representation, between source and target modalities (i.e., modality shift), transferred models often suffer from significantly degraded performance (i.e., negative transfer) (Guo et al., 2022; Oh et al., 2022).
 Our experiments reveal that state-of-the-art (SOTA) baselines (Table 1) fail to reach oracle performance (Gong et al., 2019) and suffer from substantial overfitting when applied to human sensing tasks with few data, falling notably short of the goal, as shown in Figure 2(a).
-These highlight the underlying challenge of latent feature misalignment across modalities.
+These highlight the underlying challenge of latent feature distribution misalignment across modalities.
 Also, these methods often neglect resource constraints, rendering them impractical for edge deployment, especially in advanced multi-source settings (Zhao et al., 2020; Yue et al., 2021a; Lee et al., 2019).
-While methods such as neural architecture search (NAS) (Wen et al., 2023; Han et al., 2021), pruning (Frankle & Carbin, 2019; Shen et al., 2022), or quantization (Polino et al., 2018), aim to reduce resource demands, they often fail to maintain performance under modality shift and few data.
+While model structuring methods such as neural architecture search (NAS) (Wen et al., 2023; Han et al., 2021), pruning (Frankle and Carbin, 2019; Shen et al., 2022), or quantization (Polino et al., 2018), aim to reduce resource demands, they often fail to maintain performance under modality shift and few data.
 As shown in Figure 2(c), our experiments using pruning (Shen et al., 2022) demonstrate significant model accuracy loss by 50.5% on average.
-Figure 1: Overview.
-XTransfer transfers source models across modalities with few sensor data through model repairing (SRR pipeline) and layer recombining (LWS control).
-LWS control first segments source models into layers and operates layer-wise search across pools.
-At each pool, the pre-search check decides which layers need repairing, then SRR pipeline repairs them and LWS control selects layers of interest.
-These layers are incrementally recombined during the search, resulting in a compact model for enabling human sensing at the edge.
-Subfigures (a)–(c) illustrate the feature space evolution before and after repairing.
 Driven by the increasing availability of high-quality pre-trained models from public repositories (e.g., PyTorch Hub (Hub, 2022)), we propose XTransfer, a first-of-its-kind method enabling modality-agnostic, few-shot model transfer.
-It flexibly leverages public pre-trained models and transfers knowledge across different modalities (e.g., image, text, audio, or sensing) to human sensing tasks using only few sensor data, significantly reducing the costs of large-scale data collection and training from scratch.
-Its resource-efficient design further optimizes edge deployment, enabling both scalability and adaptability to diverse human sensing applications.
-XTransfer uniquely repairs, searches and recombines layers of interest from pre-trained models under cross-modality FSL settings (see Appendix A), resulting in compact models.
-To mitigate modality shift and prevent overfitting, it enables model repairing using a novel Splice-Repair-Removal (SRR) pipeline (see Section 4).
-SRR pipeline transforms a few sensor data to broadly adapt pre-trained layers by repairing layer-wise misalignment and harmonizing latent feature distributions across modalities.
-XTransfer also features layer recombining using a novel Layer-Wise Search (LWS) control mechanism (see Section 5) to identify and recombine layers of interest by conducting a layer-wise search.
-Building on the insight that not all repaired layers contribute, LWS control discards such layers and recombines only useful ones to optimize model compactness using principles from NAS (White et al., 2023).
-To further accelerate the search in multi-source settings, a pre-search check strategy is designed to efficiently avoid unnecessary repairing.
-The main contributions are summarized as: 1) a first-of-its-kind method, XTransfer, that enables modality-agnostic, few-shot model transfer with resource-efficient design for human sensing on edge systems, as shown in Figure 1; 2) a novel SRR pipeline that effectively mitigates modality shift and prevents overfitting, harmonizes latent feature distributions across modalities; 3) a novel LWS control mechanism that efficiently searches layers of interest from multiple sources in a layer-wise manner to create compact models; and 4) extensive experimental evaluations (see Section 6) under cross-modality FSL settings, benchmarking SOTA baselines (Table 1) on 8 public datasets with corresponding pre-trained source models across diverse modalities (e.g., image, text, audio, and sensing) and 7 target datasets (Table 2) using 3 commercial edge devices.
-The results show that XTransfer achieves a significantly enhanced accuracy-to-resource (ATR) ratio (see Section 6.1), outperforming single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively, while reducing layer search time by 2.1 to 4 times and substantially lowering the costs of data collection, model training, and edge deployment, demonstrating broad adaptability.
-Figure 2: Preliminary study.
-(a) reveal baseline performance gap  [^2] We test a pre-trained ResNet18 on miniImageNet as source and HHAR dataset as target under a 5-shot setting (see Appendix A)..
-(b) shows the average similarity and FSL difficulty across all sensing datasets to source modalities (e.g., Image, Text, Sensing) using default reshaping (see Appendix B.1).
-2 distinct areas represent similarity levels (A–hard, B–normal).
-Key findings: 1) compared to CUB, similarity levels across modalities are notably low, e.g., Text and Sensing fall into Area A, indicating a significant modality shift; 2) compared to Image-Unpair (i.e., no class pairing) in Area A, Image surprisingly falls into Area B, indicating that pairing classes may enhance cross-modality similarity (motivating our techniques in Section 4.1); 3) compared to Text and Sensing, Image exhibits more stable standard deviations and lower FSL difficulty, suggesting better potential for broad adaptability.
-(c) shows model pruning bottlenecks.
+It flexibly leverages public pre-trained models and transfers knowledge across modalities (e.g., image, text, audio, or sensing) to human sensing tasks using only few labeled sensor data with no additional unlabeled data (i.e., cross-modality FSL settings, Appendix A.1), significantly reducing the costs of large-scale data collection and training from scratch.
+Its resource-efficient design also optimizes both cloud performance and edge deployment, enabling scalable and adaptable method across human sensing applications.
+Instead of relying on shared semantic spaces or paired cross-modal data, XTransfer mitigates modality shift by model repairing at the level of layer-wise latent feature distribution misalignment using activation-based statistics.
+This layer-wise formulation is motivated by our observation (Section 3) that modality shift “damages” intermediate representations unevenly across layers, making global adaptation with few data unstable and prone to overfitting.
+It is enabled by a Splice–Repair–Removal (SRR) pipeline (Section 4) that (i) splices heterogeneous layer inputs, (ii) repairs layer misalignment via an anchor-based generative transfer module in a reduced-orthogonal feature space, and (iii) removes unnecessary channels after repair.
+Building on the insight that not all repaired layers contribute, XTransfer further features layer recombining using a Layer-Wise Search (LWS) control (Section 5) that selects and recombines layers of interest (i.e., useful repaired layers) through an NAS-inspired, resource-constrained layer-wise search, incrementally enhancing the repairing and optimizing model structure.
+To accelerate and stabilize LWS in multi-source settings with large candidate pools, we incorporate a pre-search check strategy coupled with a dynamic search range mechanism to avoid unnecessary repairing, balancing accuracy and efficiency.
+The main contributions are summarized as:
+  • A first-of-its-kind method, XTransfer, that enables modality-agnostic, few-shot model transfer with resource-efficient design for human sensing on edge systems.
+  • A SRR pipeline that mitigates modality shift and prevents overfitting via layer-wise repairing of latent feature distribution misalignment.
+  • A LWS control that enables efficient, stable layer-wise search, selectively recombining layers of interest while discarding others for model restructuring.
+  • Extensive experimental evaluations (Section 6) under cross-modality FSL settings, benchmarking SOTA baselines (Table 1) on various source and target datasets (Table 2), showing consistent accuracy and efficiency gains.
+Figure 1: Preliminary study under cross-modality FSL settings  [^1] We test a pre-trained ResNet18 on miniImageNet as source and HHAR dataset as target under a 5-shot setting with LOOCV enabled (details in Appendix A.2)..
+(a) reveals baseline performance gap.
+(b) shows average similarity and FSL difficulty across all target sensing datasets in Table 2 to source modalities (e.g., Image, Text, Sensing) using default reshaping (Appendix B.1) and benchmarking (Oh et al., 2022) (Appendix A.2 for details).
+Two distinct areas represent similarity levels (A–hard, B–normal).
+Key findings: 1) compared to CUB, similarity levels across modalities are notably low, e.g., Text and Sensing fall into Area A, indicating a significant modality shift; 2) compared to Image-Unpair (i.e., no class pairing) in Area A, Image surprisingly falls into Area B, indicating that pairing classes  [^2] The number of classes in the source model should be greater than that in the target.
+may enhance cross-modality similarity; 3) Image exhibits more stable standard deviations and lower FSL difficulty, suggesting better potential for model transfer.
+(c) shows significant model accuracy loss using pruning.
 2 Related Work
@@
 However, testing on benchmarks (Liang et al., 2021; Oh et al., 2022), our results reveal that they suffer from severe overfitting under cross-modality FSL settings.
-LLM-based data augmentation (Leng et al., 2024; 2023) helps reduce data collection costs, but remains modality-specific and orthogonal to our focus.
-Existing attempts on human sensing tasks (Gong et al., 2019; Wang et al., 2024; Thukral et al., 2025; Yin et al., 2024) remain constrained by a strong reliance on learning in the same modality, resulting in costly source data collection and training with limited adaptability.
-In contrast, XTransfer enables layer-wise model repairing to safely mitigate modality shifts using only few sensor data, achieving significant cost reduction with broad adaptability.
-Resource constraints optimization.
-Model compression techniques such as model pruning (Frankle & Carbin, 2019; Shen et al., 2022), quantization using low-precision data types (Polino et al., 2018), model merging by fine-tuning shared layers (Padmanabhan et al., 2023), one-size-fits-all model (Cai et al., 2020), and NAS techniques such as model structure optimization (Wen et al., 2023; Han et al., 2021) or search using few data (Eustratiadis et al., 2024; Xu et al., 2022), have shown success in reducing model resource overhead.
+LLM-based data augmentation (Leng et al., 2024, 2023) helps reduce data collection costs, but remains modality-specific and orthogonal to our focus.
+Recent advances introduce foundation models (FM) (Abbaspourazad et al., 2024; Weng et al., 2024) or prompt-based FM adaptation (Li et al., 2025) for human sensing, but these methods assume either a pre-existing target-modality FM or access to large-scale unlabeled and paired data for feature alignment.
+Existing attempts in human sensing FSL (Gong et al., 2019; Wang et al., 2024; Thukral et al., 2025; Yin et al., 2024) remain constrained by a strong reliance on learning in the same modality, resulting in costly source data collection and training.
+In contrast, XTransfer enables layer-wise model repairing to safely mitigate modality shifts using only few sensor data, achieving significant cost reduction.
+Model structuring for edge deployment.
+Model structuring techniques such as pruning (Frankle and Carbin, 2019; Shen et al., 2022), quantization using low-precision data types (Polino et al., 2018), model merging by fine-tuning shared layers (Padmanabhan et al., 2023), one-size-fits-all model (Cai et al., 2020), and NAS techniques such as model restructuring (Wen et al., 2023; Han et al., 2021) or search using few data (Eustratiadis et al., 2024; Xu et al., 2022), have shown success in reducing model resource overhead.
 However, they strongly rely on either sufficient target datasets or minimal modality shifts to maintain model performance.
-XTransfer uniquely applies layer recombining to optimize model structures while strengthening layer-wise dependence, by efficiently searching and recombining layers of interest from pre-trained source models at scale.
-It not only creates compact models for streamlined edge deployment but also further enhances model repairing.
-3 Preliminary
+XTransfer uniquely applies layer recombining to optimize model structures while strengthening repaired layer-wise dependence at scale under cross-modality FSL settings.
+It not only restructures models for streamlined edge deployment but also enhances model repairing.
 Figure 2: Design insights.
 (a) Layer-wise accuracy convergence using baselines is disrupted due to modality shift.
-(b) A notable MMC shift emerges and grows with increasing layer index, i.e., accuracy increases while MMC shift drops in Area A and begins to drop as MMC shift largely grows in Area B.
-(c) After repairing, layer S-score improves, but stagnation occurs at certain layers.
-Insights in cross-modality FSL.
-We apply a standard cross-domain FSL benchmarking (Oh et al., 2022) to gain insights.
-It uses Earth Mover’s Distance (EMD) to measure similarity between source models from 3 modalities (i.e., Image, Text, Sensing) and targets including 5 human sensing and 2 well-known image datasets (Table 2).
-It also assesses FSL difficulty by fine-tuning source models on targets.
-Since source models involve a large number of classes  [^3] The number of classes in the source model should be greater than that in the target.
-(e.g., 100 classes in miniImageNet), we propose a class pairing strategy that pairs source and target classes exhibiting similar latent feature distributions across modalities.
-The key insights are shown in Figure 2(b).
+(b) In area A, accuracy rises as MMC shift stays low, indicating a small latent feature gap.
+In area B, MMC shift notably increases with layer index, where excessive latent feature deviation begins to reduce accuracy.
+(c) After repairing, S-score improves but stagnation occurs at certain layers.
+3 Preliminary
 Layer-wise analysis.
-We perform a layer-wise analysis to gain insights, focusing on accuracy changes and latent feature distributions at each layer.
-Since effective training relies on learning discriminative features (Islam et al., 2021), we benchmark latent feature distribution clustering and use Mean Magnitude of Channels (MMC) (Luo et al., 2022) of each layer as the default metric to quantify layer-wise performance, as it offers a lightweight measure for channel importance.
-We define a single layer or a dependent layer block (i.e., L-Blocks in Table 3)—where layers are not structurally separable—as the default segmentation unit (i.e., L-unit).
-Figure 3(a) shows that negative transfer (Zhang et al., 2023b) disrupts accuracy convergence across layers (i.e., layer-wise dependence), causing layer-wise misalignment and accuracy drops.
-These misalignments correlate with MMC shifts, where greater shifts correspond to larger accuracy drops shown in Figure 3(b), indicating inconsistent latent feature interpretation during fine-tuning.
-To track MMC shifts, we essentially employ the Silhouette score (S-score) (Shahapure & Nicholas, 2020), which captures inter- and intra-class distances and reflects the discriminability of layer-wise latent features.
-Challenges.
-We aim to minimize layer-wise MMC shifts under cross-modality FSL settings, but several technical challenges arise.
-First, it is hindered by structural and semantic gaps between sources and target data, where classes and semantics often differ.
-Second, the pre-trained feature space at each layer is highly complex due to the multidimensional and noisy nature of MMC across source models from different modalities.
-This complexity increases with layer depth, as the strong layer-wise dependence formed during source model training becomes increasingly rigid, making it difficult to repair layer-wise misalignment using only few sensor data.
-Third, Figure 3(c) shows that S-scores may stagnate or even drop at certain layers (i.e., ineffective layers), impairing overall convergence.
-It indicates that not all repaired layers contribute positively, motivating our LWS control design to selectively recombine layers of interest while discarding others.
-The selection must also account for on-device resource constraints to control layer growth, while the rapidly expanding search space with increasing layers and source models further challenges the layer search.
+It is widely used to diagnose representation quality and guide model restructuring (Sahoo et al., 2025).
+We examine how accuracy and latent feature discriminability change across layers.
+Since effective training relies on discriminative latent representations (Islam et al., 2021), we quantify layer-wise latent features using Mean Magnitude of Channels (MMC) (Luo et al., 2022) as a lightweight activation-based statistics.
+We also measure class clustering via the Silhouette score (S-score) (Shahapure and Nicholas, 2020) computed on MMC-based representations, which captures inter-/intra-class distances and reflects layer-wise discriminability.
+We segment the backbone into L-units, defined as a single layer or an inseparable dependent layer block (L-Blocks in Table 3) that cannot be structurally decoupled (e.g., residual blocks in ResNet).
+The experimental setup is detailed in Appendix A.3, and additional evidence on S-score/accuracy correlation is provided in Appendix A.4.
+Figure 2(a) shows that negative transfer (Zhang et al., 2023b) disrupts accuracy convergence across layers (i.e., impaired layer-wise dependence), leading to misalignment and accuracy drops under few-shot fine-tuning.
+These degradations correlate with MMC shifts, where larger shifts correspond to larger accuracy drops (Figure 2(b)).
+We adopt MMC as a modality-agnostic metric as it is defined purely on layer activations without assuming shared semantic spaces or paired data, remaining lightweight and stable under FSL settings, compared to kernel-based distribution metrics such as Maximum Mean Discrepancy (MMD) (Wang et al., 2023).
+Problem setup.
+We aim to minimize layer-wise MMC shifts under cross-modality FSL settings.
+This poses several challenges that motivate the design of SRR pipeline and LWS control: (i) cross-modality transfer must handle structural differences (e.g., input formats and feature shapes) and semantic discrepancies (e.g., non-overlapping label spaces), requiring an objective defined purely on latent representations (e.g., activation-based MMC statistics) rather than shared semantic spaces or paired data; (ii) the latent feature distribution misalignment is complicated by noisy representations and strong layer-wise dependencies inherited from pre-training, making few-shot repair brittle; and (iii) not all layers remain useful after repairing (Figure 2(c)), selectively recombining useful layers is critical and must be done efficiently and stably in multi-source settings with large candidate pools.
+Figure 3 overviews XTransfer.
+Figure 3: Overview.
+XTransfer transfers source models across modalities with few sensor data through model repairing (SRR pipeline) and layer recombining (LWS control).
+LWS control first segments source models into layers and operates layer-wise search across pools.
+At each pool, the pre-search check decides which layers need repairing, then SRR pipeline repairs them and LWS control selects layers of interest.
+These layers are incrementally recombined during the search, restructuring models for enabling human sensing at the edge.
+Subfigures (a)–(c) illustrate the feature space evolution before and after repairing.
 4 Model Repairing
-Enabled by SRR pipeline, this process safely repairs layer-wise misalignment by minimizing MMC shifts, effectively restoring layer performance from source models using few sensing inputs under cross-modality FSL settings.
-Three stages are designed, as shown in Figure 1.
-In the Splice stage, default reshaping (see Appendix B.1)—using fixed down/up-sampling—cannot be fine-tuned and may lose important features, especially when layer shapes vary during layer-wise search.
-We propose a compact connector placed between sensing inputs and layers, consisting of a Pre-header (adaptive convolutional layer), Resizer, and one encoder-decoder pair to bridge cross-modality structural gaps.
-In the Repair stage, each connector is fine-tuned by the generative transfer module to generate adaptive channels, preserve key sensing features, and minimize MMC shifts.
-Finally, the Removal stage applies channel removal (see Appendix B.2) to further enhance layer-wise performance.
+Enabled by SRR pipeline, this process safely repairs layer-wise misalignment by minimizing MMC shifts, restoring pre-trained layer performance under cross-modality FSL settings.
+As shown in Figure 3, SRR comprises three stages.
+Splice: instead of default reshaping using fixed up/down-sampling (Appendix B.1), which is non-trainable and can discard features, especially when layer shapes vary during layer-wise search, we propose a trainable, compact connector placed between heterogeneous layers to enforce shape compatibility, consisting of a Pre-header (adaptive convolutional layer), Resizer, and one encoder-decoder pair.
+Repair: each connector is fine-tuned by the generative transfer module to adapt layer channels and minimize MMC shifts.
+Removal: layer channel removal (Appendix B.2) is applied to further strengthen repairing.
 4.1 Repairing in complex feature space
-Since the original latent features from source models are well-trained for high discriminability, our intuition is to safely fine-tune few sensor data by aligning the misinterpreted latent features with the original features (i.e., anchors) to minimize MMC shifts.
+Since source layers are already highly discriminative, our intuition is to safely adapt few sensor data by aligning their misaligned latent feature distributions to the "original" source-layer distributions, which serve as anchors.
 We hence propose cross-modality anchor-based alignment, which aligns source MMCs as anchors (i.e., anchor MMCs) with sensing MMCs.
-To handle complex feature spaces, our insight is to reduce MMC dimensionality by preserving key channels and excluding irrelevant or noisy ones.
+Since the raw MMC feature space is high-dimensional, noisy and correlated across channels, our insight is to reduce MMC dimensionality by preserving key layer channels while suppressing noise.
 Reduced-orthogonal feature space.
-We project anchor MMCs into a lower-dimensional, orthogonal feature space using principal component analysis (PCA) (Valpola, 2015) (i.e., anchor PCA space).
-Unlike the raw MMC feature space, where channel dimensions are often noisy and correlated, PCA enforces orthogonality across components, removing redundancy and suppressing noise.
-This enables a compact subspace (i.e., defaulting to 2 components) where the projected data captures maximum variance and forms highly clustered distributions.
-The optimized component weights (i.e., projection coefficients) highlight the most important channels contributing to layer performance.
+For each layer, we project anchor MMCs into a low-dimensional, orthogonal feature space via PCA (Valpola, 2015) (i.e., anchor PCA space).
+PCA removes redundancy by decorrelating channel dimensions and suppresses noise, yielding a compact subspace (we use 2 components by default) that preserves maximal variance and typically forms highly clustered class distributions.
+The resulting projection coefficients (i.e., component weights) also highlight the most important layer channels that contribute to layer performance.
 Given high discriminability of the original layers, the projected anchor MMCs preserve high clustering performance, with class distributions tightly centered (i.e., Anchor-center), as shown in Figure 3(a).
-Once initialized for each layer, this anchor PCA space is reused to project sensing MMCs.
+Once initialized per layer, the same anchor PCA space is reused to project sensing MMCs.
 Feature space alignment.
-After projecting sensing MMCs, we observe that the sample distribution for each class, along with their centroid (i.e., Sensing-center), appears disorganized compared to the well-clustered Anchor-center at each layer, as shown in Figure 1(a).
-Specifically, the scale and orientation of Sensing-center are notably different from those of Anchor-center, caused by MMC shifts.
+After projection, we observe that the sample distribution for each class, along with their centroid (i.e., Sensing-center), appears disorganized compared to the well-clustered Anchor-center at each layer, as shown in Figure 3(a).
+Due to MMC shifts, the scale and orientation of Sensing-center are notably different from those of Anchor-center.
 To align the scale, we first define a scale function based on the mean shift in inter-class distance $\mathrm{InterD}(\cdot)$, formulated as $S={{Mean}({InterD}(Pro(fs_{ij})))}\mathbin{/}{{Mean}({InterD}(Pro(ft_{ij})))}$, where $Pro(\cdot)={PCA}({MMC}(\cdot))$ denotes the space project function, $fs$ and $ft$ represent the latent features from both source and sensing modalities, respectively.
 $i$ denotes the index of pre-trained models, and $j$ denotes the layer index in each model $i$.
-Next, to align the orientation, we design a rotation alignment function that minimizes the cosine angle (Nguyen & Bai, 2010) by fine-tuning a multi-dimensional rotation matrix $M_{rot}]$ with dimensions matching the number of PCA components:
-$\begin{split}\arg\min_{[M_{rot}]}Cos(Cen(Pro(fs_{ij})),Cen(Pro(ft_{ij})*S)\cdot[M_{rot}])\end{split}$ (1)
+Next, to align the orientation, we design a rotation alignment function that minimizes the cosine angle (Nguyen and Bai, 2010) by fine-tuning a multi-dimensional rotation matrix $M_{rot}]$ with dimensions matching the number of PCA components:
+$\scriptsize\begin{split}\arg\min_{[M_{rot}]}Cos(Cen(Pro(fs_{ij})),Cen(Pro(ft_{ij})*S)\cdot[M_{rot}])\end{split}$ (1)
 where $Cos(\cdot)$ denotes the cosine similarity function, and $Cen(\cdot)$ calculates the centroid of the projected distribution. Figure 3(b) shows the aligned results.
 Anchor class pairing.
-Building on the key finding that class pairing boosts latent feature alignment across modalities (see Section 3), we propose selecting source classes with the highest S-scores as anchor classes to pair with sensing classes to facilitate the anchor-based alignment.
-We minimize pairing shift (i.e., the sum distance across each pair) via a linear sum assignment solved using the standard Hungarian algorithm, outputting the optimal pairing set $pairST$.
+Building on the key finding that class pairing boosts latent feature alignment across modalities in Figure 2(b), we propose selecting source classes with the highest S-scores as anchor classes to pair with sensing classes by minimizing a pairing shift objective (i.e., sum of centroid distances across pairs).
+This is solved as a linear sum assignment via the standard Hungarian algorithm, outputting the optimal pairing set $pairST$.
+This pairing is fully optimized by similarity in the given feature space.
 4.2 Repairing layer-wise dependence
 Generative transfer module.
-To effectively restore layer-wise dependence, we design a generative transfer module that enhances the anchor-based alignment through generative learning.
-It consists of a generator (i.e., connector) and a discriminator—each frozen pre-trained layer paired with its anchor PCA space (Figure 1).
+To restore layer-wise dependence, we design a generative transfer module that optimizes the generator (i.e., connector) using an adversarial objective.
+The discriminator operates on each frozen pre-trained layer together with its anchor PCA space (Figure 3).
 Freezing preserves stable anchors and prevents overfitting.
-For each layer, given the optimal pairing set $pairST$, the generator is fine-tuned using the discriminator’s loss to minimize MMC shifts, and uniquely transforms sensor inputs to align with anchors.
+Given each pairing set $pairST$, the connector is fine-tuned to minimize MMC shifts, uniquely transforming sensor inputs to align with anchors.
 Anchor-based repair loss.
-Building on the generative transfer module, we propose an anchor-based repair loss function ($loss_{srr}$) to harmonize latent feature distributions across modalities.
-Given each layer’s anchor PCA space, our objective is twofold: 1) minimize the distance ($D(\cdot)$) between the centroid of projected anchor MMC distribution and that of the projected sensing MMC distribution for each class pair $(c_{s},c_{t})\in pairST$ (i.e., positive loss), and 2) maximize the distance between sensing MMC samples with different labels $(c_{t}^{-})$ (i.e., negative loss).
+We propose an anchor-based repair loss ($loss_{srr}$) to harmonize layer-wise latent feature distributions across modalities in the anchor PCA space.
+For each class pair $(c_{s},c_{t})\in pairST$, our objective (i) minimizes the distance $D(\cdot)$ between the centroids of the projected anchor and sensing MMC distributions (i.e., positive loss), and (ii) maximizes the distance between sensing MMC samples with different labels $(c_{t}^{-})$ (i.e., negative loss).
 Since the projected anchor MMCs are highly clustered, they can provide the maximum margin ($M_{max}$) between anchor clusters.
 To further improve the negative loss, we update the function by calculating the anchor-based margin $margin^{c}=InterD(Pro(fs^{c}_{ij}))-IntraD(Pro(fs^{c}_{ij}))$, where the $M_{max}$ is selected, and $IntraD(\cdot)$ measures the intra-class distance.
-Hence, our loss function is formulated as:
-$\begin{split}loss_{srr}=&\frac{1}{N}\sum^{N}_{(c_{s},c_{t})\in pairST}D(Cen(Pro(fs^{c_{s}}_{ij})),Cen(Pro(ft^{c_{t}}_{ij})))\\ +&\frac{1}{K}\sum^{K}_{(c_{t},c_{tk}^{-})}ReLU(M_{max}-D(Pro(ft^{c_{t}}_{ij}),Pro(ft^{c_{tk}^{-}}_{ij})))\end{split}$ (2)
+The loss function is formulated as:
+$\scriptsize\begin{split}loss_{srr}=&\frac{1}{N}\sum^{N}_{(c_{s},c_{t})\in pairST}D(Cen(Pro(fs^{c_{s}}_{ij})),Cen(Pro(ft^{c_{t}}_{ij})))\\ +&\frac{1}{K}\sum^{K}_{(c_{t},c_{tk}^{-})}ReLU(M_{max}-D(Pro(ft^{c_{t}}_{ij}),Pro(ft^{c_{tk}^{-}}_{ij})))\end{split}$ (2)
 where $K$ denotes the number of negative samples, and $N$ denotes the number of classes.
@@
 5 Layer Recombining
-Working in conjunction with SRR pipeline for repairing each layer, XTransfer leverages layer recombining enabled by LWS control to: 1) perform the pre-search check that accelerates the search process, and 2) effectively select and recombine layers of interest to form a compact model.
+Working with the SRR pipeline, XTransfer performs layer recombining via LWS control to (i) perform the pre-search check that accelerates the search, (ii) stabilize the search using the dynamic search range mechanism, and (iii) select and recombine layers of interest to restructure models.
 5.1 Search control design
-Following NAS principles (White et al., 2023), poor design choices (e.g., search space, actions, strategy) can disrupt the balance between model performance and resource efficiency.
-As layers of interest grow, resource overhead can increase exponentially, making dynamic search control essential under on-device resource constraints.
-To effectively manage, we define the search space as all layers from source models, structured into search pools containing layer candidates (i.e., repaired layers), each with width $I$ (number of models) and depth $J$ (number of L-units), forming a window of size $I\times J$.
-Search actions (init, continue, skip, cross) are defined to guide layer recombination within each designated search pool.
-Init selects the starting layer with a default fixed depth of 1.
-Continue and skip guide intra-model layer recombination, while cross enables inter-model layer recombination.
-Search strategy under resource constraints.
+Inspired by NAS (White et al., 2023), effective layer-wise search relies on properly designing the search space, actions, and strategy under on-device resource constraints.
+We hence define the search space as all layers from source models, organized into search pools containing layer candidates (i.e., repaired layers).
+Each pool forms a window of size $I\times J$, where $I$ is the number of source models (width) and $J$ is the number of L-units (depth).
+We also define four search actions: init selects a starting layer (default depth $1$); continue and skip perform intra-model recombination; and cross enables inter-model recombination.
+Resource-constrained search strategy.
 To effectively control search actions within each search pool, we propose a layer value function based on the S-score $Score(\cdot)$ to evaluate each repaired layer performance and the convergence between adjacent layers after each SRR process $SRR(\cdot)$.
-It is formulated as $V_{ij}=Score(SRR(Pro(ft_{ij}),MT_{ij},L_{ij}),Y_{t}))$, where $MT_{ij}$ denotes the connector and $L_{ij}$ represents each repaired layer.
-Under resource constraints, only the highest-valued repaired layer in each pool is selected.
-LWS control then checks whether its S-score exceeds that of the previous layer to maintain layer-wise dependence.
-If successful, the layer $L_{ij}$ is identified as the layer of interest and recombined, and the next pool window begins at layer index $j+1$.
-Otherwise, the current pool is discarded, and the search window shifts forward.
-Considering the constraints, we design a resource-constrained function that works in conjunction with $V_{ij}(\cdot)$, which is formulated as $RC(n)=exp(n/(L-2)-2)+1,n\in[1,~L]$, where $n$ denotes the current recombined layer index, and $L$ denotes the maximum number of L-units.
-Each layer’s adjusted overhead, $R(n)_{ij}$, is computed as $R(n)_{ij}=(Res(L_{ij})*RC(n))/Max(Res(P))$, where $Res(\cdot)$ calculates actual cost (e.g., FLOPs), and $Max(Res(P))$ is the peak cost in current pool $P$.
+It is formulated as $V_{ij}=Score(SRR(Pro(ft_{ij}),MT_{ij},L_{ij}),Y_{t}))$, where $MT_{ij}$ is the connector and $L_{ij}$ is each repaired layer.
+Under resource constraints, LWS control selects the highest-valued candidate in each pool and verifies that it improves layer-wise dependence by checking whether its S-score exceeds that of the previously selected layer.
+If selected, $L_{ij}$ becomes a layer of interest and is recombined, and the next pool starts from layer index $j+1$.
+Otherwise, the pool is discarded, and the window shifts forward.
+To incorporate the constraints, we define a resource-constrained function formulated as $RC(n)=exp(n/(L-2)-2)+1,n\in[1,~L]$ where $n$ is the current recombined layer index and $L$ is the maximum number of L-units.
+Each candidate’s adjusted overhead is computed as $R(n)_{ij}=(Res(L_{ij})*RC(n))/Max(Res(P))$, where $Res(\cdot)$ measures actual cost (e.g., weighted FLOPs and memory), and $Max(Res(P))$ is the maximum cost in current pool $P$.
 The resource-constrained layer value function becomes $VR_{ij}=V_{ij}\mathbin{/}R(n)_{ij}$.
@@
 Pre-search check.
-LWS control is designed to handle large search spaces from deep or multi-source models.
-While fine-tuning the connectors in SRR pipeline is efficient, the large search space can considerably slow down LWS control.
-We hence propose the pre-search check strategy to avoid unnecessary repair steps before applying SRR.
-We track S-score progression across three states—after anchor pairing (Anchor) (i.e., repairing objective), before SRR (Before), and after SRR (After) (i.e., repaired).
-SRR pipeline consistently improves S-scores from Before to After, approaching the Anchor level, indicating a strong correlation.
-Based on this, our intuition is to estimate repaired S-score (i.e., repair rate) of each layer based on Before and Anchor states to decide whether to bypass unnecessary repairs.
-Experiments show the actual repair rate grows exponentially with the recombined layer index (see Appendix B.4).
-We hence design a repair rate growth model formulated as $rate_{n}^{est}=exp(a*n)+b,n\in[1,L]$, where $a$ and $b$ are optimized via non-linear regression to minimize MSE to fit $rate_{n}^{est}$, with the objective function defined as $\arg\min_{a,b}\frac{1}{n^{{}^{\prime}}}\sum_{n=0}^{n^{{}^{\prime}}}(rate_{n}^{est}-rate_{n})^{2}$, where $n^{{}^{\prime}}$ is the number of recombined layers.
-The estimation begins after init.
-Once a new layer is selected, LWS control updates the estimation.
-Due to the limited updates, the estimation may be inaccurate.
-To mitigate, we also develop a dynamic search range mechanism that dynamically adjusts the search range in each pool (details in Appendix B.5).
-6 Experiment
+Although SRR fine-tunes connectors efficiently, applying SRR to every candidate in a large pool can still slow down LWS control.
+We hence propose the pre-search check strategy that avoids unnecessary repairs before applying SRR.
+We track S-score progression across three states: after anchor pairing (Anchor) (i.e., repairing objective), before SRR (Before), and after SRR (After).
+Empirically, SRR improves S-scores from Before to After, approaching the Anchor level, indicating a strong correlation (see Appendix B.3 for details).
+Based on this insight, our intuition is to estimate repaired S-score (i.e., repair rate) of each layer based on Before and Anchor states to decide whether to bypass unnecessary repairs.
+Experiments show the actual repair rate grows exponentially with the recombined layer index (Appendix B.4).
+We hence design a repair rate growth model formulated as $rate_{n}^{est}=exp(a*n)+b,n\in[1,L]$, where $a$ and $b$ are optimized via non-linear regression to minimize MSE to fit $rate_{n}^{est}$, with the objective defined as $\arg\min_{a,b}\frac{1}{n^{{}^{\prime}}}\sum_{n=0}^{n^{{}^{\prime}}}(rate_{n}^{est}-rate_{n})^{2}$, where $n^{{}^{\prime}}$ is the number of recombined layers.
+Estimation starts after init and is updated after each newly selected layer.
+Dynamic search range mechanism.
+Using $rate_{n}^{est}$, LWS can initially apply a narrow search range (e.g., only the top-1 candidate by estimated repair rate) to filter most candidates before SRR.
+However, since $rate_{n}^{est}$ is updated sparsely, estimation errors may cause missed optimal candidates.
+To stabilize search while retaining efficiency, we develop a dynamic search range mechanism that adjusts the search range, which is formulated as $range_{n}=(1+\hat{u}^{\pm}rate_{n}^{est})*range_{n-1},\,range_{n}\in[0.2,1]$, where the range starts at 0.5 and adjusts based on the previous range.
+Here, $\hat{u}$ is a signed unit vector.
+If $rate_{n-1}^{est}<rate_{n-1}$, it indicates that the estimation might be inaccurate, and $\hat{u}$ will be positive to increase the search range to repair more candidates and stabilize accuracy.
+Otherwise, the range will be decreased to bypass unnecessary repairs and improve efficiency.
 Table 1: State-of-the-art (SOTA) baselines.
   Baselines Group Description
-  TL (Jiang et al., 2022) Single-source A standard transfer learning method by freezing all layers
-FT (Oh et al., 2022) Single-source A standard fine-tuning method by updating all layers
-ProtoNet (Chen et al., 2019) Single-source A distance-based FSL method
-MAML (Chen et al., 2019) Single-source A learning-based FSL method
-DAPN (Zhao et al., 2021) Single-source A cross-domain FSL method using adversarial ProtoNet
-SemiCMT (Chen et al., 2024) Single-source A cross-modal method using contrastive alignment
-GPT2 (Zhou et al., 2023) Single-(LLM) A large language model based method using fine-tuning
-MDDA (Zhao et al., 2020) Multi-source A cross-domain FSL method using knowledge distillation
-MCW (Lee et al., 2019) Multi-source A cross-domain FSL method using maximal correlation analysis
-MetaSense (Gong et al., 2019) Oracle-(GT) An advanced FSL method based on MAML with sufficient target data
+  Transfer Learning (TL) (Jiang et al., 2022) Single-source A standard transfer learning method that reuses a source model by freezing all layers.
+Fine-tuning (FT) (Oh et al., 2022) Single-source A conventional fine-tuning method that updates all layers of a pre-trained model using available target data.
+ProtoNet (Chen et al., 2019) Single-source A distance-based FSL method that constructs class prototypes by averaging support set and classifies query set in the feature space.
+MAML (Chen et al., 2019) Single-source A gradient-based meta-learning method that learns a source model with shared initialization parameters from scratch.
+DAPN (Zhao et al., 2021) Single-source A cross-domain FSL method that uses adversarial ProtoNet with a domain discriminator to reduce latent feature distribution shift.
+SemiCMT (Chen et al., 2024) Single-source A cross-modal learning method that applies contrastive self-supervised learning during model adaptation.
+GPT2 (Zhou et al., 2023) Single-(LLM) A LLM-based method that enables in-context learning after fine-tuning on few target data.
+MDDA (Zhao et al., 2020) Multi-source A multi-domain FSL method that uses adversarial discriminative adaptation and multi-source knowledge distillation.
+MCW (Lee et al., 2019) Multi-source A multi-domain FSL method that uses maximal correlation analysis to fuse multiple source domains for few-shot model adaptation.
+MetaSense (Gong et al., 2019) Oracle-(GT) An advanced FSL method based on MAML that uses extensive target data to train from scratch, serving as an oracle upper bound.
+6 Experiment
+6.1 Experiment setup
 Table 2: Source and target datasets specifications.
-  Dataset Modality Subject Class Type Shape Privacy
+  Dataset Modality Subject Class Train Size Type Shape Privacy
   Source
-miniImageNet (Vinyals et al., 2016) Image - 100 - 3x84x84 Public
-miniDomainNet (Zhou et al., 2021) Image - 126 - 3x224x224 Public
-Office-31 (Saenko et al., 2010) Image - 31 - 3x224x224 Public
-Office-Home (Venkateswara et al., 2017) Image - 65 - 3x224x224 Public
-Caltech-101 (Fei-Fei et al., 2004) Image - 101 - 3x224x224 Public
-Newsgroup (Thulasidasan et al., 2021) Text - 20 - 100x1x1000 Public
-VoxCeleb (Nagrani et al., 2017) Audio - 100 Time-S 1x512x300 Public
-OPPORTUNITY (Roggen et al., 2010) IMU 12 11 Time-S 77x1x100 Public
+miniImageNet (Vinyals et al., 2016) Image - 100 50k - 3x84x84 Public
+miniDomainNet (Zhou et al., 2021) Image - 126 137.5k - 3x224x224 Public
+Office-31 (Saenko et al., 2010) Image - 31 3.3k - 3x224x224 Public
+Office-Home (Venkateswara et al., 2017) Image - 65 12.7k - 3x224x224 Public
+Caltech-101 (Fei-Fei et al., 2004) Image - 101 7.3k - 3x224x224 Public
+Newsgroup (Thulasidasan et al., 2021) Text - 20 15k - 100x1x1000 Public
+VoxCeleb (Nagrani et al., 2017) Audio - 100 14.9k Time-S 1x512x300 Public
+OPPORTUNITY (Roggen et al., 2010) IMU 12 11 3.7k Time-S 77x1x100 Public
 Target
-HHAR (Stisen et al., 2015) IMU 9 6 Time-S 6x1x256 Public
-WESAD (Schmidt et al., 2018) ECG 15 3 Time-S 10x1x256 Public
-Gesture IMU/PPG 10 8 Time-S 7x1x200 Private
-Writing Ultrasound 10 10 Time-S 1x1x256 Private
-Emotion mmWave 10 7 Spectrum 3x224x224 Private
-BP mmWave 10 - Doppler 1x1x499 Private
-ChestX (Wang et al., 2017) Image - 5 - 1x1024x1024 Public
+HHAR (Stisen et al., 2015) IMU 9 6 - Time-S 6x1x256 Public
+WESAD (Schmidt et al., 2018) ECG 15 3 - Time-S 10x1x256 Public
+Gesture IMU/PPG 10 8 - Time-S 7x1x200 Private
+Writing Ultrasound 10 10 - Time-S 1x1x256 Private
+Emotion mmWave 10 7 - Spectrum 3x224x224 Private
+BP mmWave 10 - - Doppler 1x1x499 Private
+ChestX (Wang et al., 2017) Image - 5 - - 1x1024x1024 Public
+We benchmark XTransfer against SOTA baselines, grouped into single- and multi-source methods (Table 1).
+Evaluation is conducted on 8 widely-used public datasets under cross-modality FSL settings (Table 2), using their corresponding public pre-trained models as sources, from diverse modalities (e.g., image, text, audio, and sensing) with various backbones such as ResNet18 from PyTorch Hub (Hub, 2022) (Table 3).
+To evaluate XTransfer, we develop 4 real-world applications and build testbeds on 3 commercial edge devices (e.g., smartwatch, smartphone, Raspberry Pi, Appendix Table 5 for details).
+Based on the testbeds, we collect 4 human sensing datasets (marked as Private in Table 2) as targets, from 40 participants in real-world settings, with ethics approval from the Human Research Ethics Committee of our institute.
+Additionally, we include 2 public sensing datasets and 1 image dataset as targets for standard benchmarking.
+We apply standard Leave-One-Out Cross-Validation (LOOCV) (Gong et al., 2019) for evaluating human sensing datasets, and adopt accuracy-to-resource (ATR) ratio, defined as $ATR=Accuracy\mathbin{/}(\alpha Norm(FLOPs))+(1-\alpha)Norm(Params))$ for evaluating resource-accuracy efficiency, where $\alpha$ is set to 0.5 by default (Appendix A for more details on our experiment setup).
+Table 3: DL model backbone specifications.
   Backbone Layers L-Blocks GFLOPs Params Kernel Shape
@@
 GPT2 (Zhou et al., 2023) 4x6 6 0.41 82.9M 1D By Input
-Table 3: DL model backbone specifications.
-6.1 Experiment setup
-We benchmark XTransfer against SOTA baselines, grouped into single- and multi-source methods (Table 1).
-Evaluation is conducted on 8 widely-used public datasets under cross-modality FSL settings (Table 2), using their corresponding public pre-trained models as sources, from diverse modalities (e.g., image, text, audio, and sensing) with various backbones such as ResNet18 from PyTorch Hub (Hub, 2022) (Table 3).
-To evaluate XTransfer, we develop 4 real-world applications and build testbeds on 3 commercial edge devices (e.g., smartwatch, smartphone, Raspberry Pi).
-Based on these, we collect 4 human sensing datasets (marked as Private in Table 2) as targets, from 40 participants in real-world settings, with ethics approval from the Human Research Ethics Committee of our institute.
-Additionally, we include 2 public sensing datasets and 1 image dataset as targets for standard benchmarking.
-We apply standard Leave-One-Out Cross-Validation (LOOCV) (Gong et al., 2019) for evaluating human sensing datasets, and adopt accuracy-to-resource (ATR) ratio, defined as $ATR=Accuracy\mathbin{/}(\alpha Norm(FLOPs))+(1-\alpha)Norm(Params))$ for evaluating resource-accuracy efficiency, where $\alpha$ is set to 0.5 by default (see Appendix A for details).
 6.2 XTransfer performance
-Figure 4: Ablation study evaluating the performance of model repairing and layer recombining.
 Repair performance.
-We first evaluate how our Repair loss reduces MMC shift to enhance the S-score of each layer with different MMC dimensionalities in the SRR pipeline.
-We use HHAR and the standard pre-trained ResNet18 on miniImageNet in a 5-shot setting.
-The channel removal is disabled in this experiment.
-We test with 4 different levels of MMC dimensionality (i.e., number of channels) ranging from 64 to 512.
+We first evaluate how our Repair loss reduces MMC shift to enhance each layer’s S-score with different MMC dimensionalities.
+We use HHAR and the standard pre-trained ResNet18 on miniImageNet under 5-shot setting.
+The layer channel removal is disabled.
+We test with 4 different levels of MMC dimensionality (i.e., number of layer channels) ranging from 64 to 512.
 We employ N-Pair loss (Zhang et al., 2023c) and Triplet loss (Weller et al., 2022) in FSL sensing tasks as baselines.
@@
 It indicates that Repair loss effectively enhances layer performance.
+Figure 4: Ablation study evaluating the performance of model repairing and layer recombining.
 Efficient search performance.
@@
 We also utilize MetaSense in Table 1 as the oracle baseline.
-We compare different configurations: ResNet18 (Single), Multi-ResNet18 (Multi) without efficient search enabled, and Multi-ResNet18 (Multi-Efficient) with efficient search enabled.
-As expected, Figure 4(b) shows that Multi-Efficient achieves a substantial time reduction ranging from 2.1 to 4 times compared to Multi in 5- and 10-shot settings, while outperforming the oracle baseline.
-Additionally, Multi-Efficient improves average accuracy by 7.8% over Multi across the three settings.
-While Multi-Efficient takes a longer search time than Single, it successfully achieves superior accuracy, highlighting the benefits of using multiple source models in XTransfer.
+We compare different configurations: ResNet18 (Single) and Multi-ResNet18 (Multi) without efficient search enabled, Multi-ResNet18 (Multi-Pre) with only pre-search check using top-1 layer selection enabled, and Multi-ResNet18 (Multi-Efficient) with both pre-search check and dynamic search range enabled.
+As shown in Figure 4(b), Multi-Pre achieves a lowest search time on average across all n-shot settings, but it leads to degraded accuracy by 3.36% and 7.94% on average, respectively, compared to Multi and Multi-Efficient.
+It also fails to reach oracle-level accuracy and presents the highest standard deviation on average.
+The results indicate that relying solely on the pre-search check introduces search stability issues due to inaccurate or drifted repair value estimation.
+In contrast, Multi-Efficient provides a strong balance between accuracy and efficiency.
+It reduces search time by 2.1 to 4 times compared to Multi in 5- and 10-shot settings, while outperforming the oracle baseline.
+While Multi-Efficient takes a longer search time than both Single and Multi-Pre, it successfully achieves superior accuracy and search stability, highlighting the benefits of using multiple source models and dynamic search range mechanism.
+Notably, even the search space is expanded 5 times more compared to Single, Multi-Efficient requires only 2.1 times more search time, indicating strong scalability enabled by the efficient search.
 The results also reveal that both Multi-Efficient and Multi achieve lower standard deviation on average compared to Single, suggesting that using multiple source models offers commendable search stability.
-In short, the results prove that the proposed efficient search significantly accelerates the layer search process and achieves better model accuracy.
-Impact of different sources.
-Since the selection of pre-trained source models plays a crucial role in XTransfer, we evaluate how different source models from various modalities affect the accuracy and resource overhead of the output models.
-We use HHAR, 2 pre-trained ResNet18s on both miniImageNet (Image) and VoxCeleb (Audio), and 2 pre-trained ResNet18-1Ds on both Newsgroup (Text) and OPPORTUNITY (Sensing), as shown in Table 2.
-We keep the other experimental setups unchanged.
-Figure 4(c) shows that the output models using Image source achieve the highest accuracy of 70.5% on average, compared to that of 69.6%, 64.3%, and 67% achieved by the models using Text, Audio, Sensing as sources, respectively.
-This suggests that the latent features learned from Image have better discriminability than those from other modalities.
-Since XTransfer prioritizes accuracy, we select Image source models by default.
-More ablation study results refer to Appendix C.
-6.3 Comprehensive comparison
-Cross-modality FSL performance.
-We evaluate the overall performance of XTransfer regarding both accuracy and ATR across 5 sensing datasets and 1 image dataset as targets (Table 2).
-We compare XTransfer against a range of SOTA baselines shown in Table 1.
-We also mark our methods as Our-Single and Our-Multi, using pre-trained ResNet18 on miniImageNet and Multi-ResNet18 with 5 corresponding image datasets as sources.
-In addition, we evaluate the fine-tuning performance of GPT2-small for comparison and align the parameters with the work (Zhou et al., 2023).
-All methods on target datasets are evaluated under the same cross-modality FSL settings (see Appendix A), with MetaSense (Gong et al., 2019) serving as the oracle baseline, trained on Conv5 with sufficient target sensor data (Table 3).
-For consistency, we also train the oracle using ResNet18 on the Emotion dataset to match 2D kernels.
-As shown in Table 4 and 5, both our methods outperform the baselines regarding model accuracy and ATR, and successfully reach or surpass the oracle in model accuracy across all target datasets in 5- and 10-shot settings.
-Specifically, both Our-Single and Our-Multi achieve a significantly higher ATR on average, surpassing the existing single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively.
-In 3-shot setting, our method achieves slightly lower accuracy than the oracle on HHAR, Gesture, and Emotion.
-This indicates that extremely low samples may not fully represent the test data distribution, leading to less stable feature alignment.
-Since the layers are reused from pre-trained sources, the performance also depends on their quality and semantic overlap with targets.
-Notably, the flexibility of XTransfer allows further gains in such low-shot cases by using additional source models with higher quality or closer relevance.
-In summary, the results show that XTransfer achieves SOTA accuracy while substantially reducing data and resource costs, providing a scalable and adaptable method for human sensing at the edge.
-Training resource usage.
-We also examine the training resource usage of XTransfer by measuring GPU utilization (Util) in the cloud.
-We focus on the above baselines that require model re-training or adaptation for transfer, e.g., MAML (Chen et al., 2019), DAPN (Zhao et al., 2021), and SemiCMT (Chen et al., 2024)), excluding methods that rely solely on fine-tuning or distance-based techniques.
-The experimental setup remains the same as above.
-As shown in Table 6, both our methods achieve the lowest GPU Util across all n-shot settings compared to baselines.
-Notably, both Our-Single and Our-Multi achieve a substantial reduction in GPU Util by 2.6 to 3.6 times and 2.5 to 3.5 times on average, respectively, compared to baselines.
-This is due to our layer-wise method, which progressively repairs and recombines only a subset of layers rather than training the entire model at once.
-It also highlights the effectiveness of our compact connector design.
-Despite leveraging multiple source models, Our-Multi shows no significant increase (e.g., only 3.67%) in training resource usage, further highlighting its efficiency and scalability.
-Hence, XTransfer significantly reduces cloud resource usage, enabling scalable model transfer.
+In short, the results prove that the proposed efficient search significantly accelerates the layer search process while improving accuracy and stability.
 HHAR (%) WESAD (%) Gesture (%) Writing (%) Emotion (%) ChestX (%)
@@
 Table 4: Comparison in model accuracy. Top-2 results are highlighted in bold.
-HHAR WESAD Gesture Writing Emotion ChestX
-Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
-ProtoNet (Chen et al., 2019) 0.51 0.46 0.50 0.44 0.49 0.52 0.51 0.56 0.59 0.69 0.79 0.73 0.51 0.49 0.55 0.22 0.23 0.25
-DAPN (Zhao et al., 2021) 0.49 0.51 0.55 0.57 0.61 0.63 0.46 0.49 0.51 0.75 0.79 0.80 0.48 0.50 0.59 0.24 0.24 0.26
-MAML (Chen et al., 2019) 0.37 0.42 0.42 0.40 0.58 0.66 0.37 0.41 0.39 0.32 0.40 0.35 0.22 0.27 0.31 0.22 0.20 0.22
-SemiCMT (Chen et al., 2024) 0.33 0.39 0.49 0.37 0.40 0.44 0.24 0.34 0.43 0.27 0.39 0.48 0.24 0.34 0.47 0.25 0.26 0.28
-GPT2 (Zhou et al., 2023) 0.11 0.09 0.12 0.13 0.09 0.09 0.13 0.14 0.17 0.14 0.19 0.20 - - - - - -
-Our-Single 1.42 1.57 2.09 2.76 2.61 2.51 1.55 1.90 1.76 2.11 1.88 1.84 0.82 1.07 1.26 0.83 1.04 1.24
-MDDA (Zhao et al., 2020) 0.03 0.04 0.04 0.08 0.08 0.08 0.02 0.03 0.03 0.03 0.02 0.02 0.03 0.03 0.03 0.04 0.04 0.04
-MCW (Lee et al., 2019) 0.06 0.05 0.05 0.06 0.07 0.06 0.04 0.05 0.05 0.06 0.06 0.06 0.05 0.05 0.05 0.04 0.04 0.05
-Our-Multi 1.71 1.47 1.64 3.00 2.70 2.75 1.20 1.25 1.63 1.93 1.96 1.53 0.96 1.23 1.73 0.68 1.24 0.83
-Table 5: Comparison in ATR ratio. Top-2 results are highlighted in bold
-HHAR (%) WESAD (%) Gesture (%) Writing (%) Emotion (%) ChestX (%)
-Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
-DAPN (Zhao et al., 2021) 48.85 49.19 51.79 44.62 49.49 48.81 42.73 48.71 50.59 43.88 56.63 57.51 44.67 59.73 63.12 46.66 47.32 48.48
-MAML (Chen et al., 2019) 59.59 79.32 89.72 52.58 70.27 77.52 53.11 66.75 73.21 72.56 85.47 92.22 63.39 72.78 75.01 44.52 55.46 61.36
-SemiCMT (Chen et al., 2024) 55.60 73.43 78.91 38.83 50.33 67.89 57.02 72.16 98.04 75.14 79.12 81.79 60.30 63.43 68.60 41.54 50.05 63.43
-Our-Single 18.27 18.90 19.48 15.95 16.46 16.98 20.67 21.71 22.51 19.92 20.67 21.08 18.24 19.20 19.96 16.97 17.58 18.80
-Our-Multi 18.31 20.77 22.06 18.00 19.42 23.10 20.49 22.49 24.49 18.68 19.25 21.69 16.45 18.18 19.85 15.33 17.69 19.62
-Table 6: Comparison in GPU utilization of each method during training. Top-2 results are highlighted in bold
+Impact of different sources.
+Since the selection of pre-trained source models plays a crucial role in XTransfer, we evaluate how different source models from various modalities affect the accuracy and resource overhead of the output models.
+We use HHAR, 2 pre-trained ResNet18s on both miniImageNet (Image) and VoxCeleb (Audio), and 2 pre-trained ResNet18-1Ds on both Newsgroup (Text) and OPPORTUNITY (Sensing), as shown in Table 2.
+We keep the other experimental setups unchanged and include 2-shot settings.
+Figure 4(c) shows that the output models using Image source achieve the highest accuracy of 70.5% on average in 3- to 10-shot settings, compared to that of 69.6%, 64.3%, and 67% achieved by using Text, Audio, Sensing sources, respectively.
+This suggests that the latent features learned from Image, trained on larger datasets with broader class semantics, provide higher discriminability than those learned from other modalities.
+Notably, Image achieves lower average accuracy in 2- and 3-shot settings, suggesting that extremely low target data may not fully represent the test data distribution, leading to less stable alignment when using Image source.
+In contrast, both Text and Sensing result in higher average accuracy and reach the oracle, indicating that performance varies with source model quality and semantic relevance to the target.
+These results also show that XTransfer can flexibly reuse diverse source models across modalities to improve ultra low-shot performance.
+Moreover, the output models using Text or Sensing sources require only 0.16G FLOPs on average, as 1D kernels require much less computation than 2D kernels.
+Since XTransfer prioritizes accuracy, we select Image source models by default.
+More ablation results see Appendix C.
+6.3 Comprehensive comparison
+Cross-modality FSL performance.
+We evaluate the overall performance of XTransfer regarding both accuracy and ATR across 5 sensing datasets and 1 image dataset as targets (Table 2).
+We compare XTransfer against a range of SOTA baselines shown in Table 1.
+We report two variants with different sources: Our-Single, using a ResNet18 pre-trained on miniImageNet, and Our-Multi, using Multi-ResNet18 with five corresponding image datasets.
+We additionally report GPT2-small fine-tuning for comparison, following the parameter setting in (Zhou et al., 2023).
+All methods on target datasets are evaluated under the same cross-modality FSL settings (Appendix A.1), where only few target data are fine-tuned with sources.
+MetaSense (Gong et al., 2019) serves as the oracle, trained on Conv5 with sufficient target sensor data (Table 3).
+For consistency, we also train the oracle using ResNet18 on the Emotion dataset to match 2D kernels.
+As shown in Table 4 and Appendix Table 12, both our variants outperform the baselines regarding model accuracy and ATR, and successfully reach or surpass the oracle in model accuracy across all target datasets in 5- and 10-shot settings.
+Specifically, both Our-Single and Our-Multi achieve a significantly higher ATR on average, surpassing the existing single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively.
+While our variants are slightly below the oracle on HHAR, Gesture, and Emotion under 3-shot settings, XTransfer can further improve by reusing source models with higher quality or closer semantic relevance, as shown in our source-impact study.
+Overall, XTransfer achieves SOTA accuracy while substantially reducing data and resource costs, providing a scalable and adaptable method for human sensing at the edge.
+More results see Appendix C.
 7 Conclusion
-This paper presents XTransfer, a novel method that enables modality-agnostic few-shot model transfer with resource-efficient design for human sensing applications.
-XTransfer harnesses the power of pre-trained models as free sources and requires only few sensor data to create compact models.
-It significantly reduces the costs associated with sensor data collection, model training, and edge deployment, offering a scalable and adaptable method for advancing human sensing.
+This paper presents XTransfer, a novel method that enables modality-agnostic few-shot model transfer with resource-efficient design.
+XTransfer reuses the power of pre-trained models as free sources and requires only few sensor data to restructure models.
+It significantly reduces the costs associated with sensor data collection, model training, and edge deployment, offering a scalable and adaptable method for advancing human sensing applications.
 Appendix
 The appendix is organized as follows:
-  • $\bullet$
   • Section A illustrates more details on our experiment setup, including implementation details, developed real-world sensing applications, experimental setups, dataset statistics, evaluation metrics, and computation resources we used for our experiments.
-  • $\bullet$
   • Section B introduces additional technique details of different components in XTransfer.
-  • $\bullet$
   • Section C gives more experimental results about ablation study and benchmarking.
@@
 Unlike FSL, cross-domain FSL focuses primarily on meta-testing and transfers pre-trained source models across domains (i.e., different users), while remaining within the same modality (Oh et al., 2022).
-In this work, the cross-modality FSL settings are defined similarly to cross-domain FSL, but extend it to transfer pre-trained source models across different modalities.
+In this work, the cross-modality FSL settings are defined similarly to cross-domain FSL, but extend it to transferring pre-trained source models across different modalities, using only few labeled target data and without assuming any additional unlabeled target data.
 A.2 Setups in preliminary study
 As shown in Figure 2(a), we test a standard overfitting metric, defined as $\frac{\left|Accuracy_{train}-Accuracy_{test}\right|}{Accuracy_{train}}$ (López et al., 2022).
-In Figure 2(b), we test 3 source models (using a pre-trained ResNet18 from PyTorch Hub (Hub, 2022)) from 3 modalities, e.g., miniImageNet (Image), Newsgroup (Text), and OPPORTUNITY (Sensing), shown in Table 2.
+In Figure 2(b), we apply a standard cross-domain FSL benchmarking (Oh et al., 2022) to gain insights.
+It uses Earth Mover’s Distance (EMD) to measure similarity between source models and targets.
+We test 3 source models (using a pre-trained ResNet18 from PyTorch Hub (Hub, 2022)) from 3 modalities, e.g., miniImageNet (Image), Newsgroup (Text), and OPPORTUNITY (Sensing), shown in Table 2.
 We also use 5 sensing datasets (HHAR, WESAD, Writing, Emotion, and Gesture) as targets, and employ 2 well-known image datasets (CUB and ChestX) as reference targets to the Image, detailed in Table 2.
-We use a default 5-shot cross-modality FSL setting with LOOCV enabled.
+It also assesses FSL difficulty by fine-tuning source models on targets under 5-shot settings with LOOCV enabled.
+Since source models usually involve a large number of classes (e.g., 100 classes in miniImageNet), we propose a class pairing strategy that pairs source and target classes exhibiting similar latent feature distributions across modalities.
 A.3 Setups in layer-wise analysis
@@
 A.4 Layer-wise metric
-To precisely track the MMC shift changes at each layer during fine-tuning, we essentially use the Silhouette score (S-score) (Shahapure & Nicholas, 2020) as the layer-wise metric to evaluate how well the features are being fine-tuned.
+To precisely track the MMC shift changes at each layer during fine-tuning, we essentially use the Silhouette score (S-score) (Shahapure and Nicholas, 2020) as the layer-wise metric to evaluate how well the features are being fine-tuned.
 The S-score is in the range from -1 to 1 and being better if closing to 1.
@@
 It uses bilinear interpolation (Mastyło, 2013) (i.e., Resizer) to resize the height and width, and a fixed convolutional layer (i.e., Fixed header) with a $1\times 1$ kernel to match the input channels (Szegedy et al., 2014), placed between source models and sensor data.
-B.2 PCA-based channel removal
-In the Removal stage, channel removal is designed to further enhance layer-wise performance while optimizing resource usage.
-This is achieved by removing unnecessary channels at each layer.
-Existing pruning techniques (Shen et al., 2022; Frankle & Carbin, 2019) typically rely on channel importance metrics, such as the L2 norm, calculated from MMC.
-However, due to MMC shift, these metrics may inaccurately assess channel importance, resulting in degraded layer performance.
-To address this, we propose using PCA-based channel importance, derived from the optimized component weights during the repair process, as a more reliable metric for channel removal.
-Our objective is to remove unnecessary channels $\widetilde{ch}$ while preserving the maximum S-score ($Score(\cdot)$), formulated as $\arg\max_{(\widetilde{ch}\,\in CH)}Score(PCA(MMC(ft^{r}_{ij}),\widetilde{ch}),Y_{t})$, where $CH$ denotes the set of all channels, $r$ denotes the repaired latent features $ft_{ij}$, and $Y_{t}$ denotes the target labels.
+B.2 PCA-based layer channel removal
+In the Removal stage, layer channel removal is designed to further enhance layer-wise performance while optimizing resource usage.
+This is achieved by removing unnecessary layer channels at each layer.
+Existing pruning techniques (Shen et al., 2022; Frankle and Carbin, 2019) typically rely on layer channel importance metrics, such as the L2 norm, calculated from MMC.
+However, due to MMC shift, these metrics may inaccurately assess layer channel importance, resulting in degraded layer performance.
+To address this, we propose using PCA-based layer channel importance, derived from the optimized component weights during the repair process, as a more reliable metric for layer channel removal.
+Our objective is to remove unnecessary layer channels $\widetilde{ch}$ while preserving the maximum S-score ($Score(\cdot)$), formulated as $\arg\max_{(\widetilde{ch}\,\in CH)}Score(PCA(MMC(ft^{r}_{ij}),\widetilde{ch}),Y_{t})$, where $CH$ denotes the set of all layer channels, $r$ denotes the repaired latent features $ft_{ij}$, and $Y_{t}$ denotes the target labels.
 B.3 Setups for efficient search insights
@@
 Thus, we design the repair rate growth model based on an exponential growth model (see Section 5.2).
-B.5 Dynamic search range mechanism
-Based on the estimated repair rate, LWS control can initially use a small, fixed search range to efficiently filter out most layer candidates before initiating SRR processing.
-However, because the updates to $rate_{n}^{est}$ are limited, the estimation may be inaccurate, potentially affecting selection decisions (e.g., missing the optimal layer candidate).
-To mitigate this risk, we develop a dynamic search range mechanism that adjusts the search range to balance accuracy and efficiency.
-This mechanism is formulated as $range_{n}=(1+\hat{u}^{\pm}rate_{n}^{est})*range_{n-1},\,range_{n}\in[0.2,1]$, where the range starts at 0.5 and adjusts based on the previous range.
-Here, $\hat{u}$ denotes a signed unit vector.
-If $rate_{n-1}^{est}$ is less than $rate_{n-1}$, it indicates that the estimation might be inaccurate and $\hat{u}$ will be positive to increase the range in searching for the optimal layer.
-Otherwise, the range will be reduced to improve search efficiency.
-B.6 Post fine-tuning
+B.5 Post fine-tuning
 After LWS control is completed, we add either a linear classifier or a regression layer (e.g., a fully connected layer) at the end of the model to finalize its backbone.
@@
 Figure 7: Ablation study evaluating the performance of components, search parameters, and applications.
-C.4 Removal performance
-We next evaluate how channel removal enhances the average S-score of selected layers.
+C.4 Non-linear feature alignment performance
+We essentially use PCA to construct a fixed, orthogonal anchor feature space in which alignment is measured and the anchor-based repair loss is computed at each layer across modalities.
+The proposed connector (i.e., repair generator) dynamically maps target-modality features into this anchor space, enabling the non-linear feature alignment.
+To examine the impact of using a linear PCA anchor space and to compare the alignment performance, we evaluate our repair loss with and without PCA against a Maximum Mean Discrepancy (MMD) loss baseline (Wang et al., 2023).
+To set up, we keep using the standard ResNet18 on miniImageNet (Table 3) as sources on HHAR as target (Table 2) in 3-, 5- and 10-shot settings with all configurations identical except for the alignment loss.
+As shown in Table 9, our repair loss with PCA anchor space achieves the highest accuracy of 70.5% and the lowest ATR of 1.69 on average across all n-shot settings.
+Compared to our repair loss without PCA, MMD loss performs a lower accuracy on average.
+This indicates that MMD loss is limited to aligning the latent feature distribution when the feature space is highly complex due to strong cross-modality discrepancies and only few target data.
+In addition, adversarial alignment methods such as DAPN (Zhao et al., 2021) and MDDA (Zhao et al., 2020) also face similar difficulties under these conditions, as shown in Tables 4 and 12.
+Moreover, the PCA anchor space provides a more stable basis for alignment and consistently enhances our repair loss performance.
+As a result, these findings further validate the effectiveness of our method in cross-modality FSL settings.
+Alignment Method 3-shot 5-shot 10-shot
+Accuracy (%) ATR Accuracy (%) ATR Accuracy (%) ATR
+MMD Loss (Wang et al., 2023) 55.78 1.63 65.89 2.01 68.52 1.83
+Repair Loss 65.11 2.33 69.33 2.45 72.07 2.29
+PCA + Repair Loss 58.96 1.42 71.85 1.56 80.74 2.09
+Table 9: Comparison of non-linear feature alignment methods across n-shot settings.
+C.5 Layer-wise convergence comparison
+To further examine layer-wise convergence across the above alignment methods, we observe the feature alignment process in the initial layer.
+To normalize heterogeneous losses, we follow the strategy in (Li et al., 2020) and compute the loss descent (LD) relative to the initial loss, enabling a fair comparison of convergence rates across different loss functions (i.e., how much each loss decreases from its starting point), defined as $LD(t)=(L(0)-L(t))\mathbin{/}L(0)\times 100\%$, where $t$ denotes the epoch index.
+All experimental settings remain the same as in the previous setup.
+As shown in Table 10, our repair loss with the PCA anchor achieves the largest loss descent of 28.87%, which is 1.79 and 1.87 times faster than the repair loss without PCA and the MMD loss, respectively.
+Interestingly, MMD loss presents a large initial drop but quickly plateaus, while our repair loss starts more gradually and accelerates over training.
+In contrast, the PCA-based repair maintains a steady and consistent descent across epochs, highlighting the benefits of aligning features in a fixed, orthogonal anchor space.
+These findings further support that our method provides both effective feature alignment and practical training efficiency, consistent with the efficiency study in Section C.10.
+Epoch (#) #10 #20 #30 #40 #50 #60 #70 #80 #90 #100
+MMD Loss (Wang et al., 2023) 6.46% 10.0% 11.63% 12.81% 13.5% 14.3% 14.66% 14.98% 15.27% 15.41%
+Repair Loss 0.03% 1.67% 2.66% 4.93% 5.26% 7.16% 8.06% 10.4% 14.66% 16.14%
+PCA + Repair Loss 5.12% 13.45% 16.54% 19.23% 20.55% 21.3% 25.96% 26.21% 27.6% 28.87%
+Table 10: Comparison of layer-wise convergence rates across alignment methods.
+C.6 Removal performance
+We next evaluate how layer channel removal enhances the average S-score of selected layers.
 We continue to use the same setup as above under 5-shot settings.
-We employ L2 norm-based (i.e., L2-based) model pruning (Frankle & Carbin, 2019; Shen et al., 2022) as the baseline and compare it with the setting where channel removal is disabled (i.e., Non-Pruning).
-Figure 7(a) demonstrates that the proposed channel removal outperforms both Non-Pruning and L2-based pruning across all settings, achieving an average S-score of 0.1 by selected layers.
-L2-based pruning achieves a lower average S-score than Non-Pruning, indicating that the channel importance may be inaccurately estimated, resulting in degraded performance.
-These results show that the proposed channel removal further improves layer performance by accurately removing unnecessary channels at each layer.
-C.5 Impact of search depth
+We employ L2 norm-based (i.e., L2-based) model pruning (Frankle and Carbin, 2019; Shen et al., 2022) as the baseline and compare it with the setting where layer channel removal is disabled (i.e., Non-Pruning).
+Figure 7(a) demonstrates that the proposed layer channel removal outperforms both Non-Pruning and L2-based pruning across all settings, achieving an average S-score of 0.1 by selected layers.
+L2-based pruning achieves a lower average S-score than Non-Pruning, indicating that the layer channel importance may be inaccurately estimated, resulting in degraded performance.
+These results show that the proposed layer channel removal further improves layer performance by accurately removing unnecessary layer channels at each layer.
+C.7 Impact of search depth
 In this experiment, we investigate how different search depths affect the accuracy and resource overhead (e.g., FLOPs) of output models.
@@
 Therefore, we set D3 as the default search depth, as it provides the best balance between resource overhead and accuracy of the output models.
-C.6 Impact of model backbones
+C.8 Impact of model backbones
 The current implementation of XTransfer uses homogeneous pre-trained source models.
@@
 Table 11: Performance of different source model backbones in terms of accuracy and FLOPs.
-C.7 Training efficiency
-To test the practical training efficiency in the cloud, we further evaluate training time across the selected baselines.
-We use the same experimental setups as in Table 6, Section 6.3.
-For MAML (Chen et al., 2019) and SemiCMT (Chen et al., 2024), which require model re-training from pre-trained source models to adapt to the target modality classes, training time includes both re-training and fine-tuning.
-For DAPN (Zhao et al., 2021), it reflects model adaptation process on few target data.
-The results in Table 12 show that both Our-Single and Our-Multi achieve significant training time reductions compared to MAML.
-Our-Single also matches or exceeds DAPN and SemiCMT in training efficiency.
-Notably, even when using up to 5 source models, Our-Multi remains practical training efficiency, requiring only 2.05 times more training time on average compared to Our-Single across all datasets and shot settings.
-This further supports the practicability and scalability of our method.
-HHAR (mins) WESAD (mins) Gesture (mins) Writing (mins) Emotion (mins) ChestX (mins)
-Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
-DAPN (Zhao et al., 2021) 2.38 4.62 11.78 1.49 2.54 6.78 3.36 6.12 15.72 4.12 7.71 20.97 2.99 7.34 18.86 2.14 4.57 12.04
-MAML (Chen et al., 2019) 124.28 125.66 183.20 71.71 77.07 115.87 118.79 129.79 191.95 165.06 129.27 194.36 94.81 98.40 153.20 66.20 72.17 106.97
-SemiCMT (Chen et al., 2024) 3.02 8.27 10.23 1.82 2.48 5.63 4.81 9.42 12.83 10.10 13.92 18.75 3.39 7.32 11.48 2.84 5.28 8.33
-Our-Single 3.98 5.57 7.95 2.12 2.99 4.33 5.64 7.80 11.24 7.94 10.54 14.69 5.58 7.43 9.88 3.65 4.75 6.84
-Our-Multi 9.25 15.33 15.25 4.40 5.38 7.21 12.49 15.12 24.37 15.35 17.36 32.38 11.72 13.39 19.75 8.41 9.33 13.61
-Table 12: Comparison in training time of each method.
-C.8 Regression task performance
+C.9 Regression task performance
 We also evaluate the effectiveness of XTransfer in a regression task using the BP dataset, which includes both systolic (SBP) and diastolic (DBP) sensor data shown in Table 2 collected by the blood pressure monitoring application.
@@
 These results demonstrate that XTransfer is highly efficient in handling regression tasks with few sensor data.
-C.9 On-device performance
+C.10 Training efficiency
+To test the practical training efficiency in the cloud, we further evaluate training time across the selected baselines.
+We use the same experimental setups as in Table 14, Section 6.3.
+For MAML (Chen et al., 2019) and SemiCMT (Chen et al., 2024), which require model re-training from pre-trained source models to adapt to the target modality classes, training time includes both re-training and fine-tuning.
+For DAPN (Zhao et al., 2021), it reflects model adaptation process on few target data.
+The results in Table 13 show that both Our-Single and Our-Multi achieve significant training time reductions compared to MAML.
+Our-Single also matches or exceeds DAPN and SemiCMT in training efficiency.
+Notably, even when using up to 5 source models, Our-Multi remains practical training efficiency, requiring only 2.05 times more training time on average compared to Our-Single across all datasets and shot settings.
+This further supports the practicability and scalability of our method.
+HHAR WESAD Gesture Writing Emotion ChestX
+Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
+ProtoNet (Chen et al., 2019) 0.51 0.46 0.50 0.44 0.49 0.52 0.51 0.56 0.59 0.69 0.79 0.73 0.51 0.49 0.55 0.22 0.23 0.25
+DAPN (Zhao et al., 2021) 0.49 0.51 0.55 0.57 0.61 0.63 0.46 0.49 0.51 0.75 0.79 0.80 0.48 0.50 0.59 0.24 0.24 0.26
+MAML (Chen et al., 2019) 0.37 0.42 0.42 0.40 0.58 0.66 0.37 0.41 0.39 0.32 0.40 0.35 0.22 0.27 0.31 0.22 0.20 0.22
+SemiCMT (Chen et al., 2024) 0.33 0.39 0.49 0.37 0.40 0.44 0.24 0.34 0.43 0.27 0.39 0.48 0.24 0.34 0.47 0.25 0.26 0.28
+GPT2 (Zhou et al., 2023) 0.11 0.09 0.12 0.13 0.09 0.09 0.13 0.14 0.17 0.14 0.19 0.20 - - - - - -
+Our-Single 1.42 1.57 2.09 2.76 2.61 2.51 1.55 1.90 1.76 2.11 1.88 1.84 0.82 1.07 1.26 0.83 1.04 1.24
+MDDA (Zhao et al., 2020) 0.03 0.04 0.04 0.08 0.08 0.08 0.02 0.03 0.03 0.03 0.02 0.02 0.03 0.03 0.03 0.04 0.04 0.04
+MCW (Lee et al., 2019) 0.06 0.05 0.05 0.06 0.07 0.06 0.04 0.05 0.05 0.06 0.06 0.06 0.05 0.05 0.05 0.04 0.04 0.05
+Our-Multi 1.71 1.47 1.64 3.00 2.70 2.75 1.20 1.25 1.63 1.93 1.96 1.53 0.96 1.23 1.73 0.68 1.24 0.83
+Table 12: Comparison in ATR ratio. Top-2 results are highlighted in bold
+HHAR (mins) WESAD (mins) Gesture (mins) Writing (mins) Emotion (mins) ChestX (mins)
+Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
+DAPN (Zhao et al., 2021) 2.38 4.62 11.78 1.49 2.54 6.78 3.36 6.12 15.72 4.12 7.71 20.97 2.99 7.34 18.86 2.14 4.57 12.04
+MAML (Chen et al., 2019) 124.28 125.66 183.20 71.71 77.07 115.87 118.79 129.79 191.95 165.06 129.27 194.36 94.81 98.40 153.20 66.20 72.17 106.97
+SemiCMT (Chen et al., 2024) 3.02 8.27 10.23 1.82 2.48 5.63 4.81 9.42 12.83 10.10 13.92 18.75 3.39 7.32 11.48 2.84 5.28 8.33
+Our-Single 3.98 5.57 7.95 2.12 2.99 4.33 5.64 7.80 11.24 7.94 10.54 14.69 5.58 7.43 9.88 3.65 4.75 6.84
+Our-Multi 9.25 15.33 15.25 4.40 5.38 7.21 12.49 15.12 24.37 15.35 17.36 32.38 11.72 13.39 19.75 8.41 9.33 13.61
+Table 13: Comparison in training time of each method.
+HHAR (%) WESAD (%) Gesture (%) Writing (%) Emotion (%) ChestX (%)
+Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
+DAPN (Zhao et al., 2021) 48.85 49.19 51.79 44.62 49.49 48.81 42.73 48.71 50.59 43.88 56.63 57.51 44.67 59.73 63.12 46.66 47.32 48.48
+MAML (Chen et al., 2019) 59.59 79.32 89.72 52.58 70.27 77.52 53.11 66.75 73.21 72.56 85.47 92.22 63.39 72.78 75.01 44.52 55.46 61.36
+SemiCMT (Chen et al., 2024) 55.60 73.43 78.91 38.83 50.33 67.89 57.02 72.16 98.04 75.14 79.12 81.79 60.30 63.43 68.60 41.54 50.05 63.43
+Our-Single 18.27 18.90 19.48 15.95 16.46 16.98 20.67 21.71 22.51 19.92 20.67 21.08 18.24 19.20 19.96 16.97 17.58 18.80
+Our-Multi 18.31 20.77 22.06 18.00 19.42 23.10 20.49 22.49 24.49 18.68 19.25 21.69 16.45 18.18 19.85 15.33 17.69 19.62
+Table 14: Comparison in GPU utilization of each method during training. Top-2 results are highlighted in bold
+C.11 Training resource usage
+We then examine the training resource usage of XTransfer by measuring GPU utilization (Util) in the cloud.
+We focus on the above baselines that require model re-training or adaptation for transfer, e.g., MAML (Chen et al., 2019), DAPN (Zhao et al., 2021), and SemiCMT (Chen et al., 2024)), excluding methods that rely solely on fine-tuning or distance-based techniques.
+The experimental setup remains the same as above.
+As shown in Table 14, both our methods achieve the lowest GPU Util across all n-shot settings compared to baselines.
+Notably, both Our-Single and Our-Multi achieve a substantial reduction in GPU Util by 2.6 to 3.6 times and 2.5 to 3.5 times on average, respectively, compared to baselines.
+This is due to our layer-wise method, which progressively repairs and recombines only a subset of layers rather than training the entire model at once.
+It also highlights the effectiveness of our compact connector design.
+Despite leveraging multiple source models, Our-Multi shows no significant increase (e.g., only 3.67%) in training resource usage, further highlighting its efficiency and scalability.
+Hence, XTransfer significantly reduces cloud resource usage, enabling scalable model transfer.
+C.12 On-device performance
 […]